### v1

# Modular Compilation for Quantum Chiplet Architectures

Abstract.
As quantum computing technology continues to mature, industry is adopting modular quantum architectures to keep quantum scaling on the projected path and meet performance targets. However, the complexity of chiplet-based quantum devices, coupled with their growing size, presents an imminent scalability challenge for quantum compilation. Contemporary compilation methods are not well-suited to chiplet architectures — in particular, existing qubit allocation methods are often unable to contend with inter-chiplet links, which don’t necessary support a universal basis gate set. Furthermore, existing methods of logical-to-physical qubit placement, swap insertion (routing), unitary synthesis, and/or optimization are typically not designed for qubit links of wildly varying levels of duration or fidelity. In this work, we propose SEQC, a complete and parallelized compilation pipeline optimized for chiplet-based quantum computers, including several novel methods for qubit placement, qubit routing, and circuit optimization. SEQC attains up to a 36% increase in circuit fidelity, accompanied by execution time improvements of up to 1.92×1.92$\times$1.92 ×. Additionally, owning to its ability to parallelize compilation, SEQC achieves consistent solve time improvements of 24×2-4$\times$2 - 4 × over a chiplet-aware Qiskit baseline.
Quantum computing, quantum compilation, modular architectures.
copyright: nonejournalyear: 2024doi: XXXXXXX.XXXXXXX
1. Introduction
Classical computer systems, through the decades of their existence, have become increasingly distributed. Physical, technological, and economic constraints have prevented single “monolithic” systems from scaling to the point that they could meet the demand for high performance and yet remain practical. Today, distributed architectures are prevalent, and their latest renditions, from cloud computing to chiplet-based digital processors, are ubiquitous.
We postulate that quantum computing is on a similar path. While improvements in superconducting quantum hardware have led to the debut of processors with $1000+$ qubits (ibm_condor), practical implementations of quantum computation will require millions of physical qubits (paler2019really). The number of qubits required by a quantum algorithm, the depth of the algorithm (i.e., execution time, or number of gates on the critical path), the auxiliary and syndrome qubits required to reach sufficiently low error rates, and the size of matter qubits when all control hardware is included, all suggest that many more qubits are required than are likely to fit in a single die (vanMeter_2010_DQC). The overhead for error correction, cooling, dilution, control systems, I/O lines, challenges associated with verification and adequate chip thermalization, and realistic resources such as lossy waveguides, limited qubit fabrication yields as qubit capacity grows (smith_2022_chiplets), and finite chip sizes, make the prospect of a single “monolithic” quantum processor very expensive or entirely unrealistic by current standards. These constraints force large-scale quantum systems to adopt a physically distributed architecture.
We are already observing signs of this shift. Recent developments in quantum chip linking, including flip-chip architectures (gold2021entanglement) and low-loss coaxial cables (niu2023low), suggest that modular designs are the most viable for scaling quantum computers (smith_2022_chiplets). Many contemporary or upcoming leading quantum systems adopt chiplet-based modular quantum processors, for example using high-bandwidth quantum links between nearest-neighbor quantum chiplets (e.g., carrier-chip couplers in Rigetti Aspen-M (Rigetti_AspenM; gold2021entanglement), or m-couplers in IBM Crossbill (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers)), or lower-fidelity, lower-bandwidth, but longer-distance flexible coupling of discrete chips (e.g., l-couplers in IBM Flamingo (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers)), or a combination of the above (e.g., c-, m-, and l-couplers in IBM Starling (Mandelbaum_2024_IBMCouplers)), or multi-chip connectivity through tunable couplers and routing chips (Field_2024_RigettiModular). Most major companies have set their sights at modular designs to meet quantum scaling targets in practical ways, and modular quantum processors dominate their latest roadmaps (IBM_2023_Roadmap; IonQ_2024_Roadmap; Rigetti_2024_Roadmap).
Unfortunately, current compilation infrastructure is not capable of reasoning well about this quantum interconnect heterogeneity. The mapping of a quantum program’s logical quantum gates into the native gates supported by the underlying quantum processor, and the mapping of logical qubits to the physical qubits of the quantum hardware, largely determine the number and type of native quantum gates required for the computation, the circuit depth, and its execution time—long programs may not complete successfully as qubits decohere. Different mappings can result in vastly different compiled quantum circuits, with diverse characteristics along all these axes. These details of efficiency can make or break an algorithm in today’s noisy intermediate-scale quantum systems (NISQ) era, so quantum compilers aggressively optimize for all of them. To make matters more challenging, each qubit, coupler, and gate have diverse error profiles that are highly variable both spatially and temporally (Tannu_2019_ErrorVariability; Murali_2019_NoiseMapping; dasgupta2021stability). Unlike classical compilation that is done only once for a given architecture, quantum programs must be recompiled every time before execution, as the ideal physical qubits to execute on change between runs. So, quantum compilers today are faced with the daunting task of optimizing quantum programs across multiple dimensions with often conflicting demands, in a continuously changing environment. Coupled with the hardness of synthesizing unitaries (which is exponential in the number of qubits) and the need for quick and frequent compilation, modern quantum compilers have no option other than to rely on heuristics to perform the task. These heuristics result in compilation complexity of $O(n^{2})$ for $n$ total qubits in a quantum processor.
Hardware modularity adds significant complexity to this already hard task, as inter-chiplet links are typically inferior compared to intra-chiplet ones (laracuente2022modeling; smith_2022_chiplets), connectivity across chiplets is often limited (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers), and not all basis gates are necessarily supported across chiplets (niu2023low). In fact, popular quantum software stacks today (e.g., Qiskit (qiskit2024)) are not even cognizant of the existence of hardware modularity. Hardware modularity, though, also presents an opportunity. Inspired by classical compilation, in this paper we leverage hardware modularity to achieve compilation modularity.
Compilation in classical systems is typically performed independently and in parallel for each source file, producing one object file per source. The individual object files are then linked together to construct the executable. We propose a compilation framework for modular quantum processors that works in a similar fashion: it stratifies, i.e., splits, the source quantum circuit into subcircuits small enough to fit in each chiplet, and maps subcircuits to chiplets, and then in parallel elaborates each subcircuit and compiles it for its target chiplet. This Stratify-Elaborate Quantum Compiler (SEQC) stratifies a source program only once for a given chiplet architecture, and performs only the elaboration step recurrently before each execution. In essence, SEQC replaces the recurrent $O(n^{2})$ compilation step for an $n$-qubit quantum processor, with several parallel $O(k^{2})$ elaboration steps for $k$-qubit chiplets. As the qubit capacity $n$ of quantum processors grows exponentially, today’s $O(n^{2})$ compilation latency rises even faster. We expect, however, that the number of qubits $k$ per chiplet will remain relatively stable or grow much slower, as it seems to be the case for the foreseeable future (IBM_2023_Roadmap), and thus the SEQC recurrent compilation latency is expected to remain relatively stable. The stratification step is $O(n^{2})$, so the end-to-end complexity remains the same, but stratification is performed only once; the recurrent compilation in SEQC is only $O(k^{2})$, and barely growing with new processor designs.
Additionally, as SEQC is cognizant of hardware modularity, it can stratify the source quantum circuit into subcircuits and map them to chiplets to minimize inter-chiplet communication. As we show in this paper, SEQC produces circuits with shorter execution times and significantly fewer inter-chiplet gates compared to today’s stock compilers, leading to much higher fidelity execution. More importantly, as the number of qubits in a processor grows, SEQC achieves even higher performance in these figures of merit.
In summary, the contributions of this paper are as follows:
  • We make stock compilers aware of hardware modularity, thereby allowing them to correctly compile circuits for modular architectures with limited cross-chiplet gate support.
  • We design and implement SEQC, a Stratify-Elaborate Quantum Compiler for modular architectures. SEQC performs compilation in two stages, with the first stage (stratification, or chiplet splitting) performed only once for a given architecture, and the second stage (elaboration, or chiplet compilation) performed in parallel for each chiplet. Only this second stage needs to be performed before each execution, and thus SEQC’s compilation time is largely unaffected by the growth of qubit counts in future quantum processors.
  • We design and implement in SEQC several novel methods for qubit placement, qubit routing, and circuit optimization.
  • We evaluate SEQC and show it compiles circuits with up to 36% higher circuit fidelity and up to 1.92×1.92$\times$1.92 × lower execution time, while consistently achieving 24×2-4$\times$2 - 4 × faster compilation time compared to a chiplet-aware Qiskit baseline.
2. Background and Motivation
To execute a program on a target device, the logical program representation must be transformed such that it conforms to the idiosyncratic physical constraints for that particular device. This process of logical-to-physical translation represents a critical part of compilation and is present even in exotic computing technologies, e.g., quantum. In quantum computing, programs are (usually) represented as quantum circuits, and devices are subject to domain-specific physical constraints, such as decoherence time (mike_and_ike), qubit topology (the connectivity between physical qubits) (qubit_allocation; qubit_routing), and the set of basis gates (the family of operations physically implemented on the device). Thus, quantum circuit compilation is necessary to actually run quantum programs on real-world devices.
Compilation is often divided into distinct stages such as qubit allocation (layout and routing), basis translation, and optimization. The relations between each stage and the physical constraints of both monolithic and chiplet architectures are described below.
2.1. Qubit Allocation
Qubit allocation addresses the topology constraints of quantum devices, where physical qubits on a device have limited connectivity to other qubits on the device. It is critical that after this stage, which often requires the modification or addition of circuit gates, the resulting optimized and technology-dependent circuit is functionally equivalent to the original, technology-independent algorithm (smith2019quantum). The qubit allocation problem has been previously shown to be NP-complete (qubit_allocation); thus, algorithms often adopt various strategies to make the problem more tractable. One common method is to further divide qubit allocation into two substages: layout and routing.
Logical-to-physical qubit layout, also known as qubit placement, involves deriving an injective map from the virtual qubits of the quantum circuit to the physical qubits of the quantum device (qubit_routing; sabre_swap). In other words, the qubit layout stage decides the initial position of logical qubits on the device. In contrast, qubit routing operates from an initial qubit layout and inserts SWAP operations to align a quantum circuit’s gates to a given qubit topology. Consider, for example, the state-of-the-art in qubit allocation: the SABRE heuristic search algorithm (sabre_swap). In SABRE, the distinction of the layout and routing stages motivates an iterative method for solving qubit allocation. Namely, an initial layout is used to perform qubit routing, which is used to generate a new initial layout.
Because qubit routing remains NP-hard (qubit_routing), another method of simplifying the problem is to assume qubit links are roughly equivalent, particularly with respect to average fidelity and two-qubit gate duration. This assumption can manifest implicitly when a given device topology is represented as an unweighted, undirected graph. For monolithic architectures, one could debate the marginal tradeoff between performance and solution quality; however, for modular architectures, the cross-chiplet connections are far worse compared to intra-chiplet connections (smith_2022_chiplets). Thus, we could expect the solutions generated from the simplified model to be worse than a more physically accurate model, for example, in circumstances where the shortest path in terms of the quantity or depth of inserted SWAPs diverges from the shortest path in terms of fidelity, gate duration, or some combination thereof.
2.2. Basis Translation
Basis translation converts the gates of an input quantum circuit into gates that are physically supported on hardware. Monolithic quantum architectures typically support a so-called universal basis gate set, where any unitary quantum operation can be constructed using solely the gates in the set (mike_and_ike). Moreover, it has been shown that basis translation can be performed efficiently given the target gate set is universal, and the quantum circuit contains a fixed, finite number of qubits (Solovay–Kitaev). Unfortunately, modular quantum architectures depend upon specialized links to connect otherwise-independent chiplets, and these links often don’t support universal operations (niu2023low). Thus, when algorithms like SABRE depend on the expectation of universality among qubit links, they may generate incompatible outputs when targeting modular architectures.
2.3. Optimization
Quantum circuit optimization seeks to prune extraneous operations by finding opportunities to combine, eliminate, and/or parallelize quantum gates (optimization1). Through commutation and cancellation identities, it is possible to search for optimization opportunities by iteratively transforming the quantum circuit without affecting the underlying operation (optimization1). However, in general, optimization is highly time- and resource-intensive, with the problem scaling exponentially with respect to number of qubits and/or circuit depth (qfactor). As a result, partitioning, the division of a quantum circuit into blocks of quantum gates, is a common tactic among optimization techniques (qfactor; qgo; bqskit). One major consideration in this process is demarcation, i.e., where and when to draw the line between blocks. Namely, one forfeits all optimization opportunities along boundaries of demarcation, so they must be carefully chosen to minimize negative impact on the output solution. From this perspective, modular architectures are naturally synergistic with partitioning, since cross-chiplet connections present a physically motivated boundary between operations.
3. A Strawman Chiplet Compiler
(a) A physical inter-chiplet link exists between halo qubits $A_{0}$ and $B_{0}$. A physical intra-chiplet link exists between $B_{0}$ and $B_{1}$.
(b) A logical CNOT gate between chiplets $A$ and $B$ which is not implementable on hardware (indicated in red).
(c) To prepare for an inter-chiplet SWAP, the logical qubit on $B_{0}$ must be moved to a nearest neighbor, $B_{1}$ in this case, with a SWAP operation.
(d) The inter-chiplet SWAP between chiplets $A$ and $B$ completes transformation of the CNOT gate to an intra-chiplet gate.
(e) Once the CNOT gate has been successfully placed, the inserted SWAPs must be undone to not affect future gates in the circuit.
Figure 1. Peephole correction for incompatible (non-SWAP) gates placed on inter-chiplet halo edges.
It was challenging to establish an appropriate baseline for state-of-the-art quantum circuit compilation on chiplet-based quantum devices. Not only are modular quantum architectures relatively nascent, but compilation schemes for monolithic devices are largely incompatible with chiplet devices, as was discussed in greater detail in Section 2.
Thus, we took it upon ourselves to design a baseline compilation scheme for chiplet architectures, against which we compare our more-optimized design presented in Section 4. This “strawman” compiler augments existing compiler designs with the minimal modifications necessary to generate valid circuit outputs for chiplet architectures. Specifically, we insert a “peephole” compiler pass immediately after qubit routing, which corrects for the placement of two-qubit gates on inter-chiplet links as shown in Figure 1. This algorithm is inefficient, since it inserts a total of 4 SWAP gates — 2 inter-chiplet and 2 intra-chiplet — for every inter-chiplet gate. Unfortunately, this degree of overhead is inevitable when the stock, unmodified compiler is ignorant of chiplet-specific device constraints.
4. The Design of SEQC
We incorporate the hierarchical elements of chiplet architectures into our compiler design by adopting a two-stage procedure. The first, stratification stage, represents a one-time overhead to elevate a quantum circuit to the chiplet level. The second, elaboration stage, represents a recurring cost to complete the compilation of a quantum circuit given the most up-to-date backend properties.
4.1. Stratification Stage
(a) A $3$-chiplet topology with $2$ qubits per chiplet ($6$ qubits in total).
(b) Mapping of a $6$-qubit quantum circuit to $3$ subcircuits. Gates represent arbitrary cross-subcircuit operations.
(c) Mapping of subcircuits to physical chiplets. Since chiplets $A$ and $C$ are not physically connected, certain gates (shown in red) necessitate inter-chiplet routing.
(d) Inter-chiplet routing is performed to address cross-chiplet gates that either don’t conform to the physical inter-chiplet topology or the basis gate set (aren’t SWAP gates).
Figure 2. Showcase of Inter-Chiplet Compilation in the Stratification Stage
In this stage, we map a quantum circuit to the inter-chiplet device topology, e.g., the structure in Figure 2(a), by stratifying it into smaller, interconnected subcircuits. This can be accomplished with the following intermediate steps, as shown in Figures 2(b)2(d). First, qubit-to-subcircuit mapping designates which logical qubits belong to which logical subcircuits. Then, as part of chiplet allocation, we define a mapping of logical subcircuits to physical chiplets on the target device and route inter-chiplet gates to conform with constraints of physical-chiplet connectivity.
4.1.1. Qubit-to-Subcircuit Mapping
To perform qubit-to-subcircuit mapping (Figure 2(b)), we leverage a form of simulated annealing. Simulated annealing is a form of stochasic optimization, notably used in VLSI algorithms for place-and-route (annealing1989). We use simulated annealing to find a qubit assignment that minimizes the number of connections between subcircuits. As our chiplet allocation algorithm acts conservatively—performing the minimal number of changes necessary to make a circuit valid—it will produce better results if given a circuit with minimal connectivity between subcircuits. For our cost function, we count the number of inter-subcircuit gates. To permute the solution, we swap two random qubits in different subcircuits. As different simulated annealing runs are independent, running the algorithm using different seeds is embarrassingly parallel. We exploit this parallelism by running many trials at once, each in a different process with a different seed, then choosing the best result across all trials and seeds.
4.1.2. Chiplet Allocation
To perform chiplet-level allocation, shown in Figures 2(c) and 2(d), we extend and modify the SABRE algorithm for initial mapping (sabre_swap) with application-specific layout and routing passes. For layout, we initialize each iterative allocation cycle with a random placement of logical subcircuits into physical chiplets. For routing, we developed a new heuristic for deciding between candidate SWAPs.
By default, SABRE attempts to place two-qubit gates on a valid edge in the graph of a device’s hardware topology, i.e, the physical distance between the qubits of a gate should be $1$. For inter-chiplet routing, we do not wish to factor intra-qubit topology into our distance calculation. Furthermore, we hope to move all gates to be within the same chiplet, i.e., the distance between qubits should be 00 instead of $1$. Thus, when selecting between potential SWAPs, we categorize and prioritize SWAPs in a three-tier approach:
  1. (1)
    Symbiotic SWAPs: A SWAP between two multi-chiplet gates is considered symbiotic if it reduces the distance of both gates, i.e., each gate benefits from the SWAP. In theory, we should expect symbiotic SWAPs to be the most efficient, and thus highest-priority, SWAP possible.
  2. (2)
    Commensalistic SWAPs: When a SWAP can benefit one multi-chiplet gate without harming any other, we consider it to be commensalistic. Such behavior occurs most commonly on idle qubits in the quantum circuit, although it can occur between two multi-chiplet gates.
  3. (3)
    Parasitic SWAPs: The remaining category of SWAPs to consider benefit one gate while harming another. Since this variety of SWAP is actively counterproductive towards a gate, it should be, and likely will be, avoided in most circumstances. However, given a particularly onerous circuit, it may be necessary in such rare instances. For example, clear one gate may free physical qubits to help route others.
With these changes, our algorithm is able to correctly and efficiently place and route circuits for any valid device topology.
4.2. Elaboration Stage
(a) A $3$-qubit chiplet with a line topology and connected to a neighboring chiplet by $q_{2}$.
(b) A $3$-qubit subcircuit with $2$ inter-chiplet SWAPs.
[…]
(d) A greedy policy takes the shortest path for routing inter-chiplet gates to their nearest halo qubit.
(e) Qubit routing for the remaining intra-chiplet gates is performed as normal, except with restrictions to ensure inter-chiplet SWAPs are kept static.
(f) The remaining compilation states (basis translation and optimization) are not affected by the additional constraints imposed by inter-chiplet gates.
Figure 3. Showcase of Intra-Chiplet Compilation in the Elaboration Stage.
Following the stratification stage, each chiplet can be compiled almost completely independently from one another. More precisely, the compilation process within each chiplet is largely identical to that of a monolithic quantum architecture. The only exception is the presence of inter-chiplet SWAPs, which must be lowered from the chiplet level to the qubit level. Once this assignment is made, these SWAPs must remain fixed and immutable to subsequent compilation stages. Despite this overhead, restricting the task of compilation to a narrower domain of qubits facilitates both parallelization (as each subcircuit can be processed in parallel) and problem size reduction (as there are less qubits to consider) of the most computationally intensive compilation stages (routing and optimization). To accomplish these tasks, we divide the elaboration stage into two substages: qubit allocation and parallel basis translation and optimization. Figure 3 illustrates this with an example of elaboration for a single chiplet in Figure 3(a) and its corresponding subcircuit in Figure 3(b).
4.2.1. Qubit Allocation
Like other qubit allocation schemes, we incorporate layout (see Figure 3(c)) and routing (see Figure 3(e)) stages. The presence of inter-chiplet gates necessitates several key changes to the SABRE (sabre_swap) qubit allocation algorithm to achieve correctness and/or maximum parallelism. First, the placement of inter-chiplet gates is a non-parallelizable task (due to cross-chiplet interactions) that strongly influences the layout and routing passes, which could otherwise be fully parallelized across chiplets. To account for this, we introduce a serial stage (shown in Figure 3(d)) between the parallelized layout and routing passes that greedily places inter-chiplet gates on the nearest valid edge between halo qubits. A greedy policy was chosen because inter-chiplet gates have more restrictions on valid placements compared to intra-chiplet gates, while also experiencing substantially higher error rates (smith_2022_chiplets). Once the positions of the inter-chiplet gates are settled, they must remain fixed in place, imposing an additional constraint on qubit routing. To promote this behavior in SABRE, we extended its cost heuristic to weigh the device topology by fidelity. This causes higher error links to assume higher costs, discouraging SABRE from moving the high-error inter-chiplet gates. As a result, SEQC achieves better solutions in the common case. However, there is nothing to prevent SABRE from incorrectly modifying any inter-chiplet SWAPs. Thus, we need to impose hard restrictions to ensure correctness. We establish two disjoint gates sets. One gate set operates strictly within a chiplet. This is the traditional gate set for a monolithic processor. The other gate set is the only one permitted to operate across chiplets, and specifically only implements inter-chiplet SWAPs, i.e., SWAP gates between two halo qubits. By restricting each gate set to its own disjoint part of the overall device topology, we ensure correctness across all cases.
4.2.2. Parallel Translation and Optimization
The remaining compilation stages of basis translation and circuit optimization can be easily parallelized at the chiplet granularity, as shown in Figure 3(f). Basis translation, in particular, is embarrassingly parallel, as every gate can be translated independently from every other gate. Meanwhile, as discussed in Section 2, optimization algorithms already employ partitioning and parallelization to create tractable subproblems from intractably large circuits, facilitating a smaller problem space, lower memory usage, and faster performance. In modular quantum architectures, we physically motivate partitions through the natural chiplet boundaries, enabling us to reap the same resource and performance benefits of partitioning without compromising on solution quality.
5. Methodology
To experimentally validate our compiler for modular quantum architectures, we ran a benchmark suite on a wide spectrum of circuit sizes for both SEQC and the strawman baseline. Both SEQC and the baseline compiler were implemented in the Qiskit SDK (qiskit2024) v1.2.41.2.41.2.41.2.4 and Python $3.12.6$. Elaboration time results were taken on a machine with a $128$-core Ampere Altra Max ARM64 processor and $256$ GB of DDR4 RAM. When given a parallelizable workload, SEQC attempts to allocate its tasks evenly among the available physical cores, rounded to the neared whole number. Finally, the simulated annealing technique used in the qubit-to-subcircuit stage uses a starting temperature $T_{0}=200$K, a rate of change of $0.005T$ each round, $\frac{T}{50}$ permutations per round, and a total of $5$ trials per core.
5.1. Metrics
We compare SEQC to the strawman baseline over a wide range of common and/or relevant metrics, discussed below. We aggregate results in a manner similar to the SPEC family of benchmarks (spec2017), by taking the geometric mean of the relative performance ratio against a baseline technique across a benchmark suite of quantum circuits. The resultant score is then derived for each of the evaluated metrics on multiple backends of varying numbers of chiplets.
Fidelity. The top priority for quantum circuit execution is to maximize fidelity. Without a guarantee of correctness, the outputs of the quantum circuit are useless, and any speedup or quantum advantage is rendered moot. Fidelity is impacted by the decoherence constraints of individual qubits, as well as the errors on a gate-by-gate basis (mike_and_ike). Due to simulation and hardware limitations, we utilize the estimated success probability (ESP) metric (esp) to approximate fidelity. Specifically, we track the per-qubit ESP and report an aggregate ESP for each benchmark from the geometric mean of the accumulated measurements.
Non-recurring stratification time. This represents the time taken in the stratification stage, which will be present for our compiler and absent for the baseline. The stratification time represents a one-time overhead to prepare a quantum circuit for execution on a modular architecture. This result can be reused for future executions on that device or any device with an equivalent or more expansive topology.
Recurring elaboration time. The error characteristics of a quantum device can change on a hour-by-hour basis, or even a run-to-run basis (Tannu_2019_ErrorVariability; Murali_2019_NoiseMapping; dasgupta2021stability). Thus, to generate the highest-quality circuits with the most up-to-date characteristics possible, a given quantum circuit should be recompiled for every execution. Since the qubit and chiplet topologies remain static per-device, it is possible under the SEQC framework to reuse the results from the stratification stage, and only incur overhead from repeating the elaboration stage.
Solve time. This metric is relevant only to the baseline strawman chiplet compiler, and represents the time it takes to compile a quantum circuit. The baseline compiler solve time is a recurring cost, i.e., this compilation step has to be performed every time a circuit needs to execute on a quantum machine.
Estimated execution time. The per-shot execution time of a quantum circuit is determined by the critical path of gate operations weighted by gate duration. In addition to performance concerns, execution time indirectly affects fidelity on NISQ-era devices due to decoherence. In other words, it is imperative that quantum circuits execute within the decoherence time to ensure high-fidelity outcomes.
Number of inter-chiplet gates. The inter-chiplet gates of a chiplet architecture contribute the most to error and execution time relative to other gates (at least by a factor of 4×4$\times$4 × (smith_2022_chiplets)). As a result, the number of inter-chiplet gates should represent a strong proxy metric for the quality of a compiler solution. Accordingly, minimizing the number of inter-chiplet gates should result in higher fidelity, faster-executing quantum circuits.
Circuit depth. Defined as the unweighted critical path of a quantum circuit (mike_and_ike), the circuit depth is often employed as a backend-agnostic metric for execution time that can be measured statically. Notably, it is used in SABRE (sabre_swap) as part of its cost heuristic.
Gate count. Similar to circuit depth, the total number of gates in a circuit can also be used as a static, backend-agnostic metric for certain properties of the aforementioned circuit. For instance, it may provide an indication of the amount of accumulated error, or, like circuit depth, the execution time. SABRE (sabre_swap) uses gate count to decide between randomly seeded allocation trials.
5.2. Benchmarks
For our benchmark suite, we leverage a subset of circuits in the Supermarq (supermarq) suite of quantum circuits. The circuits we select are chosen to be representative of practical, real-world applications for quantum computing, while also being feasible to compile for a large number of qubits. In particular, our benchmark suite comprises the following quantum circuits.
[…]
VQE. The variational quantum eigensolver (VQE) (vqe) is another hybrid quantum-classical variational algorithm that is often used in chemistry applications. In our experiments, we choose a $2$-layer ansatz whose parameters are once again randomly generated.
Hamiltonian Simulation. Quantum devices are naturally well-suited to simulating Hamiltonians. The Supermarq benchmark selects the 1D Transverse Field Ising Model (TFIM) system to model, which is relevant for phase transitions in magnetic materials.
5.3. Simulated Quantum Device Specifications
We perform our experiments on mock backends that are generated to conform to the chiplet architecture proposed in Smith et al. (smith_2022_chiplets). The hardware specifications of the qubits and intra-chiplet connections are sourced from Acharya et al. (backend_specs). Table 1 details the full specifications. To generate the expected fidelity of the inter-chiplet SWAPs, we perform a simulated random benchmarking trial and calculate the error rate of a SWAP gate with these backend specifications. Following Smith et al. (smith_2022_chiplets), we model inter-chiplet SWAPs on this backend with 4×4$\times$4 × the intra-chiplet error rate. Similarly, we approximate the duration of the inter-chiplet SWAP gate as 4×4$\times$4 × the duration of an intra-chiplet SWAP.
T1 (backend_specs) $20\times 10^{-6}$ seconds
T2 (backend_specs) $30\times 10^{-6}$ seconds
Frequency (backend_specs) $6\times 10^{9}$ Hz
(a) Qubit Properties.
Instruction Duration (ns) Error
(backend_specs) $25$ $0.109\%$
SX (backend_specs) $25$ $0.109\%$
$\text{R}_{\text{z}}(\phi)$ (backend_specs) 00 $0.00\%$
CZ (intra-chiplet) (backend_specs) $34$ $0.605\%$
SWAP (inter-chiplet) $702.4$ $10.23\%$
Reset (backend_specs) $500$ $0.186\%$
Measure (backend_specs) $500$ $0.196\%$
(b) Instruction Properties.
Table 1. Specifications of Simulated Chiplet Backend.
The topology of the simulated device is composed of a grid lattice of symmetric chiplet modules, where each chiplet features a heavy-hexagon qubit topology (heavy_hex), as shown in Figure 4. The intra-chiplet topology is constructed such that the heavy-hex structure is preserved for the global device topology (ignoring chiplet division). The tessellation requirements of each chiplet restrict valid chiplet sizes to multiples of $10$; however, for our experiments, we limit the size of each chiplet to the minimum of $10$ qubits due to hardware limitations.
Note that for each valid chiplet count, there may exist multiple valid chiplet topologies. In this work, we generate our backends to use the “most-square” topology for a given chiplet count. For example, a $12$-chiplet machine would be generated with a grid of $3\times 4$ chiplets.
Figure 4. Construction of a grid inter-chiplet lattice from a heavy-hexagon qubit lattice, where the heavy-hex lattice is maintained for each chiplet.
6. Experimental Results
Figures 57 present measurements of the recurring compilation time (i.e., solve time) of the baseline strawman chiplet compiler, and SEQC’s two compilation stages: the non-recurring stratification stage and the recurring elaboration stage.
Figure 5. SEQC stratification time normalized to the 20-qubit (i.e., 2-chiplet) trial.
The average stratification time of SEQC across the benchmark suite obeys a quadratic trajectory ($R^{2}=0.9971$) with respect to the circuit size and number of chiplets, as shown in Figure 5. We note that the compilation time of the baseline, presented in Figure 6, also scales quadratically ($R^{2}=0.9964$). Although stratification time represents a one-time cost, the quadratic scaling of stratification points toward the practicality of SEQC. The stratification time depends heavily on the complexity of the input circuit. Simple circuits, such as GHZ or HamiltonianSimulation, exhibit a very gradual increase in stratification time as circuit size increases, while others, such as BitCode and PhaseCode, are much steeper. This observation holds true for the solve time of the baseline compiler as well. Collectively, these results indicate that the computational complexity of the baseline compiler and the stratification stage of SEQC are affected similarly by the same quantum circuit characteristics.
Figure 6. Baseline compilation time normalized to the 20-qubit (i.e., 2-chiplet) trial.
Figure 7. Relative elaboration time of SEQC normalized to the strawman baseline for devices with 10-qubit chiplets (lower is better).
From Figure 7, we observe an improved SEQC elaboration time compared to the baseline solve time, leading to a roughly 24×2-4$\times$2 - 4 × speedup, up until $800$-qubit (i.e., $80$-chiplet) trials. The observed speedup reflects the expected performance improvement that our stratify-elaborate technique should garner from parallelization and problem-space reduction. For $81$-chiplet experiments and above, our results begin to reflect a far more marginal speedup compared to baseline, reflecting the limitations of our evaluation hardware. Namely, as the number of chiplets begins to approach the number of physical cores and hardware threads on our hardware, SEQC experiences difficulties in evenly distributing its tasks. For example, under-allocating tasks causes poor core utilization, while over-allocation leads to process thrashing, either way resulting in poorer elaboration time. However, in all cases, SEQC outperforms the baseline on average.
On a circuit-by-circuit basis, we notice higher relative performance with parallel workloads (like BitCode and PhaseCode) compared to serial workloads (GHZ and HamiltonianSimulation). Notably, above the $81$-chiplet threshold where the classical machine we use for compilation no longer suffices, elaboration time exceeds the baseline solve time for the serial workloads. This is because serial circuits are fundamentally trivial to solve, and thus benefit less from the parallelism offered by SEQC. We verified experimentally that in these cases, SEQC spends almost all of its time on thread startup and teardown, rather than useful work. Moreover, the raw compilation time is trivially small in these cases, owning to the simplicity of the circuits, thus relative performance differences have little practical significance.
Figure 8. Relative estimated fidelity (ESP) of SEQC normalized to the strawman baseline for devices with 10-qubit chiplets (higher is better).
Figure 8 presents the relative estimated fidelity of SEQC compared to the baseline with respect to the number of qubits per circuit. Notably, as the number of chiplets / the circuit size increases, SEQC improves its relative fidelity against the baseline. We begin to see higher fidelity circuits in the geometric mean once the circuit size exceeds $36$ chiplets, up to a maximum of $36\%$ higher fidelity. All benchmarks reflect higher relative performance for the $50$-chiplet trials and above.
Figure 9. Estimated per-shot execution time of circuits produced by SEQC normalized to circuits produced by the strawman baseline for devices with 10-qubit chiplets (lower is better).
Figure 9 presents the estimated per-shot execution time of the circuits produces by SEQC compared to the circuits produced by the baseline compiler, as a function of the number of qubits per circuit. As explained above, we approximate the duration of an inter-chiplet SWAP as 4×4$\times$4 × the duration of an intra-chiplet SWAP (smith_2022_chiplets). Under this model, we observe execution time benefits up to a 1.92×1.92$\times$1.92 × improvement over baseline at $750$ qubits ($75$ chiplets). Not only are improvements in execution time desirable for quantum circuit throughput, they could also result in lower decoherence and thus improved circuit fidelity.
Figure 10. Relative number of inter-chiplet gates for SEQC normalized to the strawman baseline for devices with 10-qubit chiplets (lower is better).
Figure 10 shows the number of inter-chiplet swaps in the circuits generated by SEQC relative to the circuits generated by the baseline, with respect to the number of qubits per circuit. For essentially all tested circuits, SEQC achieves a significant reduction in the number of inter-chiplet swap gates, appearing to asymptotically converge on a roughly 4×4$\times$4 × reduction, with the greatest reduction of 4.6×4.6$\times$4.6 × at $800$ qubits / $80$ chiplets. This aligns with our model of inter- vs intra-chiplet gates, where inter-chiplet SWAPs are assumed to have 4×4$\times$4 × the error (smith_2022_chiplets) and 4×4$\times$4 × the gate duration.
Figure 11. Relative circuit depth of SEQC ormalized to the strawman baseline for devices with 10-qubit chiplets (lower is better).
Figure 12. Relative gate count of SEQC normalized to the strawman baseline for devices with 10-qubit chiplets (lower is better).
Relative circuit depth and gate count, shown in Figures 11 and 12, respectively, share similar behavior and analysis. At small circuit sizes, SEQC has higher circuit depth and gate count than the baseline. The gap shrinks as the circuit size increases, although SEQC always tends to perform worse on these metrics on average. This behavior is to be expected, since the baseline directly optimizes for circuit depth and gate count, while SEQC prioritizes other metrics, e.g., reducing inter-chiplet gates.
6.1. Discussion
The discrepancy between the circuit depth and gate count results compared to the execution time and fidelity results, respectively, are indicative of the differences between SEQC and the contemporary quantum compilers. Circuit depth and gate count represent the heuristics used in the baseline, especially SABRE (sabre_swap), and they are fundamentally unweighted metrics. In other words, circuit depth and gate count only consider the number of gates (in the critical path and in total) irrespective of each one’s unique underlying properties. Meanwhile, SEQC is able to more directly estimate execution time and fidelity, the practically relevant metrics, by constructing weighted heuristics.
With the context of our static metrics (number of inter-chiplet gates, circuit depth, and gate count), it is possible to justify the improvements in relative fidelity and relative per-shot execution time as the circuit size increases. Namely, as expected, the unweighted heuristics in SABRE tend to under-represent the error and gate delay contribution of inter-chiplet gates. This effect gets exacerbated for wider or deeper circuits, which increase the number of opportunities for cross-chiplet interactions.
Overall, we observe that SEQC yields circuits with more desirable characteristics when executing on modular quantum architectures, while at the same time it achieves faster compilation times. Also, as the number of qubits increase, the performance improvement of SEQC widens compared to the baseline, especially for the most important metrics of fidelity, compiled circuit execution time, and number of inter-chiplet gates. These trends hold promise in allowing for faster and better compilation for future large-scale modular quantum devices than we can achieve with today’s frameworks.
7. Related Work
Even though quantum computing is still nascent, there is a significant body of work on techniques for compiling quantum circuits for NISQ processors. Here, we focus on the work that best relates to our contributions.
MECH (mech) sacrifices some ancillary qubit resources to build an ”adjustable multi-entry communication highway” through software. By entangling ancillary qubits and distributing them throughout the chiplet architecture’s topology, the work virtualizes those qubits into fast network-on-chip routing channels without hardware modifications. Future work to extend SEQC with MECH’s methodology may enable combining MECH’s techniques with SEQC’s allocation procedures to gain the benefits of each.
Noting the scaling challenges and fidelity issues associated with large, monolithic quantum machines, CutQC (cutqc) establishes a hybrid computing paradigm which enables smaller-scale quantum computers to evaluate quantum circuits they would have otherwise had too few qubits to execute. To establish said paradigm, the work introduces a novel circuit-cutting technique for partitioning larger quantum circuits into small subcircuits. Classical post-processing is performed on the results of the subcircuit evaluations, merging them to find the result that would be generated by the uncut circuit. Circuit cutting as performed in CutQC served as an inspiration for SEQC’s stratification step, but unlike CutQC, SEQC does not perform any classical “stitching” of the results afterwards. It is interesting to note that our first version of SEQC was formulating the problem using constrained optimization as well, but the slowdown of the stratification step was unacceptable and impractical for large quantum circuits, and thus was abandoned in favor of the heuristics we currently employ in our compiler.
Bandic et al.(bandic2024profiling) also note the benefits of modularity. Their study is motivated by application-specific optimization – they seek techniques that allow a quantum circuit’s structure to influence hardware architecture. In their work, the relationship between circuit structure and mapping efficiency is defined, and this information optimizes the architecture of single- and multi-core quantum processors.
Piveteau and Sutter (circuit_knitting) investigate the benefits of using quasiprobability simulation to allow subcircuits cut from larger quantum circuits to perform their quantum operations independently, simulating halo qubit behavior where the subcircuits connect. When augmented with classical communication between circuits to reduce the simulation overhead incurred by sampling, they demonstrate that resulting circuits reach fixed accuracy with less simulation time. Unlike SEQC that targets quantum compilation, this technique has been developed to partition large quantum circuits into subcircuits that fit on smaller devices, at the cost of a simulation overhead to classically combine them.
Lin et al. (let_every_qubit) recognize that each 2-qubit gate in a quantum circuit has a dynamic behavioral profile which depends on manufacturing and environmental factors, and present a compilation technique which uses each 2-qubit set’s calibration data to determine potentially nonstandard 2Q gate operations which are much more ideal for that set to perform than their standard counterparts—essentially, choosing a more favorable basis. By characterizing these nonstandard gates, using their unitaries to synthesize standard gates, and essentially allowing 2Q gates to perform standard operations along bases more compatible with their respective physical states, the technique improves fidelity. This technique is orthogonal to SEQC and could work synergistically with it.
By mapping circuit partitioning to a minimum cut problem, Bandic et al. (bandic2023mapping) use the QUBO model with quantum solvers to generate a set of cuts intended to minimize swaps between circuit partitions. QUBO employs quantum optimization algorithms on quantum annealers to accelerate the solution process and remain practical, as the search space for transmission qubits is $2^{n}$. SEQC, in contrast, does not require acceleration through quantum means, but at the cost of not guaranteeing an optimal solution (it is rather a “best-effort” approach that works well in practice).
Through discussion of scalability issues impeding the development of larger monolithic quantum circuits, Ovide et al. (ovide2023mapping) motivate the necessity of mapping quantum circuits to multi-core architectures and present a commonly agreed-upon likely computing paradigm which leverages both quantum and classical methods to foster inter-chip communication, circumventing the scalability challenges inherent to monolithic quantum processors. While this work advocates for a two-level, hierarchical approach for quantum circuit mapping, it leaves the actual problem of qubit placement and then routing unsolved. SEQC provides an algorithm and implementation, addressing both of these challenges.
Using iterative optimization via the Hungarian algorithm, Hungarian Qubit Assignment (HQA) (escofet2023hungarian) seeks to optimize the qubit-to-core mapping problem. Because the Hungarian algorithm is here extended with the ability to adjust a qubit’s cost matrix according to behavior multiple timeslices ahead, the model can leverage lookahead to optimize further the number of inter-core swaps. However, HQA operates under the assumption that all inter-chiplet communications have the same cost, meaning chiplets must be fully connected. SEQC, on the other hand, is able to work with any arrangement of chiplets and inter-chiplet links (as long as there is a path between any two chiplets), taking into account the distance between pairs of chiplets to generate improved circuits.
Iterating on their previous work on Hungarian Qubit Assignment, Escofet et al. (escofet2024revisiting) develop a characterization technique for arbitrary quantum circuits which yields theoretical upper and lower optimization bounds. This is significant, since it enables optimality assessment for solutions to the qubit mapping problem.
Pastor et al. (pastor2024circuit) leverage deep reinforcement learning to partition quantum circuits for execution across multiple processing cores. Using a DRL model enhanced with policy-masking mechanisms and a lookahead-capable observation array, the work finds valid qubit-circuit mappings for each of timeslice while prioritizing the reduction of inter-core swaps. Notably, it only considers inter-core swaps for optimization, deeming intra-core single-and-multiple qubit operations so inexpensive that their optimization should serve as fodder for inter-core swap reductions. This is sensible, but leaves optimization potential on the table, since it only seeks to generate valid mappings within each timeslice. Additionally, similar to HQA, this technique treats all inter-chiplet communications equally, further missing optimization opportunities with non-fully connected chiplet topologies. SEQC, in contract, is cognizant of the differing costs of moving information between each pair of chiplets, and optimizes accordingly.
To solve the qubit allocation (mapping) problem, Liu et al. (liu_2023_tackling) develop an algorithm that combines partitioning, permutation-aware synthesis, and a modified version of SABRE. In a bottom-up approach, this technique partitions a logical circuit into blocks, performs permutation-aware synthesis on each block in parallel (solving for all possible qubit topologies), and finally uses an augmented SABRE algorithm to perform initial layout and routing between blocks. In contrast, SEQC takes a top-down approach, where SWAP permutations between chiplets are settled prior to any compilation tasks performed within each block. As a consequence, we only need to generate solutions for one qubit layout per block, rather than synthesizing all possible block permutations.
Parallelizing SEQC’s recurrent elaboration step was inspired by traditional classical compilation techniques, which has a decades-long history and is now commonplace (Mickunas_1978_ParallelCompilation; Gross_1989_ParallelCompilation). Besides traditional compilation, though, we can also draw parallels between our work and prior research on data movement in the classical space. Data movement is widely understood to be the primary performance bottleneck in the classical architecture space. In the last two decades, work has proliferated on dataflow and spatial architectures, much of which focuses on mapping problems. Consider neural network architectures, for example, which seek to map their physical topologies as closely as possible to problem topologies. Eyeriss (Eyeriss_2016_Chen) approaches the dataflow mapping problem by proposing a Row-Stationary architecture, and then optimizing the mapping of problem to processing element such that elements which are spatially local in the problem are spatially local in hardware during execution. This is not unlike qubit swap optimization, and in fact is performed for many of the same reasons—like inter-chiplet swaps, data movements between more distant classical programming elements in a dataflow architecture are simply more expensive, and minimizing their frequency yields performance improvements.
8. Conclusion
Physical, technological and economic considerations make the prospect of a single “monolithic” large-scale quantum processor very expensive or entirely unrealistic by current standards. These constraints force future quantum systems to adopt a physically distributed architecture, based on modular quantum chiplets. As such, modular quantum processors already dominate the roadmaps of several major quantum companies.
However, the complexity of chiplet-based quantum devices, coupled with their growing size, presents an imminent scalability challenge for quantum compilation. Existing qubit allocation methods are often unable to contend with inter-chiplet links, which don’t necessary support a universal basis gate set, and existing methods of logical-to-physical qubit placement, routing, unitary synthesis, and quantum circuit optimization, are typically not designed for qubit links of wildly varying levels of duration or fidelity.
In this work, we first modify stock contemporary compilers to become aware of hardware modularity, thereby allowing them to correctly compile circuits for modular architectures with limited cross-chiplet gate support. We then propose SEQC, a Stratify-Elaborate Quantum Compiler for modular architectures. SEQC performs compilation in two stages, with the first stage (stratification, or chiplet splitting) performed only once for a given architecture, and the second stage (elaboration, or chiplet compilation) performed in parallel for each chiplet. Unlike contemporary quantum compilers that have to compile a quantum circuit anew before each execution, only the elaboration stage of SEQC needs to be performed recurrently before each execution. As its computational complexity is a function of the quantum chiplet size, not the entire quantum processor size, and chiplet sizes grow slowly, SEQC’s recurrent compilation time is largely unaffected by the relentless growth of qubit counts in future quantum processors.
We design and implement several novel methods for qubit placement, qubit routing, and circuit optimization in SEQC, with both hardware modularity and compilation parallelization in mind. Owning to these techniques and its parallelization strategy, SEQC compiles circuits that exhibit up to a 36% increase in circuit fidelity and up to 1.92×1.92$\times$1.92 × lower execution time, while consistently achieving 24×2-4$\times$2 - 4 × faster compilation time over a chiplet-aware Qiskit baseline.
[…]

### v2

@@
 Contemporary compilation methods are not well-suited to chiplet architectures — in particular, existing qubit allocation methods are often unable to contend with inter-chiplet links, which don’t necessary support a universal basis gate set.
-Furthermore, existing methods of logical-to-physical qubit placement, swap insertion (routing), unitary synthesis, and/or optimization are typically not designed for qubit links of wildly varying levels of duration or fidelity.
+Furthermore, existing methods of logical-to-physical qubit placement, swap insertion (routing), unitary synthesis, and/or optimization are typically not designed for qubit links of significantly varying levels of duration or fidelity.
 In this work, we propose SEQC, a complete and parallelized compilation pipeline optimized for chiplet-based quantum computers, including several novel methods for qubit placement, qubit routing, and circuit optimization.
 SEQC attains up to a 36% increase in circuit fidelity, accompanied by execution time improvements of up to 1.92×1.92$\times$1.92 ×.
-Additionally, owning to its ability to parallelize compilation, SEQC achieves consistent solve time improvements of 24×2-4$\times$2 - 4 × over a chiplet-aware Qiskit baseline.
+Additionally, owing to its ability to parallelize compilation, SEQC achieves consistent solve time improvements of 24×2-4$\times$2 - 4 × over a chiplet-aware Qiskit baseline.
 Quantum computing, quantum compilation, modular architectures.
@@
 While improvements in superconducting quantum hardware have led to the debut of processors with $1000+$ qubits (ibm_condor), practical implementations of quantum computation will require millions of physical qubits (paler2019really).
-The number of qubits required by a quantum algorithm, the depth of the algorithm (i.e., execution time, or number of gates on the critical path), the auxiliary and syndrome qubits required to reach sufficiently low error rates, and the size of matter qubits when all control hardware is included, all suggest that many more qubits are required than are likely to fit in a single die (vanMeter_2010_DQC).
+The number of qubits required by a quantum algorithm, the depth of the algorithm (i.e., execution time or gate count on the critical path), the auxiliary and syndrome qubits required to reach sufficiently low error rates, and the size of qubits with all control hardware, all suggest that many more qubits are required than are likely to fit in a single die (vanMeter_2010_DQC).
 The overhead for error correction, cooling, dilution, control systems, I/O lines, challenges associated with verification and adequate chip thermalization, and realistic resources such as lossy waveguides, limited qubit fabrication yields as qubit capacity grows (smith_2022_chiplets), and finite chip sizes, make the prospect of a single “monolithic” quantum processor very expensive or entirely unrealistic by current standards.
@@
 Coupled with the hardness of synthesizing unitaries (which is exponential in the number of qubits) and the need for quick and frequent compilation, modern quantum compilers have no option other than to rely on heuristics to perform the task.
-These heuristics result in compilation complexity of $O(n^{2})$ for $n$ total qubits in a quantum processor.
+These heuristics result in a compilation complexity that we experimentally observed to be $O(n^{2})$ for $n$ total qubits in a quantum processor, as shown in Figure 8.
 Hardware modularity adds significant complexity to this already hard task, as inter-chiplet links are typically inferior compared to intra-chiplet ones (laracuente2022modeling; smith_2022_chiplets), connectivity across chiplets is often limited (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers), and not all basis gates are necessarily supported across chiplets (niu2023low).
@@
 2. Background and Motivation
+(a) A chiplet quantum architecture
+(b) Input circuit from SABRE
+(c) Insertion of perfect-shuffle
+(d) Insertion of return SWAPs
+Figure 1. Peephole correction for incompatible (non-SWAP) gates placed on inter-chiplet edges. Gates that are unimplementable on hardware are shown in red.
 To execute a program on a target device, the logical program representation must be transformed such that it conforms to the idiosyncratic physical constraints for that particular device.
@@
 Moreover, it has been shown that basis translation can be performed efficiently given the target gate set is universal, and the quantum circuit contains a fixed, finite number of qubits (Solovay–Kitaev).
-Unfortunately, modular quantum architectures depend upon specialized links to connect otherwise-independent chiplets, and these links often don’t support universal operations (niu2023low).
-Thus, when algorithms like SABRE depend on the expectation of universality among qubit links, they may generate incompatible outputs when targeting modular architectures.
+Unfortunately, modular quantum architectures depend upon specialized links to connect otherwise-independent chiplets.
+In the general case, the cross-chiplet links can only be used for data movement (such as SWAPs) and not computation, meaning these links may not support universal operations (niu2023low).
+Thus, compilers that depend on the expectation of universality among qubit links (qiskit_basis_translation_errors) and do not expect these heterogenous, non-universal basis gate sets can generate incompatible outputs when targeting modular architectures.
 2.3. Optimization
@@
 3. A Strawman Chiplet Compiler
-(a) A physical inter-chiplet link exists between halo qubits $A_{0}$ and $B_{0}$. A physical intra-chiplet link exists between $B_{0}$ and $B_{1}$.
-(b) A logical CNOT gate between chiplets $A$ and $B$ which is not implementable on hardware (indicated in red).
-(c) To prepare for an inter-chiplet SWAP, the logical qubit on $B_{0}$ must be moved to a nearest neighbor, $B_{1}$ in this case, with a SWAP operation.
-(d) The inter-chiplet SWAP between chiplets $A$ and $B$ completes transformation of the CNOT gate to an intra-chiplet gate.
-(e) Once the CNOT gate has been successfully placed, the inserted SWAPs must be undone to not affect future gates in the circuit.
-Figure 1. Peephole correction for incompatible (non-SWAP) gates placed on inter-chiplet halo edges.
-It was challenging to establish an appropriate baseline for state-of-the-art quantum circuit compilation on chiplet-based quantum devices.
-Not only are modular quantum architectures relatively nascent, but compilation schemes for monolithic devices are largely incompatible with chiplet devices, as was discussed in greater detail in Section 2.
-Thus, we took it upon ourselves to design a baseline compilation scheme for chiplet architectures, against which we compare our more-optimized design presented in Section 4.
-This “strawman” compiler augments existing compiler designs with the minimal modifications necessary to generate valid circuit outputs for chiplet architectures.
-Specifically, we insert a “peephole” compiler pass immediately after qubit routing, which corrects for the placement of two-qubit gates on inter-chiplet links as shown in Figure 1.
-This algorithm is inefficient, since it inserts a total of 4 SWAP gates — 2 inter-chiplet and 2 intra-chiplet — for every inter-chiplet gate.
-Unfortunately, this degree of overhead is inevitable when the stock, unmodified compiler is ignorant of chiplet-specific device constraints.
+Figure 2. High-Level Diagram of SEQC. Elements in red reflect modifications from existing compiler designs.
+As discussed previously in Section 2, our model of modular quantum architectures includes the more general case where inter-chiplet links are can be non-universal.
+This is representative of hardware manufacturing constraints (niu2023low), where only a restricted set of operations, e.g., state transfers or SWAPs, may be supported over long-distance couplers (inter-chiplet links).
+However, many compiler designs (including designs for modular architectures, see Section 7) experience errors or crashes when presented with non-universal, heterogenous, basis gate sets.
+For example, the Qiskit compiler (qiskit2024) is explicit in its compatibility for only universal basis gate sets (qiskit_basis_translation_errors).
+While it is possible to create a Qiskit Backend that supports heterogeneous basis gate sets, the compiler itself does not take them into account.
+Specifically, the Qiskit routing stage (zou2024lightsabrelightweightenhancedsabre) (an implementation of SABRE) will freely place gates on incompatible physical links, and there are no methods in the compiler to prevent or fix consequent errors.
+For the purposes of this work, however, it was critical to establish a baseline representative of the current state of quantum compilation, against which we can compare our more-optimized design (see Section 4).
+Thus, we created a “Strawman” baseline compiler, built on top of the Qiskit compiler (qiskit2024), with the minimal modifications necessary to generate valid outputs for chiplet architectures.
+To accomplish this, we leverage a “peephole” compilation pass, which is commonly used in classical compilation to perform optimization (dragon_book).
+Peephole passes operate within a nominal peephole window and substitute its contents with a logically-equivalent replacement.
+Notable examples in the context of quantum compilation include basis translation and elimination of self-adjoint gates.
+This design strategy facilitates modular compilation, not only in the sense that it can be easily parallelized, but also because peephole passes can be added between other compilation passes without breaking compatibility.
+These qualities make a peephole pass well-suited to the Strawman baseline, since we do not want to make extensive modifications to any of Qiskit’s built-in compilation stages or passes, nor introduce a significant performance overhead.
+The peephole pass used in the Strawman compiler is inserted between the routing and translation compilation stages in Qiskit.
+As shown in Figure 1, it operates on the circuit generated by SABRE (zou2024lightsabrelightweightenhancedsabre), where it detects when non-SWAP gates are placed on non-universal inter-chiplet links and employs a cyclic qubit rotation, i.e., a perfect shuffle permutation, to move the aforementioned gates to nearby intra-chiplet links ( see Figure 1(c)).
+Overall, for every inter-chiplet gate in the original circuit generated by SABRE, the peephole pass inserts 4 additional SWAP gates—2 inter-chiplet and 2 intra-chiplet.
+Ultimately, this degree of overhead is inevitable when the stock, unmodified compiler is ignorant of chiplet-specific device constraints.
+To emphasize the rigor of the Strawman, it is worth discussing the designs that were ultimately discarded.
+For example, consider a hypothetical alternative design that skips the final peephole step shown in Figure 1(d).
+This would appear to save $2$ SWAPs, thus halving the cost per inter-chiplet gate.
+However, the hypothetical pass would also no longer be a true peephole, since it introduces a net qubit permutation that propagates to downstream gates in the circuit.
+From the perspective of modular compilation, this results in serial dependencies that prevent parallel execution of the pass.
+From a routing perspective, the resultant qubit permutations disrupt the solution generated by SABRE, leading to destructive routing errors that can only be resolved by an entirely new routing pass.
+Overall, such a hypothetical pass results in substantially worse output circuits compared to the peephole method, which we confirmed through testing.
+Other potential improvements on the peephole design that avoid the placement of return SWAPs would involve extensive modifications to Qiskit’s compilation stages or its underlying passes.
+In Section 4, we will explore such a modification of SABRE that accounts for the only-the-fly changes in qubit layout associated with inter-chiplet routing as an element of our design, in addition to numerous other components to optimize compilation for modular quantum architectures.
 4. The Design of SEQC
-We incorporate the hierarchical elements of chiplet architectures into our compiler design by adopting a two-stage procedure.
+We incorporate the hierarchical elements of chiplet architectures into our compiler design by adopting a two-stage procedure, as shown in Figure 2.
 The first, stratification stage, represents a one-time overhead to elevate a quantum circuit to the chiplet level.
@@
 4.1. Stratification Stage
-(a) A $3$-chiplet topology with $2$ qubits per chiplet ($6$ qubits in total).
-(b) Mapping of a $6$-qubit quantum circuit to $3$ subcircuits. Gates represent arbitrary cross-subcircuit operations.
-(c) Mapping of subcircuits to physical chiplets. Since chiplets $A$ and $C$ are not physically connected, certain gates (shown in red) necessitate inter-chiplet routing.
-(d) Inter-chiplet routing is performed to address cross-chiplet gates that either don’t conform to the physical inter-chiplet topology or the basis gate set (aren’t SWAP gates).
-Figure 2. Showcase of Inter-Chiplet Compilation in the Stratification Stage
+(a) A chiplet topology
+(b) Qubit-to-Subcircuit Mapping
+(c) Subcircuit-to-Chiplet Routing
+(d) Inter-Chiplet Routing
+Figure 3. Step-by-step example of inter-chiplet compilation in the Stratification stage.
 In this stage, we map a quantum circuit to the inter-chiplet device topology, e.g., the structure in Figure 3(a), by stratifying it into smaller, interconnected subcircuits.
-This can be accomplished with the following intermediate steps, as shown in Figures 2(b)2(d).
+This can be accomplished with the following intermediate steps, as shown in Figures 3(b)3(d).
 First, qubit-to-subcircuit mapping designates which logical qubits belong to which logical subcircuits.
@@
 4.1.1. Qubit-to-Subcircuit Mapping
-To perform qubit-to-subcircuit mapping (Figure 2(b)), we leverage a form of simulated annealing.
-Simulated annealing is a form of stochasic optimization, notably used in VLSI algorithms for place-and-route (annealing1989).
-We use simulated annealing to find a qubit assignment that minimizes the number of connections between subcircuits.
-As our chiplet allocation algorithm acts conservatively—performing the minimal number of changes necessary to make a circuit valid—it will produce better results if given a circuit with minimal connectivity between subcircuits.
-For our cost function, we count the number of inter-subcircuit gates.
-To permute the solution, we swap two random qubits in different subcircuits.
-As different simulated annealing runs are independent, running the algorithm using different seeds is embarrassingly parallel.
-We exploit this parallelism by running many trials at once, each in a different process with a different seed, then choosing the best result across all trials and seeds.
+Algorithm 1 Qubit-to-Subcircuit Mapping
+[…]
+[…]
+2:$q_{0},q_{1}\leftarrow\textsc{Qubits}(g)$
+3:$s_{0},s_{1}\leftarrow L[q_{0}],L[q_{1}]$
+4:Increment $C$ if $s_{0}\neq s_{1}$EndFor
+[…]
+[…]
+7:$s_{0}\leftarrow\textsc{RandChoice}(\{1,\dots,N_{C}\})$
+8:$q_{0}\leftarrow\textsc{RandChoice}(L^{\prime}[s_{0}])$
+9:$s_{1}\leftarrow\textsc{RandChoice}(\{1,\dots,N_{C}\}-\{s_{0}\})$
+10:$q_{1}\leftarrow\textsc{RandChoice}(L^{\prime}[s_{1}])$LCommentSwap $q_{0}$ and $q_{1}$
+11:$L^{\prime}[s_{0}]\leftarrow(L^{\prime}[s_{0}]\cup\{q_{1}\})-\{q_{0}\}$
+12:$L^{\prime}[s_{1}]\leftarrow(L^{\prime}[s_{1}]\cup\{q_{0}\})-\{q_{1}\}$EndFor
+[…]
+14:$L\leftarrow$ Initial Layout $\triangleright$ Current Solution
+15:$R\leftarrow L$triangleright Best Solution
+[…]
+17:$D\leftarrow\textsc{Ceil}(T\times\alpha)$triangleright Number of changes to make
+18:$L^{\prime}\leftarrow\textsc{Kick}(D,N_{C},N_{Q},L)$
+19:$\Delta\leftarrow\textsc{Cost}(Circ,L^{\prime})-\textsc{Cost}(Circ,L)$
+20:$k\leftarrow\textsc{Rand}(0,1)$triangleright Random number in $[0,1)\subset\mathbb{R}$
+21:$L\leftarrow L^{\prime}$ if $(\Delta<0)$ or $(k>\exp(\Delta/T))$
+22:$R\leftarrow L^{\prime}$ if $(\textsc{Cost}(Circ,L^{\prime})<\textsc{Cost}(Circ,R))$
+[…]
+24:\$R$EndProcedure
+[…]
+Qubit-to-subcircuit mapping produces a stratified quantum circuit, where logical qubits are grouped into subcircuits, as shown in Figure 3(b).
+An optimal qubit-to-subcircuit mapping minimizes the number of connections between subcircuits, thereby reducing the number of inter-chiplet SWAPs needed to transfer data between chiplets during execution.
+To perform this task, we leverage simulated annealing, an efficient stochastic optimization technique shown to work well on NP-complete problems, such as place-and-route in VLSI algorithms (annealing1989).
+Simulated annealing is an iterative method to optimize a state against a given cost function.
+In each iteration, the state is “kicked,” or randomly perturbed, and the new state is greedily accepted if there is a reduction in cost.
+However, the algorithm may randomly also accept worse solutions to escape from local minima wells.
+The runtime of simulated annealing is statically determined using constant tuning parameters, thus the overall complexity is determined by the complexity of the cost and kick functions, which are typically polynomial.
+Additionally, since each run of a simulated annealing is randomized and independent, it is possible to run multiple trials in parallel with different initial random seeds, then pick the best result.
+This improves the chances of finding a (near-)optimal solution.
+Our method is shown in Algorithm 1.
+We start with an initial qubit-to-subcircuit mapping, which we minimize against a cost function, defined on line 1 as the number of inter-chiplet gates given a certain layout.
+By minimizing the number of inter-chiplet gates, we implicitly reduce the inter-chiplet connectivity.
+The kick function then swaps random qubits in the layout, ensuring that each SWAP is between qubits in different subcircuits.
+In doing so, we hope to move strongly interacting qubits close-together, minimizing the number of cross-chiplet gates.
+As input, our algorithm takes the input circuit, the chiplet count $N_{C}$, the chiplet size $N_{Q}$, the starting temperature $T_{0}$, the number of changes to do per degree of temperature $\alpha$, and the cooling modifier $\beta$.
+The last three are tuning parameters that do not change across runs.
 4.1.2. Chiplet Allocation
-To perform chiplet-level allocation, shown in Figures 2(c) and 2(d), we extend and modify the SABRE algorithm for initial mapping (sabre_swap) with application-specific layout and routing passes.
-For layout, we initialize each iterative allocation cycle with a random placement of logical subcircuits into physical chiplets.
-For routing, we developed a new heuristic for deciding between candidate SWAPs.
-By default, SABRE attempts to place two-qubit gates on a valid edge in the graph of a device’s hardware topology, i.e, the physical distance between the qubits of a gate should be $1$.
-For inter-chiplet routing, we do not wish to factor intra-qubit topology into our distance calculation.
-Furthermore, we hope to move all gates to be within the same chiplet, i.e., the distance between qubits should be 00 instead of $1$.
-Thus, when selecting between potential SWAPs, we categorize and prioritize SWAPs in a three-tier approach:
-  1. (1)
-    Symbiotic SWAPs: A SWAP between two multi-chiplet gates is considered symbiotic if it reduces the distance of both gates, i.e., each gate benefits from the SWAP.
-In theory, we should expect symbiotic SWAPs to be the most efficient, and thus highest-priority, SWAP possible.
-  2. (2)
-    Commensalistic SWAPs: When a SWAP can benefit one multi-chiplet gate without harming any other, we consider it to be commensalistic.
-Such behavior occurs most commonly on idle qubits in the quantum circuit, although it can occur between two multi-chiplet gates.
-  3. (3)
-    Parasitic SWAPs: The remaining category of SWAPs to consider benefit one gate while harming another.
+Algorithm 2 Inter-Chiplet Routing
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+8:$V_{virt},V_{phys}\leftarrow\textsc{ValidGates}(Q,L)$
+[…]
+[…]
+[…]
+12:$L^{\prime}\leftarrow\textsc{Swap}(\textsc{Copy}(L),S)$
+[…]
+14:$q_{0},q_{1}\leftarrow\textsc{Qubits}(g)$
+15:$s_{0},s_{1}\leftarrow L[q_{0}],L[q_{1}]$
+[…]
+[…]
+18:$D\leftarrow\textsc{CircuitToDAG}(Circ)$
+19:$E\leftarrow$ All-pairs distances from chiplet topology $T$
+20:$Q\leftarrow\textsc{FirstLayer}(DAG)$
+21:$R\leftarrow$ empty list
+[…]
+23:$\mathcal{S}\leftarrow\textnormal{{FindSwaps}}(Q,E)$triangleright SWAP Candidates
+24:$S\leftarrow\operatorname*{arg\,min}_{s\in\mathcal{S}}\textsc{Heuristic}(s,R,Circ)$
+25:$\textsc{Swap}(L,S)$
+26:$\textsc{Step}(Q,R,L,D)$EndWhile
+27:$D^{\prime}\leftarrow$ Empty DAG
+28:Append $g$ to $D^{\prime}$ for each $g\in{R}$triangleright $R$ is ordered
+[…]
+[…]
+Chiplet allocation describes the process of transforming a stratified circuit to be compatible with the inter-chiplet topology of a modular device.
+In Figure 2, we decompose chiplet allocation into subcircuit-to-chiplet mapping and inter-chiplet routing, analogous to how qubit allocation can be decomposed into qubit layout and routing.
+Accordingly, subcircuit-to-chiplet mapping, shown in Figure 3(c), defines an initial placement of logical subcircuits to physical chiplets.
+Subsequently, inter-chiplet routing, shown in Figure 3(d), inserts inter-chiplet SWAPs to account for limited cross-chiplet gate connectivity.
+In our design, we do not explictly derive a subcircuit-to-chiplet mapping.
+Rather, we extend and modify the SABRE algorithm (sabre_swap) to apply to chiplets, thereby using iterations of inter-chiplet routing with randomized permutations to inform a subcircuit-to-chiplet mapping.
+Inter-chiplet routing seeks to consolidate the qubits of (non-SWAP) inter-chiplet gates to a single chiplet by inserting inter-chiplet SWAPs into the stratified circuit.
+For an single two-qubit, inter-chiplet gate, we represent inter-chiplet routing as a shortest-path problem, where the two qubits of the gate correspond to chiplet endpoints in the unweighted inter-chiplet connectivity graph.
+From the shortest path, we can derive an optimal sequence of SWAPs for routing, where the unweighted path length, or chiplet distance, indicates the length of the SWAP chain.
+However, in the context of a entire stratified circuit, composed of multiple inter-chiplet gates, it becomes possible to interleave otherwise efficiently reorder execution of SWAP chains to reduce the depth and/or total number of inter-chiplet SWAPs.
+Like qubit routing, we would expect it to be impractical to derive an optimal solution for this more complex picture of inter-chiplet routing at compile-time.
+Thus, we employ a similar SWAP selection scheme as SABRE (sabre_swap), where inserted SWAPs are chosen from a pool of SWAP candidates according to a cost heuristic.
+For our inter-chiplet routing algorithm, we categorize SWAPs from our candidate pool into three tiers, based on a concept of the “efficiency” of each SWAP, as shown in Algorithm 2.
+An inter-chiplet SWAP is considered symbiotic if it directly affects two inter-chiplet gates in a “beneficial” manner.
+In other words, each qubit of a symbiotic SWAP corresponds to a different inter-chiplet gate, where the insertion of the SWAP reduces the chiplet distance score of both gates.
+In theory, this outcome represents the most efficient possible insertion for a single SWAP, and thus should be the highest-priority case for our greedy inter-chiplet routing cost heuristic.
+Next, commensalistic SWAPs are those that do not “harm,” or increase that chiplet distance, of any directly impacted gates.
+These SWAPs typically occur when an idle qubit can be used as an ancilla qubit to facilitate the routing of an inter-chiplet gate without affecting other inter-chiplet gates.
+In certain device topologies, a SWAP acting on the qubits of two inter-chiplet gates can also exhibit commercialistic behavior.
+Finally, parasitic SWAPs act on the qubits of two inter-chiplet gates, where one gate benefits while the other is harmed.
 Since this variety of SWAP is actively counterproductive towards a gate, it should be, and likely will be, avoided in most circumstances.
-However, given a particularly onerous circuit, it may be necessary in such rare instances.
-For example, clear one gate may free physical qubits to help route others.
-With these changes, our algorithm is able to correctly and efficiently place and route circuits for any valid device topology.
+However, given a particularly onerous circuit, it may be necessary in such rare instances to break out of a deadlock.
+For example, clearing one gate may free physical qubits to help route others.
 4.2. Elaboration Stage
-(a) A $3$-qubit chiplet with a line topology and connected to a neighboring chiplet by $q_{2}$.
-(b) A $3$-qubit subcircuit with $2$ inter-chiplet SWAPs.
-[…]
-(d) A greedy policy takes the shortest path for routing inter-chiplet gates to their nearest halo qubit.
-(e) Qubit routing for the remaining intra-chiplet gates is performed as normal, except with restrictions to ensure inter-chiplet SWAPs are kept static.
-(f) The remaining compilation states (basis translation and optimization) are not affected by the additional constraints imposed by inter-chiplet gates.
-Figure 3. Showcase of Intra-Chiplet Compilation in the Elaboration Stage.
+(a) An intra-chiplet topology
+(b) A $3$-qubit subcircuit post-stratification.
+(c) Subcircuit after Intra-Chiplet Layout
+(d) Inter-Chiplet SWAP Lowering
+(e) Intra-Chiplet Routing
+(f) Basis Translation and Optimization
+Figure 4. Step-by-step example of intra-chiplet compilation in the Elaboration stage. Gates that cannot be implemented on hardware are shown in red.
 Following the stratification stage, each chiplet can be compiled almost completely independently from one another.
@@
 Once this assignment is made, these SWAPs must remain fixed and immutable to subsequent compilation stages.
-Despite this overhead, restricting the task of compilation to a narrower domain of qubits facilitates both parallelization (as each subcircuit can be processed in parallel) and problem size reduction (as there are less qubits to consider) of the most computationally intensive compilation stages (routing and optimization).
+Despite this overhead, restricting the task of compilation to a narrower domain of qubits facilitates both parallelization (since each subcircuit can be processed in parallel) and problem size reduction (since there are fewer qubits to consider) of the most computationally intensive compilation stages, namely routing and optimization.
 To accomplish these tasks, we divide the elaboration stage into two substages: qubit allocation and parallel basis translation and optimization.
@@
 4.2.1. Qubit Allocation
-Like other qubit allocation schemes, we incorporate layout (see Figure 3(c)) and routing (see Figure 3(e)) stages.
-The presence of inter-chiplet gates necessitates several key changes to the SABRE (sabre_swap) qubit allocation algorithm to achieve correctness and/or maximum parallelism.
+Like other qubit allocation schemes, we incorporate layout (see Figure 4(c)) and routing (see Figure 4(e)) substages.
+However, the presence of inter-chiplet gates also necessitates several key changes to the SABRE (sabre_swap) qubit allocation algorithm to achieve correctness and/or maximum parallelism.
 First, the placement of inter-chiplet gates is a non-parallelizable task (due to cross-chiplet interactions) that strongly influences the layout and routing passes, which could otherwise be fully parallelized across chiplets.
-To account for this, we introduce a serial stage (shown in Figure 3(d)) between the parallelized layout and routing passes that greedily places inter-chiplet gates on the nearest valid edge between halo qubits.
+To account for this, we introduce a serial stage, as shown in Figure 4(d), between the parallel layout and routing passes that greedily places gates on the nearest valid inter-chiplet edge.
 A greedy policy was chosen because inter-chiplet gates have more restrictions on valid placements compared to intra-chiplet gates, while also experiencing substantially higher error rates (smith_2022_chiplets).
-Once the positions of the inter-chiplet gates are settled, they must remain fixed in place, imposing an additional constraint on qubit routing.
+Once the positions of the inter-chiplet gates are settled, they must remain fixed, imposing an additional constraint on qubit routing.
 To promote this behavior in SABRE, we extended its cost heuristic to weigh the device topology by fidelity.
@@
 This is the traditional gate set for a monolithic processor.
-The other gate set is the only one permitted to operate across chiplets, and specifically only implements inter-chiplet SWAPs, i.e., SWAP gates between two halo qubits.
+The other gate set is the only one permitted to operate across chiplets, and specifically only implements inter-chiplet SWAPs.
 By restricting each gate set to its own disjoint part of the overall device topology, we ensure correctness across all cases.
@@
 The remaining compilation stages of basis translation and circuit optimization can be easily parallelized at the chiplet granularity, as shown in Figure 4(f).
-Basis translation, in particular, is embarrassingly parallel, as every gate can be translated independently from every other gate.
+Basis translation, in particular, is embarrassingly parallel, because every gate can be translated independently from every other gate.
 Meanwhile, as discussed in Section 2, optimization algorithms already employ partitioning and parallelization to create tractable subproblems from intractably large circuits, facilitating a smaller problem space, lower memory usage, and faster performance.
@@
 When given a parallelizable workload, SEQC attempts to allocate its tasks evenly among the available physical cores, rounded to the neared whole number.
-Finally, the simulated annealing technique used in the qubit-to-subcircuit stage uses a starting temperature $T_{0}=200$K, a rate of change of $0.005T$ each round, $\frac{T}{50}$ permutations per round, and a total of $5$ trials per core.
+Finally, the simulated annealing technique used in the qubit-to-subcircuit stage uses a starting temperature $T_{0}=200$, a rate of change of $0.005T$ each round, $\frac{T}{50}$ permutations per round, and a total of $10$ trials per CPU core.
 5.1. Metrics
@@
 The stratification time represents a one-time overhead to prepare a quantum circuit for execution on a modular architecture.
-This result can be reused for future executions on that device or any device with an equivalent or more expansive topology.
+This result can be reused for future executions on any device with an equivalent or more expansive topology.
 Recurring elaboration time.
@@
 5.2. Benchmarks
-For our benchmark suite, we leverage a subset of circuits in the Supermarq (supermarq) suite of quantum circuits.
-The circuits we select are chosen to be representative of practical, real-world applications for quantum computing, while also being feasible to compile for a large number of qubits.
+(a) BitCode
+(b) PhaseCode
+(c) GHZ
+(d) VQE
+(e) Hamiltonian Simulation
+Figure 5. Examples of the Benchmark Circuits
+For our benchmark suite, we leverage a subset of circuits from the Supermarq (supermarq) suite.
+The circuits are chosen to be representative of practical, real-world applications for quantum computing, while also being feasible to compile for a large number of qubits.
 In particular, our benchmark suite comprises the following quantum circuits.
@@
 […]
-[…]
+Greenberger–Horne–Zeiling (GHZ).
+Entanglement is one of the fundamental properties of quantum states and a major source of quantum advantage.
+The GHZ state represents a maximally entangled $n$-qubit state of $\frac{1}{\sqrt{2}}\left(\ket{|}0>^{\otimes n}+\ket{|}1>^{\otimes n}\right)$.
+The algorithmic circuit structure is essentially a single chain of CNOT gates, as shown in Figure 5(c) From a compilation perspective, this represents a pure exercise in qubit allocation, because the logical circuit is entirely serial.
+An $n$-qubit GHZ circuit that fully utilizes a $k$-chiplet device should necessitate $\lfloor\frac{n}{k}\rfloor-1$ logical CNOT operations to extend between chiplets.
+In turn, this will necessitate at least $2\left(\lfloor\frac{n}{k}\rfloor-1\right)$ inter-chiplet SWAPs to transfer qubit in and out of a chiplet.
 VQE.
 The variational quantum eigensolver (VQE) (vqe) is another hybrid quantum-classical variational algorithm that is often used in chemistry applications.
-In our experiments, we choose a $2$-layer ansatz whose parameters are once again randomly generated.
+In our experiments, we choose a $2$-layer ansatz whose parameters are again randomly generated, as shown in Figure 5(d).
+This results in a quantum circuit that is essentially two CNOT chains separated by single-qubit operations.
+The CNOT chains can be interleaved, which in conjunction with the single-qubit gates, creates many opportunities for parallelism (resembling pipelining in classical computing).
+Moreover, from a qubit routing perspective, the second CNOT chain forces the compiler to consider how qubit permutations in one circuit layer propagate to future layers.
 Hamiltonian Simulation.
@@
 The Supermarq benchmark selects the 1D Transverse Field Ising Model (TFIM) system to model, which is relevant for phase transitions in magnetic materials.
+The circuit for Hamiltonian Simulation, shown in Figure 5(e), resembles a slightly more complex GHZ circuit, including how it is completely serial.
+Compared to GHZ, this circuit should test the compiler’s ability to lookahead during routing to detect back-to-back gates on the same two-qubit pair, while also emphasizing parallelism originating from the efficient placement of SWAP gates.
 5.3. Simulated Quantum Device Specifications
@@
 The tessellation requirements of each chiplet restrict valid chiplet sizes to multiples of $10$; however, for our experiments, we limit the size of each chiplet to the minimum of $10$ qubits due to hardware limitations.
-Note that for each valid chiplet count, there may exist multiple valid chiplet topologies.
+Finally, note that for each valid chiplet count, there may exist multiple valid chiplet topologies.
 In this work, we generate our backends to use the “most-square” topology for a given chiplet count.
@@
 This aligns with our model of inter- vs intra-chiplet gates, where inter-chiplet SWAPs are assumed to have 4×4$\times$4 × the error (smith_2022_chiplets) and 4×4$\times$4 × the gate duration.
-Figure 11. Relative circuit depth of SEQC ormalized to the strawman baseline for devices with 10-qubit chiplets (lower is better).
+Figure 13. Relative circuit depth of SEQC normalized to the Strawman baseline for devices with 10-qubit chiplets (lower is better).
 Figure 14. Relative gate count of SEQC normalized to the Strawman baseline for devices with 10-qubit chiplets (lower is better).
-Relative circuit depth and gate count, shown in Figures 11 and 12, respectively, share similar behavior and analysis.
+Relative circuit depth and gate count, shown in Figures 13 and 14, respectively, share similar behavior and analysis.
 At small circuit sizes, SEQC has higher circuit depth and gate count than the baseline.
@@
 Their study is motivated by application-specific optimization – they seek techniques that allow a quantum circuit’s structure to influence hardware architecture.
-In their work, the relationship between circuit structure and mapping efficiency is defined, and this information optimizes the architecture of single- and multi-core quantum processors.
-Piveteau and Sutter (circuit_knitting) investigate the benefits of using quasiprobability simulation to allow subcircuits cut from larger quantum circuits to perform their quantum operations independently, simulating halo qubit behavior where the subcircuits connect.
+They define the relationship between circuit structure and mapping efficiency, utilizing this information to optimize the architecture of single- and multi-core quantum processors.
+Piveteau and Sutter (circuit_knitting) investigate the benefits of using quasiprobability simulation to allow subcircuits cut from larger quantum circuits to perform their quantum operations independently, simulating inter-chiplet link behavior where the subcircuits connect.
 When augmented with classical communication between circuits to reduce the simulation overhead incurred by sampling, they demonstrate that resulting circuits reach fixed accuracy with less simulation time.
@@
 However, HQA operates under the assumption that all inter-chiplet communications have the same cost, meaning chiplets must be fully connected.
-SEQC, on the other hand, is able to work with any arrangement of chiplets and inter-chiplet links (as long as there is a path between any two chiplets), taking into account the distance between pairs of chiplets to generate improved circuits.
+SEQC, on the other hand, is able to work with any arrangement of chiplets and inter-chiplet links so long there is a path between any two chiplets, taking into account the distance between pairs of chiplets to generate improved circuits.
 Iterating on their previous work on Hungarian Qubit Assignment, Escofet et al.
@@
 Eyeriss (Eyeriss_2016_Chen) approaches the dataflow mapping problem by proposing a row-stationary architecture, and then optimizing the mapping of problem to processing element such that elements which are spatially local in the problem are spatially local in hardware during execution.
-This is not unlike qubit swap optimization, and in fact is performed for many of the same reasons—like inter-chiplet swaps, data movements between more distant classical programming elements in a dataflow architecture are simply more expensive, and minimizing their frequency yields performance improvements.
+This is not unlike qubit swap optimization, and in fact is performed for many of the same reasons — like inter-chiplet swaps, data movements between more distant classical programming elements in a dataflow architecture are more expensive, and minimizing their frequency yields performance improvements.
 8. Conclusion
@@
 However, the complexity of chiplet-based quantum devices, coupled with their growing size, presents an imminent scalability challenge for quantum compilation.
-Existing qubit allocation methods are often unable to contend with inter-chiplet links, which don’t necessary support a universal basis gate set, and existing methods of logical-to-physical qubit placement, routing, unitary synthesis, and quantum circuit optimization, are typically not designed for qubit links of wildly varying levels of duration or fidelity.
+Existing qubit allocation methods are often unable to contend with inter-chiplet links, which don’t necessary support a universal basis gate set, and existing methods of logical-to-physical qubit placement, routing, unitary synthesis, and quantum circuit optimization, are typically not designed for qubit links that vary significantly in terms of duration or fidelity.
 In this work, we first modify stock contemporary compilers to become aware of hardware modularity, thereby allowing them to correctly compile circuits for modular architectures with limited cross-chiplet gate support.
@@
 Owning to these techniques and its parallelization strategy, SEQC compiles circuits that exhibit up to a 36% increase in circuit fidelity and up to 1.92×1.92$\times$1.92 × lower execution time, while consistently achieving 24×2-4$\times$2 - 4 × faster compilation time over a chiplet-aware Qiskit baseline.
+Acknowledgements
+This research was partially supported by NSF awards CNS-1763743 and CCF-2119069, Academic Year Undergraduate Research Award #2023-17704 by Northwestern University, and undergraduate summer research awards by the Robert R.
+McCormick School of Engineering and Applied Science and the Computer Science Department at Northwestern University.
+The authors thank Kabir Dubey for helping with earlier versions of this manuscript.
+This research was supported in part through the computational resources and staff contributions provided for the Quest high performance computing facility at Northwestern University which is jointly supported by the Office of the Provost, the Office for Research, and Northwestern University Information Technology.
+We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.
 […]

### v3

@@
 Abstract.
-As quantum computing technology continues to mature, industry is adopting modular quantum architectures to keep quantum scaling on the projected path and meet performance targets.
+As quantum computing technology matures, industry is adopting modular quantum architectures to keep quantum scaling on the projected path and meet performance targets.
 However, the complexity of chiplet-based quantum devices, coupled with their growing size, presents an imminent scalability challenge for quantum compilation.
-Contemporary compilation methods are not well-suited to chiplet architectures — in particular, existing qubit allocation methods are often unable to contend with inter-chiplet links, which don’t necessary support a universal basis gate set.
-Furthermore, existing methods of logical-to-physical qubit placement, swap insertion (routing), unitary synthesis, and/or optimization are typically not designed for qubit links of significantly varying levels of duration or fidelity.
-In this work, we propose SEQC, a complete and parallelized compilation pipeline optimized for chiplet-based quantum computers, including several novel methods for qubit placement, qubit routing, and circuit optimization.
-SEQC attains up to a 36% increase in circuit fidelity, accompanied by execution time improvements of up to 1.92×1.92$\times$1.92 ×.
-Additionally, owing to its ability to parallelize compilation, SEQC achieves consistent solve time improvements of 24×2-4$\times$2 - 4 × over a chiplet-aware Qiskit baseline.
+Contemporary compilation methods are not well-suited to chiplet architectures — in particular, existing qubit allocation methods are often unable to contend with inter-chiplet links, which don’t necessarily support a universal basis gate set.
+Furthermore, existing methods of logical-to-physical qubit placement, swap insertion (routing), unitary synthesis, and/or optimization, are typically not designed for qubit links of significantly varying latency or fidelity.
+In this work, we propose SEQC, a hierarchical parallelized compilation pipeline optimized for chiplet-based quantum systems, including several novel methods for qubit placement, qubit routing, and circuit optimization.
+SEQC attains a $9.3\%$ average increase in circuit fidelity (up to $49.99\%$).
+Additionally, owing to its ability to parallelize compilation, SEQC achieves 3.27×3.27$\times$3.27 × faster compilation on average (up to 6.74×6.74$\times$6.74 ×) over a chiplet-unaware Qiskit baseline.
 Quantum computing, quantum compilation, modular architectures.
@@
 We are already observing signs of this shift.
-Recent developments in quantum chip linking, including flip-chip architectures (gold2021entanglement) and low-loss coaxial cables (niu2023low), suggest that modular designs are the most viable for scaling quantum computers (smith_2022_chiplets).
+Recent developments in quantum chip linking, including flip-chip architectures (gold2021entanglement) and low-loss coaxial cables (niu2023low), suggest that modular designs are the most viable for scaling (smith_2022_chiplets).
 Many contemporary or upcoming leading quantum systems adopt chiplet-based modular quantum processors, for example using high-bandwidth quantum links between nearest-neighbor quantum chiplets (e.g., carrier-chip couplers in Rigetti Aspen-M (Rigetti_AspenM; gold2021entanglement), or m-couplers in IBM Crossbill (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers)), or lower-fidelity, lower-bandwidth, but longer-distance flexible coupling of discrete chips (e.g., l-couplers in IBM Flamingo (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers)), or a combination of the above (e.g., c-, m-, and l-couplers in IBM Starling (Mandelbaum_2024_IBMCouplers)), or multi-chip connectivity through tunable couplers and routing chips (Field_2024_RigettiModular).
 Most major companies have set their sights at modular designs to meet quantum scaling targets in practical ways, and modular quantum processors dominate their latest roadmaps (IBM_2023_Roadmap; IonQ_2024_Roadmap; Rigetti_2024_Roadmap).
-Unfortunately, current compilation infrastructure is not capable of reasoning well about this quantum interconnect heterogeneity.
+Unfortunately, current compilation infrastructure does not support quantum interconnect heterogeneity.
 The mapping of a quantum program’s logical quantum gates into the native gates supported by the underlying quantum processor, and the mapping of logical qubits to the physical qubits of the quantum hardware, largely determine the number and type of native quantum gates required for the computation, the circuit depth, and its execution time—long programs may not complete successfully as qubits decohere.
@@
 To make matters more challenging, each qubit, coupler, and gate have diverse error profiles that are highly variable both spatially and temporally (Tannu_2019_ErrorVariability; Murali_2019_NoiseMapping; dasgupta2021stability).
-Unlike classical compilation that is done only once for a given architecture, quantum programs must be recompiled every time before execution, as the ideal physical qubits to execute on change between runs.
+Unlike classical compilation that is done only once for a given architecture, quantum programs must be recompiled every time before execution [^1] The term “execution” here includes all shots required to produce a final result for the user.
+Different executions of the same circuit can span qubit calibration windows., as the ideal physical qubits to target change between runs.
 So, quantum compilers today are faced with the daunting task of optimizing quantum programs across multiple dimensions with often conflicting demands, in a continuously changing environment.
 Coupled with the hardness of synthesizing unitaries (which is exponential in the number of qubits) and the need for quick and frequent compilation, modern quantum compilers have no option other than to rely on heuristics to perform the task.
-These heuristics result in a compilation complexity that we experimentally observed to be $O(n^{2})$ for $n$ total qubits in a quantum processor, as shown in Figure 8.
+These heuristics result in a compilation complexity that is quadratic to the number of qubits in the quantum processor (experimentally observed in Figure 12).
 Hardware modularity adds significant complexity to this already hard task, as inter-chiplet links are typically inferior compared to intra-chiplet ones (laracuente2022modeling; smith_2022_chiplets), connectivity across chiplets is often limited (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers), and not all basis gates are necessarily supported across chiplets (niu2023low).
@@
 This Stratify-Elaborate Quantum Compiler (SEQC) stratifies a source program only once for a given chiplet architecture, and performs only the elaboration step recurrently before each execution.
-In essence, SEQC replaces the recurrent $O(n^{2})$ compilation step for an $n$-qubit quantum processor, with several parallel $O(k^{2})$ elaboration steps for $k$-qubit chiplets.
-As the qubit capacity $n$ of quantum processors grows exponentially, today’s $O(n^{2})$ compilation latency rises even faster.
-We expect, however, that the number of qubits $k$ per chiplet will remain relatively stable or grow much slower, as it seems to be the case for the foreseeable future (IBM_2023_Roadmap), and thus the SEQC recurrent compilation latency is expected to remain relatively stable.
-The stratification step is $O(n^{2})$, so the end-to-end complexity remains the same, but stratification is performed only once; the recurrent compilation in SEQC is only $O(k^{2})$, and barely growing with new processor designs.
+As the qubit capacity $n$ of quantum processors grows exponentially, the $O(n^{2})$ compilation latency of contemporary compilers rises even faster.
+SEQC replaces the recurrent $O(n^{2})$ compilation step with several parallel $O(k^{2})$ elaboration steps for $k$-qubit chiplets.
+The stratification step is $O(n^{2})$, so the end-to-end complexity remains the same, but stratification is performed only once.
+The recurrent compilation in SEQC is only $O(k^{2})$, and as the number of qubits $k$ per chiplet is expected to remain relatively stable, or grow at a much slower pace in the foreseeable future (IBM_2023_Roadmap), SEQC’s recurrent compilation latency will barely grow as processors scale.
 Additionally, as SEQC is cognizant of hardware modularity, it can stratify the source quantum circuit into subcircuits and map them to chiplets to minimize inter-chiplet communication.
-As we show in this paper, SEQC produces circuits with shorter execution times and significantly fewer inter-chiplet gates compared to today’s stock compilers, leading to much higher fidelity execution.
+As we show in this paper, SEQC produces circuits with shorter execution times and significantly fewer inter-chiplet gates compared to today’s stock compilers, leading to higher execution fidelity.
 More importantly, as the number of qubits in a processor grows, SEQC achieves even higher performance in these figures of merit.
@@
   • We design and implement in SEQC several novel methods for qubit placement, qubit routing, and circuit optimization.
-  • We evaluate SEQC and show it compiles circuits with up to 36% higher circuit fidelity and up to 1.92×1.92$\times$1.92 × lower execution time, while consistently achieving 24×2-4$\times$2 - 4 × faster compilation time compared to a chiplet-aware Qiskit baseline.
+SEQC compiles circuits with $9.3\%$ higher circuit fidelity on average (up to $49.99\%$), while consistently achieving 3.27×3.27$\times$3.27 × faster compilation time on average (up to 6.74×6.74$\times$6.74 ×) compared to a chiplet-unaware Qiskit baseline.
 2. Background and Motivation
-(a) A chiplet quantum architecture
-(b) Input circuit from SABRE
+(a) Example 2-chiplet architecture
+(b) Input circuit from qubit allocation
 (c) Insertion of perfect-shuffle
@@
 The relations between each stage and the physical constraints of both monolithic and chiplet architectures are described below.
-2.1. Qubit Allocation
+Qubit Allocation
 Qubit allocation addresses the topology constraints of quantum devices, where physical qubits on a device have limited connectivity to other qubits on the device.
-It is critical that after this stage, which often requires the modification or addition of circuit gates, the resulting optimized and technology-dependent circuit is functionally equivalent to the original, technology-independent algorithm (smith2019quantum).
-The qubit allocation problem has been previously shown to be NP-complete (qubit_allocation); thus, algorithms often adopt various strategies to make the problem more tractable.
+It is critical that after this stage, which often requires the modification or addition of circuit gates, the resulting optimized, technology-dependent circuit is functionally equivalent to the original, technology-independent algorithm (smith2019quantum).
+The qubit allocation problem has been previously shown to be NP-complete (qubit_allocation).
+Thus, algorithms often adopt various strategies to make the problem more tractable.
 One common method is to further divide qubit allocation into two substages: layout and routing.
@@
 This assumption can manifest implicitly when a given device topology is represented as an unweighted, undirected graph.
-For monolithic architectures, one could debate the marginal tradeoff between performance and solution quality; however, for modular architectures, the cross-chiplet connections are far worse compared to intra-chiplet connections (smith_2022_chiplets).
+For monolithic architectures, one could debate the marginal tradeoff between performance and solution quality.
+However, for modular architectures, the cross-chiplet connections are far worse compared to intra-chiplet connections (smith_2022_chiplets).
 Thus, we could expect the solutions generated from the simplified model to be worse than a more physically accurate model, for example, in circumstances where the shortest path in terms of the quantity or depth of inserted SWAPs diverges from the shortest path in terms of fidelity, gate duration, or some combination thereof.
-2.2. Basis Translation
+Basis Translation
 Basis translation converts the gates of an input quantum circuit into gates that are physically supported on hardware.
@@
 In the general case, the cross-chiplet links can only be used for data movement (such as SWAPs) and not computation, meaning these links may not support universal operations (niu2023low).
-Thus, compilers that depend on the expectation of universality among qubit links (qiskit_basis_translation_errors) and do not expect these heterogenous, non-universal basis gate sets can generate incompatible outputs when targeting modular architectures.
-2.3. Optimization
+Thus, compilers that depend on the expectation of universality among qubit links (qiskit_basis_translation_errors), and do not expect these heterogenous, non-universal basis gate sets, can generate incompatible outputs when targeting modular architectures.
+Optimization
 Quantum circuit optimization seeks to prune extraneous operations by finding opportunities to combine, eliminate, and/or parallelize quantum gates (optimization1).
 Through commutation and cancellation identities, it is possible to search for optimization opportunities by iteratively transforming the quantum circuit without affecting the underlying operation (optimization1).
-However, in general, optimization is highly time- and resource-intensive, with the problem scaling exponentially with respect to number of qubits and/or circuit depth (qfactor).
+However, in general, optimization is highly time and resource intensive, with the problem scaling exponentially with respect to number of qubits and/or circuit depth (qfactor).
 As a result, partitioning, the division of a quantum circuit into blocks of quantum gates, is a common tactic among optimization techniques (qfactor; qgo; bqskit).
 One major consideration in this process is demarcation, i.e., where and when to draw the line between blocks.
-Namely, one forfeits all optimization opportunities along boundaries of demarcation, so they must be carefully chosen to minimize negative impact on the output solution.
+Namely, one forfeits all optimization opportunities along boundaries of demarcation, so they must be carefully chosen to minimize the negative impact on the output solution.
 From this perspective, modular architectures are naturally synergistic with partitioning, since cross-chiplet connections present a physically motivated boundary between operations.
-3. A Strawman Chiplet Compiler
+3. Compilation for Modular Quantum Systems
 Figure 2. High-level diagram of SEQC. Elements in red reflect modifications from existing compiler designs.
-As discussed previously in Section 2, our model of modular quantum architectures includes the more general case where inter-chiplet links are can be non-universal.
+Modular quantum architectures introduce new challenges to quantum compilation that impact the correctness of solutions and the speed of compilation.
+As discussed previously in §2, our model of modular quantum architectures considers the possibility that inter-chiplet links can be non-universal.
 This is representative of hardware manufacturing constraints (niu2023low), where only a restricted set of operations, e.g., state transfers or SWAPs, may be supported over long-distance couplers (inter-chiplet links).
-However, many compiler designs (including designs for modular architectures, see Section 7) experience errors or crashes when presented with non-universal, heterogenous, basis gate sets.
+However, many compiler designs (including designs for modular architectures, see §7) experience errors or crashes when presented with non-universal, heterogenous, basis gate sets.
 For example, the Qiskit compiler (qiskit2024) is explicit in its compatibility for only universal basis gate sets (qiskit_basis_translation_errors).
@@
 Specifically, the Qiskit routing stage (zou2024lightsabrelightweightenhancedsabre) (an implementation of SABRE) will freely place gates on incompatible physical links, and there are no methods in the compiler to prevent or fix consequent errors.
-For the purposes of this work, however, it was critical to establish a baseline representative of the current state of quantum compilation, against which we can compare our more-optimized design (see Section 4).
-Thus, we created a “Strawman” baseline compiler, built on top of the Qiskit compiler (qiskit2024), with the minimal modifications necessary to generate valid outputs for chiplet architectures.
-To accomplish this, we leverage a “peephole” compilation pass, which is commonly used in classical compilation to perform optimization (dragon_book).
-Peephole passes operate within a nominal peephole window and substitute its contents with a logically-equivalent replacement.
+Moreover, compilers are called to target increasingly larger quantum devices, enabled by modularity.
+To address the mounting pressure of ever-growing compilation workloads, these compilers must be cognizant of classical hardware limitations and tradeoffs between solution quality and compilation latency.
+To ameliorate these issues, we present a family of compilers optimized for modular quantum architectures that support chiplet links with non-universal basis gate sets, and leverage modular compilation strategies.
+In Section 3.1, we propose a peephole compilation pass that can be inserted into any compiler to add support for non-universal basis gate sets.
+This would allow any compiler to be compatible with chiplet-based devices without significant changes to their underlying implementations.
+Next, in Section 3.2, we introduce the stratify-elaborate quantum compiler (SEQC), which advocates for a hierarchical two-stage compilation that produces both better circuits for execution and faster compilation times.
+3.1. Chiplet-awareness via Peephole Correction
+The design of the peephole correction pass was inspired by the “peephole” optimizations commonly used in classical compilation (dragon_book).
+Peephole passes operate within a nominal peephole window and substitute its contents with a logically equivalent replacement.
 Notable examples in the context of quantum compilation include basis translation and elimination of self-adjoint gates.
-This design strategy facilitates modular compilation, not only in the sense that it can be easily parallelized, but also because peephole passes can be added between other compilation passes without breaking compatibility.
-These qualities make a peephole pass well-suited to the Strawman baseline, since we do not want to make extensive modifications to any of Qiskit’s built-in compilation stages or passes, nor introduce a significant performance overhead.
-The peephole pass used in the Strawman compiler is inserted between the routing and translation compilation stages in Qiskit.
-As shown in Figure 1, it operates on the circuit generated by SABRE (zou2024lightsabrelightweightenhancedsabre), where it detects when non-SWAP gates are placed on non-universal inter-chiplet links and employs a cyclic qubit rotation, i.e., a perfect shuffle permutation, to move the aforementioned gates to nearby intra-chiplet links ( see Figure 1(c)).
+Peephole correction (Figure 1) operates on a circuit generated by qubit allocation, where it detects when non-SWAP gates are placed on non-universal inter-chiplet links (Figure 1(b)), and employs a cyclic qubit rotation, i.e., a perfect shuffle permutation, to move the aforementioned gates to nearby intra-chiplet links (Figure 1(c)).
+Additional “return” SWAPs are then inserted to return the qubits to their initial locations (Figure 1(d)).
 Overall, for every inter-chiplet gate in the original circuit generated by SABRE, the peephole pass inserts 4 additional SWAP gates — 2 inter-chiplet and 2 intra-chiplet.
-Ultimately, this degree of overhead is inevitable when the stock, unmodified compiler is ignorant of chiplet-specific device constraints.
-To emphasize the rigor of the Strawman, it is worth discussing the designs that were ultimately discarded.
-For example, consider a hypothetical alternative design that skips the final peephole step shown in Figure 1(d).
-This would appear to save $2$ SWAPs, thus halving the cost per inter-chiplet gate.
-However, the hypothetical pass would also no longer be a true peephole, since it introduces a net qubit permutation that propagates to downstream gates in the circuit.
+Ultimately, this degree of overhead is inevitable when the stock, unmodified compiler is unaware of chiplet-specific device constraints.
+For example, skipping the return SWAP operations would initially appear to save $2$ SWAPs, thus halving the cost per inter-chiplet gate.
+However, this hypothetical design would also no longer be a true peephole, since it introduces a net qubit permutation that propagates to downstream gates in the circuit.
 From the perspective of modular compilation, this results in serial dependencies that prevent parallel execution of the pass.
-From a routing perspective, the resultant qubit permutations disrupt the solution generated by SABRE, leading to destructive routing errors that can only be resolved by an entirely new routing pass.
-Overall, such a hypothetical pass results in substantially worse output circuits compared to the peephole method, which we confirmed through testing.
-Other potential improvements on the peephole design that avoid the placement of return SWAPs would involve extensive modifications to Qiskit’s compilation stages or its underlying passes.
-In Section 4, we will explore such a modification of SABRE that accounts for the only-the-fly changes in qubit layout associated with inter-chiplet routing as an element of our design, in addition to numerous other components to optimize compilation for modular quantum architectures.
-4. The Design of SEQC
-We incorporate the hierarchical elements of chiplet architectures into our compiler design by adopting a two-stage procedure, as shown in Figure 2.
+From a routing perspective, the resultant qubit permutations disrupt the post-routing layout of qubits, leading to destructive routing errors that can only be resolved by an entirely new routing pass.
+Overall, such a hypothetical pass results in substantially worse output circuits compared to the peephole method.
+3.2. The Stratify-Elaborate Quantum Compiler
+We incorporate the hierarchical elements of chiplet architectures into SEQC by adopting a two-stage procedure (Figure 2).
 The first, stratification stage, represents a one-time overhead to elevate a quantum circuit to the chiplet level.
 The second, elaboration stage, represents a recurring cost to complete the compilation of a quantum circuit given the most up-to-date backend properties.
-4.1. Stratification Stage
-(a) A chiplet topology
+3.2.1. Stratification Stage
+(a) Example chiplet topology
 (b) Qubit-to-subcircuit mapping
@@
 Figure 3. Step-by-step example of inter-chiplet compilation in the Stratification stage.
-In this stage, we map a quantum circuit to the inter-chiplet device topology, e.g., the structure in Figure 3(a), by stratifying it into smaller, interconnected subcircuits.
-This can be accomplished with the following intermediate steps, as shown in Figures 3(b)3(d).
+In this stage, we map a quantum circuit to the inter-chiplet device topology (Figure 3(a) in our running example) by stratifying it into smaller, interconnected subcircuits.
+This can be accomplished with the following intermediate steps (Figures 3(b)3(d)).
 First, qubit-to-subcircuit mapping designates which logical qubits belong to which logical subcircuits.
 Then, as part of chiplet allocation, we define a mapping of logical subcircuits to physical chiplets on the target device and route inter-chiplet gates to conform with constraints of physical-chiplet connectivity.
-4.1.1. Qubit-to-Subcircuit Mapping
-Algorithm 1 Qubit-to-Subcircuit Mapping
+Qubit-to-Subcircuit Mapping
+Algorithm 1 Qubit-to-Subcircuit Mapping with Sim. Annealing
 […]
@@
 […]
-Qubit-to-subcircuit mapping produces a stratified quantum circuit, where logical qubits are grouped into subcircuits, as shown in Figure 3(b).
+Qubit-to-subcircuit mapping produces a stratified quantum circuit, where logical qubits (numbers in Figure 3(b)) are grouped into subcircuits (lowercase letters).
 An optimal qubit-to-subcircuit mapping minimizes the number of connections between subcircuits, thereby reducing the number of inter-chiplet SWAPs needed to transfer data between chiplets during execution.
-To perform this task, we leverage simulated annealing, an efficient stochastic optimization technique shown to work well on NP-complete problems, such as place-and-route in VLSI algorithms (annealing1989).
+To perform this task, we investigate two methods: Simulated Annealing and Spatio-Temporal Aware Qubit Allocation (STA) (ovide2024sta).
+Simulated annealing is an efficient stochastic optimization technique shown to work well on NP-complete problems, such as place-and-route in VLSI algorithms (annealing1989).
 Simulated annealing is an iterative method to optimize a state against a given cost function.
@@
 The kick function then swaps random qubits in the layout, ensuring that each SWAP is between qubits in different subcircuits.
-In doing so, we hope to move strongly interacting qubits close-together, minimizing the number of cross-chiplet gates.
-As input, our algorithm takes the input circuit, the chiplet count $N_{C}$, the chiplet size $N_{Q}$, the starting temperature $T_{0}$, the number of changes to do per degree of temperature $\alpha$, and the cooling modifier $\beta$.
-The last three are tuning parameters that do not change across runs.
-4.1.2. Chiplet Allocation
+In doing so, we hope to move strongly interacting qubits close together, minimizing the number of cross-chiplet gates.
+As input, our algorithm takes the input circuit $Circ$, the chiplet count $N_{C}$, the chiplet size $N_{Q}$, the starting temperature $T_{0}$, the number of changes to do per degree of temperature $\alpha$, and the cooling factor $\beta$.
+The last three are constant tuning parameters across runs.
+The STA algorithm is a qubit allocation algorithm designed for trapped-ion quantum systems.
+These devices often feature modular designs, where one device is composed of “traps” arranged in a (typically) line or ring topology, where each trap contains multiple qubits (ions).
+Thus, it seems possible to adapt qubit allocation algorithms designed for trapped-ion systems to support modular transmon systems as well.
+Specifically, one component of STA, denoted as Algorithm 2 in Ovide et al.
+(ovide2024sta), assigns qubits to ion traps, which we can leverage to perform qubit-to-subcircuit mapping and generate an initial subcircuit-to-chiplet mapping.
+Chiplet Allocation
 Algorithm 2 Inter-Chiplet Routing
@@
 In Figure 2, we decompose chiplet allocation into subcircuit-to-chiplet mapping and inter-chiplet routing, analogous to how qubit allocation can be decomposed into qubit layout and routing.
-Accordingly, subcircuit-to-chiplet mapping, shown in Figure 3(c), defines an initial placement of logical subcircuits to physical chiplets.
-Subsequently, inter-chiplet routing, shown in Figure 3(d), inserts inter-chiplet SWAPs to account for limited cross-chiplet gate connectivity.
-In our design, we do not explictly derive a subcircuit-to-chiplet mapping.
-Rather, we extend and modify the SABRE algorithm (sabre_swap) to apply to chiplets, thereby using iterations of inter-chiplet routing with randomized permutations to inform a subcircuit-to-chiplet mapping.
+Accordingly, subcircuit-to-chiplet mapping (Figure 3(c)), defines an initial placement of logical subcircuits (lowercase letters) to physical chiplets (uppercase letters).
+Subsequently, inter-chiplet routing (Figure 3(d)) inserts inter-chiplet SWAPs to account for limited cross-chiplet gate connectivity.
+In our design, we use our qubit-to-subcircuit mapping to provide an initial subcircuit-to-chiplet mapping.
+Then, we extend and modify the SABRE algorithm (sabre_swap) to apply to chiplets, placing iterations of inter-chiplet routing alongside randomized qubit permutations to inform a final subcircuit-to-chiplet mapping.
+We did explore a design that extended other components of STA to perform subcircuit-to-chiplet mapping on transmon chiplet systems, but did not produce circuits meaningfully better than this iterative approach and increased compilation time, hence it was abandoned.
 Inter-chiplet routing seeks to consolidate the qubits of (non-SWAP) inter-chiplet gates to a single chiplet by inserting inter-chiplet SWAPs into the stratified circuit.
-For an single two-qubit, inter-chiplet gate, we represent inter-chiplet routing as a shortest-path problem, where the two qubits of the gate correspond to chiplet endpoints in the unweighted inter-chiplet connectivity graph.
+For a single two-qubit, inter-chiplet gate, we represent inter-chiplet routing as a shortest-path problem, where the two qubits of the gate correspond to chiplet endpoints in the unweighted inter-chiplet connectivity graph.
 From the shortest path, we can derive an optimal sequence of SWAPs for routing, where the unweighted path length, or chiplet distance, indicates the length of the SWAP chain.
-However, in the context of a entire stratified circuit, composed of multiple inter-chiplet gates, it becomes possible to interleave otherwise efficiently reorder execution of SWAP chains to reduce the depth and/or total number of inter-chiplet SWAPs.
+However, in the context of a stratified circuit, composed of multiple inter-chiplet gates, it becomes possible to interleave SWAP chains (i.e., efficiently reorder their execution) to reduce the depth and/or total number of inter-chiplet SWAPs.
 Like qubit routing, we would expect it to be impractical to derive an optimal solution for this more complex picture of inter-chiplet routing at compile-time.
@@
 These SWAPs typically occur when an idle qubit can be used as an ancilla qubit to facilitate the routing of an inter-chiplet gate without affecting other inter-chiplet gates.
-In certain device topologies, a SWAP acting on the qubits of two inter-chiplet gates can also exhibit commercialistic behavior.
+In certain device topologies, a SWAP acting on the qubits of two inter-chiplet gates can also exhibit commensalistic behavior.
 Finally, parasitic SWAPs act on the qubits of two inter-chiplet gates, where one gate benefits while the other is harmed.
-Since this variety of SWAP is actively counterproductive towards a gate, it should be, and likely will be, avoided in most circumstances.
+Since this variety of SWAPs is actively counterproductive towards a gate, it should be, and likely will be, avoided in most circumstances.
 However, given a particularly onerous circuit, it may be necessary in such rare instances to break out of a deadlock.
 For example, clearing one gate may free physical qubits to help route others.
-4.2. Elaboration Stage
+3.2.2. Elaboration Stage
 (a) An intra-chiplet topology
-(b) A $3$-qubit subcircuit post-stratification.
+(b) A $3$-qubit subcircuit post-stratification
 (c) Subcircuit after intra-chiplet layout
@@
 To accomplish these tasks, we divide the elaboration stage into two substages: qubit allocation and parallel basis translation and optimization.
-Figure 4 illustrates this with an example of elaboration for a single chiplet in Figure 4(a) and its corresponding subcircuit in Figure 4(b).
-4.2.1. Qubit Allocation
-Like other qubit allocation schemes, we incorporate layout (see Figure 4(c)) and routing (see Figure 4(e)) substages.
-However, the presence of inter-chiplet gates also necessitates several key changes to the SABRE (sabre_swap) qubit allocation algorithm to achieve correctness and/or maximum parallelism.
-First, the placement of inter-chiplet gates is a non-parallelizable task (due to cross-chiplet interactions) that strongly influences the layout and routing passes, which could otherwise be fully parallelized across chiplets.
-To account for this, we introduce a serial stage, as shown in Figure 4(d), between the parallel layout and routing passes that greedily places gates on the nearest valid inter-chiplet edge.
+Figure 4 illustrates this with an example of elaboration for a single chiplet (Figure 4(a)) and its corresponding subcircuit (Figure 4(b)).
+Qubit Allocation
+We employ a modified version of SABRE to perform qubit allocation in the elaboration stage.
+Specifically, we parallelize the SABRE routing algorithm and add constraints to maintain correctness when applied to a stratified quantum circuit with inter-chiplet gates.
+We also explored a layout pass that used portions of the STA algorithm, but the marginal benefits to solution quality were not worth the 56×5-6$\times$5 - 6 × slower elaboration time.
+The placement of inter-chiplet gates is a non-parallelizable task (due to cross-chiplet interactions) that strongly influences the layout and routing passes, which could otherwise be fully parallelized across chiplets.
+To account for this, we introduce a serial stage (Figure 4(d)) between the parallel layout and routing passes that greedily places gates on the nearest valid inter-chiplet edge.
 A greedy policy was chosen because inter-chiplet gates have more restrictions on valid placements compared to intra-chiplet gates, while also experiencing substantially higher error rates (smith_2022_chiplets).
 Once the positions of the inter-chiplet gates are settled, they must remain fixed, imposing an additional constraint on qubit routing.
-To promote this behavior in SABRE, we extended its cost heuristic to weigh the device topology by fidelity.
+To promote this behavior in SABRE, we extend its cost heuristic to weigh the device topology by fidelity.
 This causes higher error links to assume higher costs, discouraging SABRE from moving the high-error inter-chiplet gates.
@@
 Thus, we need to impose hard restrictions to ensure correctness.
-We establish two disjoint gates sets.
+We address this by establishing two disjoint gate sets.
 One gate set operates strictly within a chiplet.
@@
 By restricting each gate set to its own disjoint part of the overall device topology, we ensure correctness across all cases.
-4.2.2. Parallel Translation and Optimization
-The remaining compilation stages of basis translation and circuit optimization can be easily parallelized at the chiplet granularity, as shown in Figure 4(f).
-Basis translation, in particular, is embarrassingly parallel, because every gate can be translated independently from every other gate.
-Meanwhile, as discussed in Section 2, optimization algorithms already employ partitioning and parallelization to create tractable subproblems from intractably large circuits, facilitating a smaller problem space, lower memory usage, and faster performance.
+Parallel Translation and Optimization
+The remaining compilation stages of basis translation and circuit optimization can be easily parallelized at the chiplet granularity (Figure 4(f)).
+Basis translation, in particular, is embarrassingly parallel: gates can be translated independently from each other.
+Meanwhile, as discussed in §2, optimization algorithms already employ partitioning and parallelization to create tractable subproblems from intractably large circuits, facilitating a smaller problem space, lower memory usage, and faster performance.
 In modular quantum architectures, we physically motivate partitions through the natural chiplet boundaries, enabling us to reap the same resource and performance benefits of partitioning without compromising on solution quality.
 4. Methodology
-To experimentally validate our compiler for modular quantum architectures, we ran a benchmark suite on a wide spectrum of circuit sizes for both SEQC and the Strawman baseline.
-Both SEQC and the baseline compiler were implemented in the Qiskit SDK (qiskit2024) v1.2.41.2.41.2.41.2.4 and Python $3.12.6$.
-Elaboration time results were taken on a machine with a $128$-core Ampere Altra Max ARM64 processor and $256$ GB of DDR4 RAM.
-When given a parallelizable workload, SEQC attempts to allocate its tasks evenly among the available physical cores, rounded to the neared whole number.
-Finally, the simulated annealing technique used in the qubit-to-subcircuit stage uses a starting temperature $T_{0}=200$, a rate of change of $0.005T$ each round, $\frac{T}{50}$ permutations per round, and a total of $10$ trials per CPU core.
-5.1. Metrics
-We compare SEQC to the Strawman baseline over a wide range of common and/or relevant metrics, discussed below.
+We evaluate quantum compilers based on the quality of their solutions, their speed of compilation, and their ability to scale to large circuits/devices.
+To experimentally profile these characteristics, we benchmark a compiler on a suite of quantum circuits, detailed in §4.1, targeting a mock chiplet-based quantum backend, as specified in §4.2.
+In total, we evaluate four compilers—a reference compiler, a peephole-augmented compiler, and two versions of SEQC—on simulated backends ranging from 1 to 100 chiplets.
+All compilers are implemented in the Qiskit SDK (qiskit2024) v1.2.41.2.41.2.41.2.4 and Python $3.12.6$, and timing results are taken on a machine with a $128$-core Ampere Altra Max ARM64 processor and $256$ GB of DDR4 RAM.
+For our reference compiler, we use the one built into Qiskit, parameterized with optimization_level=3.
+Because Qiskit is not natively chiplet-aware, we run these experiments on a backend with universal inter-chiplet links (§4.2).
+To construct the peephole compiler we augment the reference Qiskit compiler by inserting our peephole pass between its routing and basis_translation stages, thus making it chiplet-aware.
+Our two SEQC versions differ in the technique they use for qubit-to-subcircuit mapping.
+Our first SEQC implementation leverages simulated annealing, initialized with a starting temperature $T_{0}=200$, a rate of change of $0.005T$ per round, $\frac{T}{50}$ permutations per round, and a total of $10$ trials per CPU core.
+Meanwhile, our second SEQC implementation leverages STA for subcircuit-to-qubit mapping and subcircuit-to-chiplet mapping.
 We aggregate results in a manner similar to the SPEC family of benchmarks (spec2017), by taking the geometric mean of the relative performance ratio against a baseline technique across a benchmark suite of quantum circuits.
 The resultant score is then derived for each of the evaluated metrics on multiple backends of varying numbers of chiplets.
-Fidelity.
-The top priority for quantum circuit execution is to maximize fidelity.
-Without a guarantee of correctness, the outputs of the quantum circuit are useless, and any speedup or quantum advantage is rendered moot.
-Fidelity is impacted by the decoherence constraints of individual qubits, as well as the errors on a gate-by-gate basis (mike_and_ike).
-Due to simulation and hardware limitations, we utilize the estimated success probability (ESP) metric (esp) to approximate fidelity.
-Specifically, we track the per-qubit ESP and report an aggregate ESP for each benchmark from the geometric mean of the accumulated measurements.
-Non-recurring stratification time.
-This represents the time taken in the stratification stage, which will be present for our compiler and absent for the baseline.
-The stratification time represents a one-time overhead to prepare a quantum circuit for execution on a modular architecture.
-This result can be reused for future executions on any device with an equivalent or more expansive topology.
-Recurring elaboration time.
-The error characteristics of a quantum device can change on a hour-by-hour basis, or even a run-to-run basis (Tannu_2019_ErrorVariability; Murali_2019_NoiseMapping; dasgupta2021stability).
-Thus, to generate the highest-quality circuits with the most up-to-date characteristics possible, a given quantum circuit should be recompiled for every execution.
-Since the qubit and chiplet topologies remain static per-device, it is possible under the SEQC framework to reuse the results from the stratification stage, and only incur overhead from repeating the elaboration stage.
-Solve time.
-This metric is relevant only to the baseline Strawman chiplet compiler, and represents the time it takes to compile a quantum circuit.
-The baseline compiler solve time is a recurring cost, i.e., this compilation step has to be performed every time a circuit needs to execute on a quantum machine.
-Estimated execution time.
-The per-shot execution time of a quantum circuit is determined by the critical path of gate operations weighted by gate duration.
-In addition to performance concerns, execution time indirectly affects fidelity on NISQ-era devices due to decoherence.
-In other words, it is imperative that quantum circuits execute within the decoherence time to ensure high-fidelity outcomes.
-Number of inter-chiplet gates.
-The inter-chiplet gates of a chiplet architecture contribute the most to error and execution time relative to other gates (at least by a factor of 4×4$\times$4 × (smith_2022_chiplets)).
-As a result, the number of inter-chiplet gates should represent a strong proxy metric for the quality of a compiler solution.
-Accordingly, minimizing the number of inter-chiplet gates should result in higher fidelity, faster-executing quantum circuits.
-Circuit depth.
-Defined as the unweighted critical path of a quantum circuit (mike_and_ike), the circuit depth is often employed as a backend-agnostic metric for execution time that can be measured statically.
-Notably, it is used in SABRE (sabre_swap) as part of its cost heuristic.
-Gate count.
-Similar to circuit depth, the total number of gates in a circuit can also be used as a static, backend-agnostic metric for certain properties of the aforementioned circuit.
-For instance, it may provide an indication of the amount of accumulated error, or, like circuit depth, the execution time.
-SABRE (sabre_swap) uses gate count to decide between randomly seeded allocation trials.
-5.2. Benchmarks
+4.1. Benchmark Suite
 (a) BitCode
@@
 (e) Hamiltonian Simulation
-Figure 5. Examples of the Benchmark Circuits
-For our benchmark suite, we leverage a subset of circuits from the Supermarq (supermarq) suite.
-The circuits are chosen to be representative of practical, real-world applications for quantum computing, while also being feasible to compile for a large number of qubits.
-In particular, our benchmark suite comprises the following quantum circuits.
+Figure 5. Examples of benchmark applications for 5-qubit circuits.
+We leverage circuits from Supermarq (supermarq) that are representative of practical, real-world applications for quantum computing, while also being feasible to compile for a large number of qubits.
+In particular, our suite comprises the following quantum circuits.
 […]
@@
 The GHZ state represents a maximally entangled $n$-qubit state of $\frac{1}{\sqrt{2}}\left(\ket{|}0>^{\otimes n}+\ket{|}1>^{\otimes n}\right)$.
-The algorithmic circuit structure is essentially a single chain of CNOT gates, as shown in Figure 5(c) From a compilation perspective, this represents a pure exercise in qubit allocation, because the logical circuit is entirely serial.
+The algorithmic circuit structure is essentially a single chain of CNOT gates (Figure 5(c)) From a compilation perspective, this represents a pure exercise in qubit allocation, because the logical circuit is entirely serial.
 An $n$-qubit GHZ circuit that fully utilizes a $k$-chiplet device should necessitate $\lfloor\frac{n}{k}\rfloor-1$ logical CNOT operations to extend between chiplets.
@@
 The variational quantum eigensolver (VQE) (vqe) is another hybrid quantum-classical variational algorithm that is often used in chemistry applications.
-In our experiments, we choose a $2$-layer ansatz whose parameters are again randomly generated, as shown in Figure 5(d).
-This results in a quantum circuit that is essentially two CNOT chains separated by single-qubit operations.
+In our experiments, we choose a $2$-layer ansatz whose parameters are again randomly generated (Figure 5(d)).
+The resultant circuit structure features two CNOT chains separated by single-qubit operations.
 The CNOT chains can be interleaved, which in conjunction with the single-qubit gates, creates many opportunities for parallelism (resembling pipelining in classical computing).
@@
 The Supermarq benchmark selects the 1D Transverse Field Ising Model (TFIM) system to model, which is relevant for phase transitions in magnetic materials.
-The circuit for Hamiltonian Simulation, shown in Figure 5(e), resembles a slightly more complex GHZ circuit, including how it is completely serial.
+The circuit for Hamiltonian Simulation (Figure 5(e)) resembles a slightly more complex GHZ circuit, including how it is completely serial.
 Compared to GHZ, this circuit should test the compiler’s ability to lookahead during routing to detect back-to-back gates on the same two-qubit pair, while also emphasizing parallelism originating from the efficient placement of SWAP gates.
@@
 Frequency (backend_specs) $6\times 10^{9}$ Hz
-(a) Qubit Properties.
+(a) Qubit Properties
 Instruction Duration (ns) Error
-(backend_specs) $25$ $0.109\%$
-SX (backend_specs) $25$ $0.109\%$
-$\text{R}_{\text{z}}(\phi)$ (backend_specs) 00 $0.00\%$
-CZ (intra-chiplet) (backend_specs) $34$ $0.605\%$
+(backend_specs) $25.0$ $0.109\%$
+SX (backend_specs) $25.0$ $0.109\%$
+$\text{R}_{\text{z}}(\phi)$ (backend_specs) $0.0$ $0.000\%$
+CZ (intra-chiplet) (backend_specs) $34.0$ $0.605\%$
+CZ (inter-chiplet) $126.0$ $2.420\%$
 SWAP (inter-chiplet) $702.4$ $10.23\%$
-Reset (backend_specs) $500$ $0.186\%$
-Measure (backend_specs) $500$ $0.196\%$
-(b) Instruction Properties.
+Reset (backend_specs) $500.0$ $0.186\%$
+Measure (backend_specs) $500.0$ $0.196\%$
+(b) Instruction Properties
+Universal gate set only. Non-universal gate set only.
 Table 1. Specifications of Simulated Chiplet Backend.
-The topology of the simulated device is composed of a grid lattice of symmetric chiplet modules, where each chiplet features a heavy-hexagon qubit topology (heavy_hex), as shown in Figure 6.
+The simulated device has a grid lattice topology of symmetric chiplet modules, where each chiplet features a heavy-hexagon qubit topology (heavy_hex) (Figure 6).
 The intra-chiplet topology is constructed such that the heavy-hex structure is preserved for the global device topology (ignoring chiplet division).
@@
 5. Experimental Results
-Figures 79 present measurements of the recurring compilation time (i.e., solve time) of the baseline Strawman chiplet compiler, and SEQC’s two compilation stages: the non-recurring stratification stage and the recurring elaboration stage.
-Figure 7. SEQC stratification time normalized to the 20-qubit (i.e., 2-chiplet) trial.
-The average stratification time of SEQC across the benchmark suite obeys a quadratic trajectory ($R^{2}=0.9971$) with respect to the circuit size and number of chiplets, as shown in Figure 7.
-We note that the compilation time of the baseline, presented in Figure 8, also scales quadratically ($R^{2}=0.9964$).
-Although stratification time represents a one-time cost, the quadratic scaling of stratification points toward the practicality of SEQC.
-The stratification time depends heavily on the complexity of the input circuit.
-Simple circuits, such as GHZ or HamiltonianSimulation, exhibit a very gradual increase in stratification time as circuit size increases, while others, such as BitCode and PhaseCode, are much steeper.
-This observation holds true for the solve time of the baseline compiler as well.
-Collectively, these results indicate that the computational complexity of the baseline compiler and the stratification stage of SEQC are affected similarly by the same quantum circuit characteristics.
-Figure 8. Baseline compilation time normalized to the 20-qubit (i.e., 2-chiplet) trial.
-Figure 9. Relative elaboration time of SEQC normalized to the Strawman baseline for devices with 10-qubit chiplets (lower is better).
-From Figure 9, we observe an improved SEQC elaboration time compared to the baseline solve time, leading to a roughly 24×2-4$\times$2 - 4 × speedup, up until $800$-qubit (i.e., $80$-chiplet) trials.
-The observed speedup reflects the expected performance improvement that our stratify-elaborate technique should garner from parallelization and problem-space reduction.
-For $81$-chiplet experiments and above, our results begin to reflect a far more marginal speedup compared to baseline, reflecting the limitations of our evaluation hardware.
-Namely, as the number of chiplets begins to approach the number of physical cores and hardware threads on our hardware, SEQC experiences difficulties in evenly distributing its tasks.
-For example, under-allocating tasks causes poor core utilization, while over-allocation leads to process thrashing, either way resulting in poorer elaboration time.
-However, in all cases, SEQC outperforms the baseline on average.
-On a circuit-by-circuit basis, we notice higher relative performance with parallel workloads (like BitCode and PhaseCode) compared to serial workloads (GHZ and HamiltonianSimulation).
-Notably, above the $81$-chiplet threshold where the classical machine we use for compilation no longer suffices, elaboration time exceeds the baseline solve time for the serial workloads.
-This is because serial circuits are fundamentally trivial to solve, and thus benefit less from the parallelism offered by SEQC.
-We verified experimentally that in these cases, SEQC spends almost all of its time on thread startup and teardown, rather than useful work.
-Moreover, the raw compilation time is trivially small in these cases, owning to the simplicity of the circuits, thus relative performance differences have little practical significance.
-Figure 10. Relative estimated fidelity (ESP) of SEQC normalized to the Strawman baseline for devices with 10-qubit chiplets (higher is better).
-Figure 10 presents the relative estimated fidelity of SEQC compared to the baseline with respect to the number of qubits per circuit.
-Notably, as the number of chiplets / the circuit size increases, SEQC improves its relative fidelity against the baseline.
-We begin to see higher fidelity circuits in the geometric mean once the circuit size exceeds $36$ chiplets, up to a maximum of $36\%$ higher fidelity.
-All benchmarks reflect higher relative performance for the $50$-chiplet trials and above.
-Figure 11. Estimated per-shot execution time of circuits produced by SEQC normalized to circuits produced by the Strawman baseline for devices with 10-qubit chiplets (lower is better).
-Figure 11 presents the estimated per-shot execution time of the circuits produces by SEQC compared to the circuits produced by the baseline compiler, as a function of the number of qubits per circuit.
-As explained above, we approximate the duration of an inter-chiplet SWAP as 4×4$\times$4 × the duration of an intra-chiplet SWAP (smith_2022_chiplets).
-Under this model, we observe execution time benefits up to a 1.92×1.92$\times$1.92 × improvement over baseline at $750$ qubits ($75$ chiplets).
-Not only are improvements in execution time desirable for quantum circuit throughput, they could also result in lower decoherence and thus improved circuit fidelity.
-Figure 12. Relative number of inter-chiplet gates for SEQC normalized to the Strawman baseline for devices with 10-qubit chiplets (lower is better).
-Figure 12 shows the number of inter-chiplet swaps in the circuits generated by SEQC relative to the circuits generated by the baseline, with respect to the number of qubits per circuit.
-For essentially all tested circuits, SEQC achieves a significant reduction in the number of inter-chiplet swap gates, appearing to asymptotically converge on a roughly 4×4$\times$4 × reduction, with the greatest reduction of 4.6×4.6$\times$4.6 × at $800$ qubits / $80$ chiplets.
-This aligns with our model of inter- vs intra-chiplet gates, where inter-chiplet SWAPs are assumed to have 4×4$\times$4 × the error (smith_2022_chiplets) and 4×4$\times$4 × the gate duration.
-Figure 13. Relative circuit depth of SEQC normalized to the Strawman baseline for devices with 10-qubit chiplets (lower is better).
-Figure 14. Relative gate count of SEQC normalized to the Strawman baseline for devices with 10-qubit chiplets (lower is better).
-Relative circuit depth and gate count, shown in Figures 13 and 14, respectively, share similar behavior and analysis.
-At small circuit sizes, SEQC has higher circuit depth and gate count than the baseline.
-The gap shrinks as the circuit size increases, although SEQC always tends to perform worse on these metrics on average.
-This behavior is to be expected, since the baseline directly optimizes for circuit depth and gate count, while SEQC prioritizes other metrics, e.g., reducing inter-chiplet gates.
-6.1. Discussion
-The discrepancy between the circuit depth and gate count results compared to the execution time and fidelity results, respectively, are indicative of the differences between SEQC and the contemporary quantum compilers.
-Circuit depth and gate count represent the heuristics used in the baseline, especially SABRE (sabre_swap), and they are fundamentally unweighted metrics.
-In other words, circuit depth and gate count only consider the number of gates (in the critical path and in total) irrespective of each one’s unique underlying properties.
-Meanwhile, SEQC is able to more directly estimate execution time and fidelity, the practically relevant metrics, by constructing weighted heuristics.
-With the context of our static metrics (number of inter-chiplet gates, circuit depth, and gate count), it is possible to justify the improvements in relative fidelity and relative per-shot execution time as the circuit size increases.
-Namely, as expected, the unweighted heuristics in SABRE tend to under-represent the error and gate delay contribution of inter-chiplet gates.
-This effect gets exacerbated for wider or deeper circuits, which increase the number of opportunities for cross-chiplet interactions.
+We present results for the Qiskit reference compiler, the peephole-augmented Qiskit (PA-Qiskit) compiler, the SEQC compiler using simulated annealing (SEQC-AN), and the SEQC compiler using STA (SEQC-STA).
+The Qiskit-based and SEQC compilers prioritize different heuristics in compilation.
+We examine how well each compiler is able to minimize its preferred heuristics, and how well that translates to practical solution quality.
+Finally, we analyze compilation time, specifically the recurring and non-recurring compilation times for repeating compilations of a single application.
+Number of inter-chiplet gates
+Figure 7. Relative number of inter-chiplet gates for SEQC normalized to the Qiskit baseline (lower is better).
+The primary heuristic we minimize in SEQC is the number of inter-chiplet gates in a quantum circuit.
+This is because we expect that the inter-chiplet gates of a chiplet architecture contribute the most to error and execution time relative to other gates (at least by a factor of 4×4$\times$4 × (smith_2022_chiplets)).
+As a result, the number of inter-chiplet gates should represent a strong proxy metric for the quality of a compiler solution, and thus minimizing the number of inter-chiplet gates should result in higher-fidelity, faster-executing quantum circuits.
+Accordingly, SEQC consistently produces results that have significantly fewer inter-chiplet gates compared to Qiskit-based compilers (Figure 7).
+Specifically, compilation with SEQC-STA results in $82.43\%$ fewer inter-chiplet gates on average compared to Qiskit (up to $92.94\%$).
+Compared to PA-Qiskit, compilation with SEQC-STA results in $64.92\%$ fewer inter-chiplet gates on average, up to $78.62\%$.
+It is worth mentioning that the inter-chiplet gates in the universal backend would be CZ gates, while the non-universal backend would be SWAP gates.
+For every inter-chiplet (CZ) gate in Qiskit, PA-Qiskit should have 2 inter-chiplet (SWAP) gates.
+However, we observe PA-Qiskit produces circuits with a roughly comparable number of inter-chiplet gates to Qiskit.
+In fact, as size increases, PA-Qiskit tends to result in fewer inter-chiplet gates than Qiskit.
+This outcome may indicate the presence of optimizations that are removing extraneous SWAPs.
+Gate count and circuit depth.
+Figure 8. Relative gate count of SEQC normalized to the Qiskit baseline for devices (lower is better).
+Figure 9. Relative circuit depth of SEQC normalized to the Qiskit baseline for devices (lower is better).
+Gate count and circuit depth are static, backend-agnostic metrics that can serve as indicators for a quantum circuit’s expected fidelity and/or execution time, where circuit depth is the critical path of a quantum circuit in terms of individual gates (mike_and_ike).
+SABRE (zou2024lightsabrelightweightenhancedsabre; sabre_swap), moreover, explicitly minimizes gate count by incorporating it into its cost functions.
+We observe that SEQC typically produces results with higher gate count and circuit depth compared to Qiskit and PA-Qiskit (Figures 8 and 9).
+This is due to SEQC prioritizing the reduction of the number of inter-chiplet gates over total gate count or circuit depth.
+Nevertheless, SEQC-STA, in particular, is able to demonstrate competitive results in these metrics for larger circuits.
+In terms of relative circuit depth, SEQC-STA has comparable or slightly better results compared to PA-Qiskit at $500$ qubits (50 chiplets) and above.
+In terms of relative gate count, SEQC-STA has comparable or slightly better results compared to PA-Qiskit at $400$ qubits (40 chiplets) and above.
+For the $810$ and $1000$ qubit ($81$ and $100$ chiplet) trials, it achieves $14.11\%$ and $8.87\%$ gate count reduction over Qiskit, respectively, despite the latter using universal inter-chiplet links.
+Fidelity
+Figure 10. Relative estimated fidelity (ESP) of our proposed compilers normalized to the Qiskit baseline (higher is better).
+Practical applications of quantum computing require some guarantee of program output correctness (measurements).
+Thus, the highest priority for quantum circuit compilation is to maximize fidelity.
+Fidelity is impacted by the decoherence constraints of individual qubits, as well as the errors on a gate-by-gate basis (mike_and_ike).
+Due to hardware limitations, we were unable to estimate fidelity though simulation.
+Instead, we leverage estimated success probability (ESP) (esp) to approximate fidelity, which can be computed statically.
+Specifically, we track the per-qubit ESP of a quantum circuit and report an averaged ESP for each benchmark from the geometric mean of the circuit measurements.
+Figure 10 presents the relative estimated fidelity of each of our compilers compared to the Qiskit compiler with respect to circuit size.
+There is a noticeable benefit in fidelity associated with the SEQC compiler, particularly with SEQC-STA.
+We observe an average improvement of $22.13\%$ over PA-Qiskit, up to a $62.87\%$ improvement.
+Moreover, it is important to mention that the Qiskit trial was ran on a backend with universal inter-chiplet links, which should translate to higher fidelity solutions.
+Nevertheless, SEQC-STA demonstrates an average improvement of $9.30\%$, up to a $49.99\%$ improvement.
+Estimated execution time.
+Figure 11. Estimated per-shot execution time of circuits produced by SEQC normalized to circuits produced by the Qiskit baseline (lower is better).
+The per-shot execution time of a quantum circuit is determined by the critical path of gate operations weighted by gate duration.
+In addition to performance concerns, execution time indirectly affects fidelity on NISQ-era devices due to decoherence.
+In other words, it is imperative that quantum circuits execute within the decoherence time to ensure high-fidelity outcomes.
+Figure 11 presents the estimated per-shot execution time of our proposed compilers relative to the reference as a function of the number of qubits per circuit.
+As explained previously, we approximate the duration of an inter-chiplet SWAP as 4×4$\times$4 × the duration of an intra-chiplet SWAP (smith_2022_chiplets).
+Under this model, we observe that the SEQC compilers consistently outperform PA-Qiskit in execution by $24.19\%$ on average ($63.98\%$ max).
+It is notable that, despite the gate set affordances given to the Qiskit compiler, SEQC-STA is able to achieve parity in execution time starting from $600$ qubits ($60$ chiplets).
+For $810$ qubits ($81$ chiplets) and above, SEQC-STA actually compiles faster-executing quantum circuits.
+Compilation time
+Figure 12. Compilation time of our evaluated compilers (lower is better).
+Figure 13. Relative recurring compilation time of our proposed compilers normalized to the Qiskit baseline (lower is better).
+Near-term quantum devices feature rapidly fluctuating error characteristics that can change hourly, or even on a run-to-run basis (Tannu_2019_ErrorVariability; Murali_2019_NoiseMapping; dasgupta2021stability).
+Thus, across the entire execution lifetime of an application (quantum circuit), regular re-compilation is necessary to generate the highest-quality circuits with the most up-to-date characteristics possible for every execution.
+Traditional quantum compilers, like Qiskit and PA-Qiskit, approach each compilation tabula rasa, where each new compilation restarts from the original algorithmic circuit.
+It is challenging to speed up repeated recompilations without fundamental changes to their design, as noise-aware routing and synthesis algorithms, which leverage updated error characteristics, usually function from a device-agnostic starting point.
+In contrast, SEQC’s stratification stage enables stateful retention of previous compilations.
+It is agnostic to the ever-changing error characteristics of a quantum device, and instead only depends on the device’s static topology.
+Therefore, a stratified quantum circuit can be perennially reused over multiple rounds of compilation, and thus SEQC only repeats the elaboration stage.
+Nevertheless, although stratification is a one-time cost for the SEQC compiler, it must be scalable in order to be practical.
+In Figure 12, we plot the non-recurring compilation time of our evaluated compilers.
+For our SEQC compilers, we plot stratification time and stratification + elaboration time.
+These can be compared to the total solve time for the Qiskit and PA-Qiskit compilers.
+We demonstrate that all compilation times obey a quadratic trajectory ($R^{2}\geq 0.9928$).
+Figure 13 shows the recurring compilation time.
+Specifically, this refers to the elaboration time of the SEQC compilers and the full compilation time for Qiskit and PA-Qiskit.
+Overall, SEQC compilers have a large speedup over Qiskit-based compilers, owing to the use of parallelism and modular compilation strategies.
+Compared to Qiskit, SEQC-STA achieves 3.27×3.27$\times$3.27 × speedup on average (up to 6.74×6.74$\times$6.74 ×), and compared to PA-Qiskit 4.69×4.69$\times$4.69 × average speedup (6.61×6.61$\times$6.61 × max).
+6. Discussion
+These results demonstrate a flaw with gate count and circuit depth as heuristics: their lack of noise-awareness.
+Circuit depth and gate count only consider the number of gates (in the critical path and in total) irrespective of each one’s unique underlying properties.
+It is expected that two-qubit gates will be significantly more noisy than single-qubit gates, and inter-chiplet gates should be significantly more noisy than intra-chiplet gates (smith_2022_chiplets).
+Failure to account for duration and error differences between gates or types of gates in quantum compilation heuristics will inevitably result in sub-optimal solutions when the discrepancies are extreme, e.g., in chiplet architectures.
+Conversely, the hierarchical structure of the SEQC framework with the stratification and elaboration stages imparts some degree of implicit noise-awareness, where minimizing the number of inter-chiplet gates takes precedence over intra-chiplet/total gates.
+As a result, SEQC can better optimize for the practically significant metrics of circuit fidelity and circuit execution time.
 Overall, we observe that SEQC yields circuits with more desirable characteristics when executing on modular quantum architectures, while at the same time it achieves faster compilation times.
-Also, as the number of qubits increase, the performance improvement of SEQC widens compared to the baseline, especially for the most important metrics of fidelity, compiled circuit execution time, and number of inter-chiplet gates.
+Also, as the number of qubits increases, the performance improvement of SEQC widens compared to the baseline, especially for the most important metrics of fidelity, compiled circuit execution time, and number of inter-chiplet gates.
 These trends hold promise in allowing for faster and better compilation for future large-scale modular quantum devices than we can achieve with today’s frameworks.
 7. Related Work
-Even though quantum computing is still nascent, there is a significant body of work on techniques for compiling quantum circuits for NISQ processors.
-Here, we focus on the work that best relates to our contributions.
 MECH (mech) sacrifices some ancillary qubit resources to build an ”adjustable multi-entry communication highway” through software.
-By entangling ancillary qubits and distributing them throughout the chiplet architecture’s topology, the work virtualizes those qubits into fast network-on-chip routing channels without hardware modifications.
-Future work to extend SEQC with MECH’s methodology may enable combining MECH’s techniques with SEQC’s allocation procedures to gain the benefits of each.
-Noting the scaling challenges and fidelity issues associated with large, monolithic quantum machines, CutQC (cutqc) establishes a hybrid computing paradigm which enables smaller-scale quantum computers to evaluate quantum circuits they would have otherwise had too few qubits to execute.
-To establish said paradigm, the work introduces a novel circuit-cutting technique for partitioning larger quantum circuits into small subcircuits.
-Classical post-processing is performed on the results of the subcircuit evaluations, merging them to find the result that would be generated by the uncut circuit.
-Circuit cutting as performed in CutQC served as an inspiration for SEQC’s stratification step, but unlike CutQC, SEQC does not perform any classical “stitching” of the results afterwards.
-It is interesting to note that our first version of SEQC was formulating the problem using constrained optimization as well, but the slowdown of the stratification step was unacceptable and impractical for large quantum circuits, and thus was abandoned in favor of the heuristics we currently employ in our compiler.
-Bandic et al.(bandic2024profiling) also note the benefits of modularity.
-Their study is motivated by application-specific optimization – they seek techniques that allow a quantum circuit’s structure to influence hardware architecture.
-They define the relationship between circuit structure and mapping efficiency, utilizing this information to optimize the architecture of single- and multi-core quantum processors.
-Piveteau and Sutter (circuit_knitting) investigate the benefits of using quasiprobability simulation to allow subcircuits cut from larger quantum circuits to perform their quantum operations independently, simulating inter-chiplet link behavior where the subcircuits connect.
-When augmented with classical communication between circuits to reduce the simulation overhead incurred by sampling, they demonstrate that resulting circuits reach fixed accuracy with less simulation time.
-Unlike SEQC that targets quantum compilation, this technique has been developed to partition large quantum circuits into subcircuits that fit on smaller devices, at the cost of a simulation overhead to classically combine them.
+By entangling ancillary qubits and distributing them throughout a chiplet, MECH virtualizes those qubits into fast network-on-chip routing channels.
 Lin et al.
-(let_every_qubit) recognize that each 2-qubit gate in a quantum circuit has a dynamic behavioral profile which depends on manufacturing and environmental factors, and present a compilation technique which uses each 2-qubit set’s calibration data to determine potentially nonstandard 2Q gate operations which are much more ideal for that set to perform than their standard counterparts—essentially, choosing a more favorable basis.
-By characterizing these nonstandard gates, using their unitaries to synthesize standard gates, and essentially allowing 2Q gates to perform standard operations along bases more compatible with their respective physical states, the technique improves fidelity.
-This technique is orthogonal to SEQC and could work synergistically with it.
-By mapping circuit partitioning to a minimum cut problem, Bandic et al.
-(bandic2023mapping) use the QUBO model with quantum solvers to generate a set of cuts intended to minimize swaps between circuit partitions.
+(let_every_qubit) present a compilation technique that utilizes calibration data to determine potentially nonstandard 2Q gates that can perform standard operations along bases more compatible with their respective physical states and improve fidelity.
+Bandic et al.(bandic2024profiling) utilize the relationship between quantum circuit structure and mapping efficiency to optimize the architecture of monolithic and modular quantum processors, on the path to application-specific hardware optimization.
+Escofet et al.
+(escofet2024revisiting) develop a characterization technique for arbitrary quantum circuits which yields theoretical upper and lower optimization bounds, enabling the optimality assessment of qubit mapping solutions.
+All these techniques are orthogonal to SEQC, and could work synergistically with it to achieve the combined benefits.
+CutQC (cutqc) enables smaller-scale quantum computers to execute quantum circuits that are otherwise too large to fit on the device.
+It partitions large quantum circuits into smaller subcircuits, executes them one at a time, and then merge the individual results at the cost of classical post-processing.
+Piveteau and Sutter (circuit_knitting) similarly allow subcircuits to execute independently, and use quasiprobability simulation to simulate inter-chiplet link behavior where the subcircuits connect, reaching fixed accuracy with less simulation time.
+SEQC’s stratification stage was inspired by CutQC, but unlike these works SEQC does not perform any classical “stitching” or simulation.
+It is interesting to note that the first version of SEQC was formulating the problem using constrained optimization as well, but the slowdown of the stratification step was unacceptable and impractical for large quantum circuits, and thus was abandoned in favor of the heuristics we currently employ in our compiler.
+Bandic et al.
+(bandic2023mapping) map circuit partitioning to min-cut and use the QUBO model with quantum solvers to generate a set of cuts intended to minimize swaps between partitions.
 QUBO employs quantum optimization algorithms on quantum annealers to accelerate the solution process and remain practical, as the search space for transmission qubits is $2^{n}$.
-SEQC, in contrast, does not require acceleration through quantum means, but at the cost of not guaranteeing an optimal solution (it is rather a “best-effort” approach that works well in practice).
-Through discussion of scalability issues impeding the development of larger monolithic quantum circuits, Ovide et al.
-(ovide2023mapping) motivate the necessity of mapping quantum circuits to multi-core architectures and present a commonly agreed-upon likely computing paradigm which leverages both quantum and classical methods to foster inter-chip communication, circumventing the scalability challenges inherent to monolithic quantum processors.
-While this work advocates for a two-level, hierarchical approach for quantum circuit mapping, it leaves the actual problem of qubit placement and then routing unsolved.
-SEQC provides an algorithm and implementation, addressing both of these challenges.
-Using iterative optimization via the Hungarian algorithm, Hungarian Qubit Assignment (HQA) (escofet2023hungarian) seeks to optimize the qubit-to-core mapping problem.
-Because the Hungarian algorithm is here extended with the ability to adjust a qubit’s cost matrix according to behavior multiple timeslices ahead, the model can leverage lookahead to optimize further the number of inter-core swaps.
-However, HQA operates under the assumption that all inter-chiplet communications have the same cost, meaning chiplets must be fully connected.
-SEQC, on the other hand, is able to work with any arrangement of chiplets and inter-chiplet links so long there is a path between any two chiplets, taking into account the distance between pairs of chiplets to generate improved circuits.
-Iterating on their previous work on Hungarian Qubit Assignment, Escofet et al.
-(escofet2024revisiting) develop a characterization technique for arbitrary quantum circuits which yields theoretical upper and lower optimization bounds.
-This is significant, since it enables optimality assessment for solutions to the qubit mapping problem.
+Instead, SEQC takes a “best-effort” approach that works well in practice, and does not require acceleration through quantum means.
+Some works on qubit mapping share a few characteristics with SEQC.
+In a bottom-up approach to qubit mapping, Liu et al.
+(liu_2023_tackling) partition a circuit into blocks that span an identical number of qubits, perform permutation-aware synthesis on each block in parallel (solving for all possible qubit topologies), and finally use an augmented SABRE algorithm to map the blocks to the target device and route between blocks.
+SEQC takes a top-down approach instead, where SWAP permutations between chiplets are settled prior to any compilation tasks performed for each chiplet (block).
+As a result, SEQC only needs to solve for one qubit layout per block, rather than all permutations.
+Hungarian Qubit Assignment (HQA) (escofet2023hungarian) extends the Hungarian algorithm with the ability to adjust a qubit’s cost matrix according to behavior multiple time slices ahead.
 Pastor et al.
-(pastor2024circuit) leverage deep reinforcement learning to partition quantum circuits for execution across multiple processing cores.
-Using a DRL model enhanced with policy-masking mechanisms and a lookahead-capable observation array, the work finds valid qubit-circuit mappings for each of timeslice while prioritizing the reduction of inter-core swaps.
-Notably, it only considers inter-core swaps for optimization, deeming intra-core single-and-multiple qubit operations so inexpensive that their optimization should serve as fodder for inter-core swap reductions.
-This is sensible, but leaves optimization potential on the table, since it only seeks to generate valid mappings within each timeslice.
-Additionally, similar to HQA, this technique treats all inter-chiplet communications equally, further missing optimization opportunities with non-fully connected chiplet topologies.
-SEQC, in contract, is cognizant of the differing costs of moving information between each pair of chiplets, and optimizes accordingly.
-To solve the qubit allocation (mapping) problem, Liu et al.
-(liu_2023_tackling) develop an algorithm that combines partitioning, permutation-aware synthesis, and a modified version of SABRE.
-In a bottom-up approach, this technique partitions a logical circuit into blocks, performs permutation-aware synthesis on each block in parallel (solving for all possible qubit topologies), and finally uses an augmented SABRE algorithm to perform initial layout and routing between blocks.
-In contrast, SEQC takes a top-down approach, where SWAP permutations between chiplets are settled prior to any compilation tasks performed within each block.
-As a consequence, we only need to generate solutions for one qubit layout per block, rather than synthesizing all possible block permutations.
+(pastor2024circuit) use deep reinforcement learning to partition quantum circuits for execution across multiple chiplets.
+Both works aim to reduce inter-chiplet swaps, but assume that all inter-chiplet communications have the same cost, i.e., chiplets are fully connected, which is both impractical for large-scale systems and misses optimization opportunities.
+SEQC, in contrast, can work with arbitrary chiplet topologies.
+Ovide et al.
+(ovide2023mapping) leverage both quantum and classical methods for inter-chiplet communication to address the scalability challenges of monolithic processors.
+While it advocates for a two-level, hierarchical approach for quantum circuit mapping, it leaves the actual problem of qubit placement and routing unsolved.
+SEQC provides an algorithm and implementation, addressing both challenges.
 Parallelizing SEQC’s recurrent elaboration step was inspired by traditional classical compilation techniques, which has a decades-long history and is now commonplace (Mickunas_1978_ParallelCompilation; Gross_1989_ParallelCompilation).
@@
 Eyeriss (Eyeriss_2016_Chen) approaches the dataflow mapping problem by proposing a row-stationary architecture, and then optimizing the mapping of problem to processing element such that elements which are spatially local in the problem are spatially local in hardware during execution.
-This is not unlike qubit swap optimization, and in fact is performed for many of the same reasons — like inter-chiplet swaps, data movements between more distant classical programming elements in a dataflow architecture are more expensive, and minimizing their frequency yields performance improvements.
+This is not unlike the target optimization for the stratification stage, and in fact is performed for many of the same reasons — like inter-chiplet swaps, data movements between more distant classical programming elements in a dataflow architecture are more expensive, and minimizing their frequency yields performance improvements.
 8. Conclusion
-Physical, technological and economic considerations make the prospect of a single “monolithic” large-scale quantum processor very expensive or entirely unrealistic by current standards.
-These constraints force future quantum systems to adopt a physically distributed architecture, based on modular quantum chiplets.
-As such, modular quantum processors already dominate the roadmaps of several major quantum companies.
-However, the complexity of chiplet-based quantum devices, coupled with their growing size, presents an imminent scalability challenge for quantum compilation.
-Existing qubit allocation methods are often unable to contend with inter-chiplet links, which don’t necessary support a universal basis gate set, and existing methods of logical-to-physical qubit placement, routing, unitary synthesis, and quantum circuit optimization, are typically not designed for qubit links that vary significantly in terms of duration or fidelity.
+Physical, technological, and economic considerations force future quantum systems to adopt a physically distributed architecture.
+Modular quantum processors already dominate the roadmaps of several major quantum companies.
+However, quantum compilers face a scalability challenge brought by the complexity and increasing scale of chiplet-based quantum devices, and their lack of chiplet awareness impedes their ability to produce high-quality results.
 In this work, we first modify stock contemporary compilers to become aware of hardware modularity, thereby allowing them to correctly compile circuits for modular architectures with limited cross-chiplet gate support.
@@
 SEQC performs compilation in two stages, with the first stage (stratification, or chiplet splitting) performed only once for a given architecture, and the second stage (elaboration, or chiplet compilation) performed in parallel for each chiplet.
-Unlike contemporary quantum compilers that have to compile a quantum circuit anew before each execution, only the elaboration stage of SEQC needs to be performed recurrently before each execution.
+Unlike contemporary quantum compilers that have to compile a quantum circuit anew before each execution to capitalize on the latest qubit calibration results, only the elaboration stage of SEQC needs to run before each execution.
 As its computational complexity is a function of the quantum chiplet size, not the entire quantum processor size, and chiplet sizes grow slowly, SEQC’s recurrent compilation time is largely unaffected by the relentless growth of qubit counts in future quantum processors.
-We design and implement several novel methods for qubit placement, qubit routing, and circuit optimization in SEQC, with both hardware modularity and compilation parallelization in mind.
-Owning to these techniques and its parallelization strategy, SEQC compiles circuits that exhibit up to a 36% increase in circuit fidelity and up to 1.92×1.92$\times$1.92 × lower execution time, while consistently achieving 24×2-4$\times$2 - 4 × faster compilation time over a chiplet-aware Qiskit baseline.
+SEQC implements several novel methods for qubit placement, qubit routing, and circuit optimization, with both hardware modularity and compilation parallelization in mind, achieving on average $9.3\%$ higher circuit fidelity (up to $49.99\%$), while attaining on average 3.27×3.27$\times$3.27 × faster compilation (up to 6.74×6.74$\times$6.74 ×) compared to a chiplet-unaware Qiskit baseline.
 Acknowledgements

### v4

@@
-Modular Compilation for Quantum Chiplet Architectures
+ab,ab.braket
 Abstract.
-As quantum computing technology matures, industry is adopting modular quantum architectures to keep quantum scaling on the projected path and meet performance targets.
-However, the complexity of chiplet-based quantum devices, coupled with their growing size, presents an imminent scalability challenge for quantum compilation.
-Contemporary compilation methods are not well-suited to chiplet architectures — in particular, existing qubit allocation methods are often unable to contend with inter-chiplet links, which don’t necessarily support a universal basis gate set.
-Furthermore, existing methods of logical-to-physical qubit placement, swap insertion (routing), unitary synthesis, and/or optimization, are typically not designed for qubit links of significantly varying latency or fidelity.
-In this work, we propose SEQC, a hierarchical parallelized compilation pipeline optimized for chiplet-based quantum systems, including several novel methods for qubit placement, qubit routing, and circuit optimization.
-SEQC attains a $9.3\%$ average increase in circuit fidelity (up to $49.99\%$).
-Additionally, owing to its ability to parallelize compilation, SEQC achieves 3.27×3.27$\times$3.27 × faster compilation on average (up to 6.74×6.74$\times$6.74 ×) over a chiplet-unaware Qiskit baseline.
+[…]
 Quantum computing, quantum compilation, modular architectures.
@@
 Recent developments in quantum chip linking, including flip-chip architectures (gold2021entanglement) and low-loss coaxial cables (niu2023low), suggest that modular designs are the most viable for scaling (smith_2022_chiplets).
-Many contemporary or upcoming leading quantum systems adopt chiplet-based modular quantum processors, for example using high-bandwidth quantum links between nearest-neighbor quantum chiplets (e.g., carrier-chip couplers in Rigetti Aspen-M (Rigetti_AspenM; gold2021entanglement), or m-couplers in IBM Crossbill (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers)), or lower-fidelity, lower-bandwidth, but longer-distance flexible coupling of discrete chips (e.g., l-couplers in IBM Flamingo (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers)), or a combination of the above (e.g., c-, m-, and l-couplers in IBM Starling (Mandelbaum_2024_IBMCouplers)), or multi-chip connectivity through tunable couplers and routing chips (Field_2024_RigettiModular).
-Most major companies have set their sights at modular designs to meet quantum scaling targets in practical ways, and modular quantum processors dominate their latest roadmaps (IBM_2023_Roadmap; IonQ_2024_Roadmap; Rigetti_2024_Roadmap).
+Many contemporary or upcoming leading quantum systems adopt chiplet-based modular quantum processors, for example using high-bandwidth quantum links between nearest-neighbor quantum chiplets (e.g., carrier-chip couplers in Rigetti Aspen-M (Rigetti_AspenM; gold2021entanglement), or m-couplers in IBM Crossbill (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers)), or lower-fidelity, lower-bandwidth, but longer-distance flexible coupling of discrete chips (e.g., l-couplers in IBM Flamingo (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers)), or a combination of the above (e.g., c-, m-, and l-couplers in IBM Starling (Mandelbaum_2024_IBMCouplers)), or multi-chip connectivity through tunable couplers and routing chips in Rigetti Cepheus (Rigetti_Cepheus; Field_2024_RigettiModular).
+Most major companies have set their sights at modular designs to meet quantum scaling targets in practical ways, and modular quantum processors dominate their latest roadmaps (IBM_2025_FTQC; IonQ_2024_Roadmap; Rigetti_2024_Roadmap).
+At the same time, no single qubit technology simultaneously offers long coherence, fast gates, high connectivity, and scalable, mature fabrication.
+Superconducting qubits offer very fast gate latencies and mature fabrication, but low coherence times and mostly local connectivity.
+Trapped ions have excellent coherence but slow speeds, and all-to-all connectivity but only for the few ions located at the C-cut sapphire window.
+Neutral atoms have excellent scalability and arrangeable connectivity, but do not reach the gate speeds and manufacturing robustness of superconducting qubits, or the coherence times of trapped ions.
+Hybrid, modular quantum architectures, comprising QPUs based on different qubit modalities, provide a path to combine the complementary strengths of each platform, enabling performance impossible for any single modality.
+For example, superconducting qubits can serve as fast computational QPUs, while ions can provide memory QPUs, and Rydberg atoms can provide ancillae and distillation functionality.
+Such heterogeneity mirrors the classical transition from monolithic CPUs to systems composed of CPUs, GPUs, NPUs, and specialized accelerators.
+As such, hybrid modular quantum systems are increasingly recognized as a pragmatic design path for long-term universal quantum computers (DARPA_HARQ; HQAN; hetarch; Xiang_superconducting_hybrid_2013).
+Efforts to realize hybrid quantum systems have predominantly leveraged quantum state transfer, i.e., SWAPs, between heterogeneous qubit technologies (Kurizki_hybrid_2015; Clerk2020; Zhou2014; PhysRevLett.108.130504).
+While universal quantum operations, e.g., CNOT or CZ operations, that interface two different quantum technologies have been demonstrated (PhysRevLett.114.080501; Xiang_superconducting_hybrid_2013; Cia_2024; Daniilidis_2013; Yu2016), these often have to compromise on gate fidelity, and have yet to be realized for all combinations of common qubit technologies, e.g, between superconducting and trapped-ion qubits.
+Thus, hybrid quantum systems may exhibit interconnect heterogeneity, where inter-module links may not be able to support universal gate sets, while intra-module couplers do.
 Unfortunately, current compilation infrastructure does not support quantum interconnect heterogeneity.
@@
 Coupled with the hardness of synthesizing unitaries (which is exponential in the number of qubits) and the need for quick and frequent compilation, modern quantum compilers have no option other than to rely on heuristics to perform the task.
-These heuristics result in a compilation complexity that is quadratic to the number of qubits in the quantum processor (experimentally observed in Figure 12).
-Hardware modularity adds significant complexity to this already hard task, as inter-chiplet links are typically inferior compared to intra-chiplet ones (laracuente2022modeling; smith_2022_chiplets), connectivity across chiplets is often limited (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers), and not all basis gates are necessarily supported across chiplets (niu2023low).
-In fact, popular quantum software stacks today (e.g., Qiskit (qiskit2024)) are not even cognizant of the existence of hardware modularity.
+These heuristics result in a compilation complexity that is quadratic to the number of qubits in the quantum processor (experimentally observed in Figure 11(a)).
 Hardware modularity, though, also presents an opportunity.
@@
 The individual object files are then linked together to construct the executable.
-We propose a compilation framework for modular quantum processors that works in a similar fashion: it stratifies, i.e., splits, the source quantum circuit into subcircuits small enough to fit in each chiplet, and maps subcircuits to chiplets, and then in parallel elaborates each subcircuit and compiles it for its target chiplet.
-This Stratify-Elaborate Quantum Compiler (SEQC) stratifies a source program only once for a given chiplet architecture, and performs only the elaboration step recurrently before each execution.
+We propose a compilation framework for modular quantum processors that works in a similar fashion: it stratifies, i.e., splits, the source quantum circuit into subcircuits small enough to fit in each module, and maps subcircuits to modules, and then in parallel elaborates each subcircuit and compiles it for its target module.
+This Stratify-Elaborate Quantum Compiler (SEQC) stratifies a source program only once for a given module architecture, and performs only the elaboration step recurrently before each execution.
 As the qubit capacity $n$ of quantum processors grows exponentially, the $O(n^{2})$ compilation latency of contemporary compilers rises even faster.
-SEQC replaces the recurrent $O(n^{2})$ compilation step with several parallel $O(k^{2})$ elaboration steps for $k$-qubit chiplets.
+SEQC replaces the recurrent $Oa(n^{2})$ compilation step with several parallel $O(k^{2})$ elaboration steps for $k$-qubit modules.
 The stratification step is $O(n^{2})$, so the end-to-end complexity remains the same, but stratification is performed only once.
-The recurrent compilation in SEQC is only $O(k^{2})$, and as the number of qubits $k$ per chiplet is expected to remain relatively stable, or grow at a much slower pace in the foreseeable future (IBM_2023_Roadmap), SEQC’s recurrent compilation latency will barely grow as processors scale.
-Additionally, as SEQC is cognizant of hardware modularity, it can stratify the source quantum circuit into subcircuits and map them to chiplets to minimize inter-chiplet communication.
-As we show in this paper, SEQC produces circuits with shorter execution times and significantly fewer inter-chiplet gates compared to today’s stock compilers, leading to higher execution fidelity.
+The recurrent compilation in SEQC is only $O(k^{2})$, and as the number of qubits $k$ per module is expected to remain relatively stable, or grow at a much slower pace in the foreseeable future (IBM_2025_FTQC), SEQC’s recurrent compilation latency will barely grow as processors scale.
+Additionally, as SEQC is cognizant of hardware modularity, it can stratify the source quantum circuit into subcircuits and map them to modules while minimizing cross-module communication.
+As we show in this paper, SEQC produces circuits with shorter execution times and significantly fewer inter-module gates compared to today’s stock compilers, leading to higher execution fidelity.
 More importantly, as the number of qubits in a processor grows, SEQC achieves even higher performance in these figures of merit.
 In summary, the contributions of this paper are as follows:
-  • We make stock compilers aware of hardware modularity, thereby allowing them to correctly compile circuits for modular architectures with limited cross-chiplet gate support.
+  • We make stock compilers aware of modularity through a “peephole” correction pass, thereby allowing them to correctly compile circuits for modular architectures with limited cross-module gate support.
   • We design and implement SEQC, a Stratify-Elaborate Quantum Compiler for modular architectures.
-SEQC performs compilation in two stages, with the first stage (stratification, or chiplet splitting) performed only once for a given architecture, and the second stage (elaboration, or chiplet compilation) performed in parallel for each chiplet.
+SEQC performs compilation in two stages, with the first stage (stratification, or module splitting) performed only once for a given architecture, and the second stage (elaboration, or module compilation) performed in parallel for each module.
 Only this second stage needs to be performed before each execution, and thus SEQC’s compilation time is largely unaffected by the growth of qubit counts in future quantum processors.
   • We design and implement in SEQC several novel methods for qubit placement, qubit routing, and circuit optimization.
-SEQC compiles circuits with $9.3\%$ higher circuit fidelity on average (up to $49.99\%$), while consistently achieving 3.27×3.27$\times$3.27 × faster compilation time on average (up to 6.74×6.74$\times$6.74 ×) compared to a chiplet-unaware Qiskit baseline.
+In particular, we classify swaps into symbiotic, commensalistic, and parasitic, based on how swap placement affects inter-module gates, and develop algorithms to promote symbiotic and avoid parasitic swap placement; we develop qubit “pinning” techniques to allow elaboration without disrupting module connectivity; and we adapt Simulated Annealing and Spatio-Temporal Aware Qubit Allocation (STA) techniques to aid in circuit stratification.
+  • We evaluate SEQC on a modular quantum system consisting of small-scale (10-qubit heavy-hex) modules, and find that it compiles circuits with $9.3\%$ higher circuit fidelity on average (up to $49.99\%$), while consistently achieving faster compilation times ($3.27\times$ on average, up to $6.74\times$) compared to a module-unaware Qiskit baseline.
+When evaluated on a system with large-scale (120-qubit grid topology) modules, SEQC achieves $23.3\%$ higher fidelity on average ($63.36\%$ max), while compiling $1.34\times$ faster on average (up to $3.37\times$).
+More importantly, SEQC outperforms Qiskit even when the latter is afforded significant allowances (no universal gate restrictions, and much lower inter-module gate errors).
 2. Background and Motivation
-(a) Example 2-chiplet architecture
-(b) Input circuit from qubit allocation
-(c) Insertion of perfect-shuffle
-(d) Insertion of return SWAPs
-Figure 1. Peephole correction for incompatible (non-SWAP) gates placed on inter-chiplet edges. Gates that are unimplementable on hardware are shown in red.
-To execute a program on a target device, the logical program representation must be transformed such that it conforms to the idiosyncratic physical constraints for that particular device.
-This process of logical-to-physical translation represents a critical part of compilation and is present even in exotic computing technologies, e.g., quantum.
-In quantum computing, programs are (usually) represented as quantum circuits, and devices are subject to domain-specific physical constraints, such as decoherence time (mike_and_ike), qubit topology (the connectivity between physical qubits) (qubit_allocation; qubit_routing), and the set of basis gates (the family of operations physically implemented on the device).
-Thus, quantum circuit compilation is necessary to actually run quantum programs on real-world devices.
-Compilation is often divided into distinct stages such as qubit allocation (layout and routing), basis translation, and optimization.
-The relations between each stage and the physical constraints of both monolithic and chiplet architectures are described below.
+Compilation is often divided into distinct qubit allocation (layout and routing), basis translation, and optimization stages.
 Qubit Allocation
@@
 It is critical that after this stage, which often requires the modification or addition of circuit gates, the resulting optimized, technology-dependent circuit is functionally equivalent to the original, technology-independent algorithm (smith2019quantum).
-The qubit allocation problem has been previously shown to be NP-complete (qubit_allocation).
-Thus, algorithms often adopt various strategies to make the problem more tractable.
-One common method is to further divide qubit allocation into two substages: layout and routing.
-Logical-to-physical qubit layout, also known as qubit placement, involves deriving an injective map from the virtual qubits of the quantum circuit to the physical qubits of the quantum device (qubit_routing; sabre_swap).
+As qubit allocation is NP-complete (qubit_allocation), it is typically divided into two substages, layout and routing, to make it more tractable.
+Logical-to-physical qubit layout, also known as qubit placement, involves deriving an injective map from the virtual qubits of the quantum circuit to the physical qubits of the quantum device (qubit_routing; sabre_swap), cognizant of its limited connectivity.
 In other words, the qubit layout stage decides the initial position of logical qubits on the device.
-In contrast, qubit routing operates from an initial qubit layout and inserts SWAP operations to align a quantum circuit’s gates to a given qubit topology.
+Qubit routing operates from an initial qubit layout and inserts SWAP operations to align a quantum circuit’s gates to a given qubit topology.
 Consider, for example, the state-of-the-art in qubit allocation: the SABRE heuristic search algorithm (sabre_swap).
-In SABRE, the distinction of the layout and routing stages motivates an iterative method for solving qubit allocation.
-Namely, an initial layout is used to perform qubit routing, which is used to generate a new initial layout.
-Because qubit routing remains NP-hard (qubit_routing), another method of simplifying the problem is to assume qubit links are roughly equivalent, particularly with respect to average fidelity and two-qubit gate duration.
+In SABRE, the distinction of the layout and routing stages motivates an iterative method for solving qubit allocation: an initial layout is used to perform qubit routing, which, in turn, is used to generate a new initial layout.
+Because qubit routing remains NP-hard (qubit_routing), the problem is often simplified by assuming qubit links are roughly equivalent, particularly with respect to average fidelity and two-qubit gate duration.
 This assumption can manifest implicitly when a given device topology is represented as an unweighted, undirected graph.
 For monolithic architectures, one could debate the marginal tradeoff between performance and solution quality.
-However, for modular architectures, the cross-chiplet connections are far worse compared to intra-chiplet connections (smith_2022_chiplets).
-Thus, we could expect the solutions generated from the simplified model to be worse than a more physically accurate model, for example, in circumstances where the shortest path in terms of the quantity or depth of inserted SWAPs diverges from the shortest path in terms of fidelity, gate duration, or some combination thereof.
+However, for modular architectures, the cross-module connections are expected to be far worse compared to intra-module connections (smith_2022_chiplets; PhysRevLett.114.080501; Zhou2014; PhysRevLett.108.130504; Yu2016).
+Thus, as we will show, the current simplified model applied on modular architectures leads to worse solutions than a more physically accurate one.
+(a) Example 2-module
+architecture
+(b) Input circuit from
+qubit allocation
+(c) Insertion of
+perfect-shuffle SWAPs
+(d) Insertion of return SWAPs
+Figure 1. Peephole correction for inter-module gates that are unimplementable on hardware (i.e., non-SWAP, shown in red).
 Basis Translation
 Basis translation converts the gates of an input quantum circuit into gates that are physically supported on hardware.
-Monolithic quantum architectures typically support a so-called universal basis gate set, where any unitary quantum operation can be constructed using solely the gates in the set (mike_and_ike).
-Moreover, it has been shown that basis translation can be performed efficiently given the target gate set is universal, and the quantum circuit contains a fixed, finite number of qubits (Solovay–Kitaev).
-Unfortunately, modular quantum architectures depend upon specialized links to connect otherwise-independent chiplets.
-In the general case, the cross-chiplet links can only be used for data movement (such as SWAPs) and not computation, meaning these links may not support universal operations (niu2023low).
+Basis translation can be performed efficiently given the target gate set is universal (mike_and_ike), and the quantum circuit contains a fixed, finite number of qubits (Solovay–Kitaev).
+Unfortunately, modular quantum architectures depend upon specialized links to connect otherwise-independent modules.
+In the general case, the cross-module links can only be used for data movement (such as SWAPs) and not computation, meaning these links may not support universal operations (niu2023low).
 Thus, compilers that depend on the expectation of universality among qubit links (qiskit_basis_translation_errors), and do not expect these heterogenous, non-universal basis gate sets, can generate incompatible outputs when targeting modular architectures.
 Optimization
-Quantum circuit optimization seeks to prune extraneous operations by finding opportunities to combine, eliminate, and/or parallelize quantum gates (optimization1).
-Through commutation and cancellation identities, it is possible to search for optimization opportunities by iteratively transforming the quantum circuit without affecting the underlying operation (optimization1).
-However, in general, optimization is highly time and resource intensive, with the problem scaling exponentially with respect to number of qubits and/or circuit depth (qfactor).
+Quantum circuit optimization seeks to prune extraneous operations by finding opportunities to combine, eliminate, and parallelize quantum gates (optimization1).
+Optimization is highly time and resource intensive, with the problem scaling exponentially to the number of qubits and circuit depth (qfactor).
 As a result, partitioning, the division of a quantum circuit into blocks of quantum gates, is a common tactic among optimization techniques (qfactor; qgo; bqskit).
@@
 Namely, one forfeits all optimization opportunities along boundaries of demarcation, so they must be carefully chosen to minimize the negative impact on the output solution.
-From this perspective, modular architectures are naturally synergistic with partitioning, since cross-chiplet connections present a physically motivated boundary between operations.
-3. Compilation for Modular Quantum Systems
-Figure 2. High-level diagram of SEQC. Elements in red reflect modifications from existing compiler designs.
+Modular architectures are naturally synergistic with partitioning, since cross-module connections present a physically motivated boundary between operations.
+3. Compilation for Hybrid Modular Quantum Systems
+While SEQC can be applied on a modular quantum system comprising hybrid QPUs, in this paper we consider the simpler case of homogeneous modules as a starting point.
+The reasons behind this decision are simple.
+First, the primary goal of this paper is to introduce and evaluate the impact of the circuit stratification compilation stage with STA and Simulated Annealing we developed, along with qubit pinning, and symbiotic swap insertion.
+The presence of diverse-modality QPUs with varying capabilities would dilute these results, and make it hard to separate their effect from the effects of heterogeneity.
+At the same time, it would make comparisons to the existing state-of-the-art hard, as, to the best of our knowledge, there are no compilers that can target multiple modalities simultaneously for different parts of the same input circuit.
+Thus, in order to isolate the impact of our techniques, and facilitate comparisons with existing tools, in this paper we describe and evaluate SEQC assuming homogeneous modules of the same technology.
+When extending SEQC to a hybrid system, one only needs to execute different elaboration stages, one for each QPU modality, but the stratification stage and the other techniques we develop will remain largely unaffected.
+Thus, for the remainder of this paper, we will be focusing on homogeneous QPUs based on transmons, and we leave qubit-specific stratification optimizations, and the implementation of elaboration stages for other qubit species for future work.
 Modular quantum architectures introduce new challenges to quantum compilation that impact the correctness of solutions and the speed of compilation.
-As discussed previously in §2, our model of modular quantum architectures considers the possibility that inter-chiplet links can be non-universal.
-This is representative of hardware manufacturing constraints (niu2023low), where only a restricted set of operations, e.g., state transfers or SWAPs, may be supported over long-distance couplers (inter-chiplet links).
+As discussed previously in §2, our model of modular quantum architectures considers the possibility that inter-module links can be non-universal.
+This is representative of hardware manufacturing constraints (niu2023low), where only a restricted set of operations, e.g., state transfers or SWAPs, may be supported over long-distance couplers (inter-module links).
 However, many compiler designs (including designs for modular architectures, see §7) experience errors or crashes when presented with non-universal, heterogenous, basis gate sets.
@@
 To address the mounting pressure of ever-growing compilation workloads, these compilers must be cognizant of classical hardware limitations and tradeoffs between solution quality and compilation latency.
-To ameliorate these issues, we present a family of compilers optimized for modular quantum architectures that support chiplet links with non-universal basis gate sets, and leverage modular compilation strategies.
+To ameliorate these issues, we present a family of compilers optimized for modular quantum architectures that support module links with non-universal basis gate sets, and leverage modular compilation strategies.
 In Section 3.1, we propose a peephole compilation pass that can be inserted into any compiler to add support for non-universal basis gate sets.
-This would allow any compiler to be compatible with chiplet-based devices without significant changes to their underlying implementations.
+This would allow any compiler to be compatible with module-based devices without significant changes to their underlying implementations.
 Next, in Section 3.2, we introduce the stratify-elaborate quantum compiler (SEQC), which advocates for a hierarchical two-stage compilation that produces both better circuits for execution and faster compilation times.
-3.1. Chiplet-awareness via Peephole Correction
+3.1. Module-awareness via Peephole Correction
 The design of the peephole correction pass was inspired by the “peephole” optimizations commonly used in classical compilation (dragon_book).
 Peephole passes operate within a nominal peephole window and substitute its contents with a logically equivalent replacement.
-Notable examples in the context of quantum compilation include basis translation and elimination of self-adjoint gates.
-Peephole correction (Figure 1) operates on a circuit generated by qubit allocation, where it detects when non-SWAP gates are placed on non-universal inter-chiplet links (Figure 1(b)), and employs a cyclic qubit rotation, i.e., a perfect shuffle permutation, to move the aforementioned gates to nearby intra-chiplet links (Figure 1(c)).
+Peephole correction (Figure 1) operates on a circuit generated by qubit allocation, where it detects when non-SWAP gates are placed on non-universal inter-module links (Figure 1(b)), and employs a cyclic qubit rotation, i.e., a perfect shuffle permutation, to move the aforementioned gates to nearby intra-module links (Figure 1(c)).
 Additional “return” SWAPs are then inserted to return the qubits to their initial locations (Figure 1(d)).
-Overall, for every inter-chiplet gate in the original circuit generated by SABRE, the peephole pass inserts 4 additional SWAP gates — 2 inter-chiplet and 2 intra-chiplet.
-Ultimately, this degree of overhead is inevitable when the stock, unmodified compiler is unaware of chiplet-specific device constraints.
-For example, skipping the return SWAP operations would initially appear to save $2$ SWAPs, thus halving the cost per inter-chiplet gate.
-However, this hypothetical design would also no longer be a true peephole, since it introduces a net qubit permutation that propagates to downstream gates in the circuit.
-From the perspective of modular compilation, this results in serial dependencies that prevent parallel execution of the pass.
-From a routing perspective, the resultant qubit permutations disrupt the post-routing layout of qubits, leading to destructive routing errors that can only be resolved by an entirely new routing pass.
-Overall, such a hypothetical pass results in substantially worse output circuits compared to the peephole method.
+Overall, for every inter-module gate in the original circuit generated by SABRE, the peephole pass inserts 4 additional SWAP gates — 2 inter-module and 2 intra-module.
+Ultimately, this degree of overhead is inevitable when the stock, unmodified compiler is unaware of module-specific device constraints.
+For example, skipping the return SWAP operations would appear to save $2$ SWAPs, and halve the cost per inter-module gate.
+However, it would also introduce a net qubit permutation that propagates downstream in the circuit.
+From the perspective of modular compilation, this creates serial dependencies that prevent parallel execution of the pass.
+From a routing perspective, the resultant qubit permutations disrupt the post-routing layout of qubits, causing destructive routing errors that can only be resolved by an entirely new routing pass.
+Overall, such a hypothetical pass would result in substantially worse output circuits compared to the peephole method.
 3.2. The Stratify-Elaborate Quantum Compiler
-We incorporate the hierarchical elements of chiplet architectures into SEQC by adopting a two-stage procedure (Figure 2).
-The first, stratification stage, represents a one-time overhead to elevate a quantum circuit to the chiplet level.
+Figure 2. High-level diagram of SEQC. Elements in red reflect modifications from existing compiler designs.
+We incorporate the hierarchical elements of module architectures into SEQC by adopting a two-stage procedure (Figure 2).
+The first, stratification stage, represents a one-time overhead to elevate a quantum circuit to the module level.
 The second, elaboration stage, represents a recurring cost to complete the compilation of a quantum circuit given the most up-to-date backend properties.
 3.2.1. Stratification Stage
-(a) Example chiplet topology
+(a) Example module topology
 (b) Qubit-to-subcircuit mapping
-(c) Subcircuit-to-chiplet routing
-(d) Inter-chiplet routing
-Figure 3. Step-by-step example of inter-chiplet compilation in the Stratification stage.
-In this stage, we map a quantum circuit to the inter-chiplet device topology (Figure 3(a) in our running example) by stratifying it into smaller, interconnected subcircuits.
+(c) Subcircuit-to-module routing
+(d) Inter-module routing
+Figure 3. Step-by-step example of inter-module compilation in the Stratification stage.
+In this stage, we map a quantum circuit to the inter-module device topology (Figure 3(a) in our running example) by stratifying it into smaller, interconnected subcircuits.
 This can be accomplished with the following intermediate steps (Figures 3(b)3(d)).
 First, qubit-to-subcircuit mapping designates which logical qubits belong to which logical subcircuits.
-Then, as part of chiplet allocation, we define a mapping of logical subcircuits to physical chiplets on the target device and route inter-chiplet gates to conform with constraints of physical-chiplet connectivity.
+Then, as part of module allocation, we define a mapping of logical subcircuits to physical modules on the target device and route inter-module gates to conform with constraints of module connectivity.
 Qubit-to-Subcircuit Mapping
-Algorithm 1 Qubit-to-Subcircuit Mapping with Sim. Annealing
-[…]
-[…]
-2:$q_{0},q_{1}\leftarrow\textsc{Qubits}(g)$
-3:$s_{0},s_{1}\leftarrow L[q_{0}],L[q_{1}]$
-4:Increment $C$ if $s_{0}\neq s_{1}$EndFor
-[…]
-[…]
-7:$s_{0}\leftarrow\textsc{RandChoice}(\{1,\dots,N_{C}\})$
-8:$q_{0}\leftarrow\textsc{RandChoice}(L^{\prime}[s_{0}])$
-9:$s_{1}\leftarrow\textsc{RandChoice}(\{1,\dots,N_{C}\}-\{s_{0}\})$
-10:$q_{1}\leftarrow\textsc{RandChoice}(L^{\prime}[s_{1}])$LCommentSwap $q_{0}$ and $q_{1}$
-11:$L^{\prime}[s_{0}]\leftarrow(L^{\prime}[s_{0}]\cup\{q_{1}\})-\{q_{0}\}$
-12:$L^{\prime}[s_{1}]\leftarrow(L^{\prime}[s_{1}]\cup\{q_{0}\})-\{q_{1}\}$EndFor
-[…]
-14:$L\leftarrow$ Initial Layout $\triangleright$ Current Solution
-15:$R\leftarrow L$triangleright Best Solution
-[…]
-17:$D\leftarrow\textsc{Ceil}(T\times\alpha)$triangleright Number of changes to make
-18:$L^{\prime}\leftarrow\textsc{Kick}(D,N_{C},N_{Q},L)$
-19:$\Delta\leftarrow\textsc{Cost}(Circ,L^{\prime})-\textsc{Cost}(Circ,L)$
-20:$k\leftarrow\textsc{Rand}(0,1)$triangleright Random number in $[0,1)\subset\mathbb{R}$
-21:$L\leftarrow L^{\prime}$ if $(\Delta<0)$ or $(k>\exp(\Delta/T))$
-22:$R\leftarrow L^{\prime}$ if $(\textsc{Cost}(Circ,L^{\prime})<\textsc{Cost}(Circ,R))$
-[…]
-24:\$R$EndProcedure
-[…]
 Qubit-to-subcircuit mapping produces a stratified quantum circuit, where logical qubits (numbers in Figure 3(b)) are grouped into subcircuits (lowercase letters).
-An optimal qubit-to-subcircuit mapping minimizes the number of connections between subcircuits, thereby reducing the number of inter-chiplet SWAPs needed to transfer data between chiplets during execution.
+An optimal qubit-to-subcircuit mapping minimizes the number of connections between subcircuits, thereby reducing the number of inter-module SWAPs needed to transfer data between modules during execution.
 To perform this task, we investigate two methods: Simulated Annealing and Spatio-Temporal Aware Qubit Allocation (STA) (ovide2024sta).
 Simulated annealing is an efficient stochastic optimization technique shown to work well on NP-complete problems, such as place-and-route in VLSI algorithms (annealing1989).
-Simulated annealing is an iterative method to optimize a state against a given cost function.
+It is an iterative method to optimize a state against a given cost function.
 In each iteration, the state is “kicked,” or randomly perturbed, and the new state is greedily accepted if there is a reduction in cost.
@@
 This improves the chances of finding a (near-)optimal solution.
-Our method is shown in Algorithm 1.
-We start with an initial qubit-to-subcircuit mapping, which we minimize against a cost function, defined on line 1 as the number of inter-chiplet gates given a certain layout.
-By minimizing the number of inter-chiplet gates, we implicitly reduce the inter-chiplet connectivity.
+Our simulated annealing method (Algorithm 1) starts with an initial qubit-to-subcircuit mapping, which we minimize against a cost function, defined as the number of inter-module gates given a certain layout.
+By minimizing the number of inter-module gates, we implicitly reduce the inter-module connectivity.
 The kick function then swaps random qubits in the layout, ensuring that each SWAP is between qubits in different subcircuits.
-In doing so, we hope to move strongly interacting qubits close together, minimizing the number of cross-chiplet gates.
-As input, our algorithm takes the input circuit $Circ$, the chiplet count $N_{C}$, the chiplet size $N_{Q}$, the starting temperature $T_{0}$, the number of changes to do per degree of temperature $\alpha$, and the cooling factor $\beta$.
+In doing so, we hope to move strongly interacting qubits close together, minimizing the number of cross-module gates.
+As input, our algorithm takes the input circuit $Circ$, the module count $N_{C}$, the module size $N_{Q}$, the starting temperature $T_{0}$, the number of changes to do per degree of temperature $\alpha$, and the cooling factor $\beta$.
 The last three are constant tuning parameters across runs.
+Algorithm 1 Qubit-to-Subcircuit Mapping with Simulated Annealing
+1: procedure Cost($Circ,L$)
+2:   $C\leftarrow 0$ $\triangleright$ Count
+3:   for $g\in\textsc{TwoQubitGates}(Circ)$ do
+4:     $q_{0},q_{1}\leftarrow\textsc{Qubits}(g)$
+5:     $s_{0},s_{1}\leftarrow L[q_{0}],L[q_{1}]$
+6:     Increment $C$ if $s_{0}\neq s_{1}$
+7:   return $C$
+8: procedure Kick($D,N_{C},N_{Q},L$)
+9:   $L^{\prime}\leftarrow\textsc{Copy}(L)$
+10:   for $i=1,\dots,D$ do
+11:     $s_{0}\leftarrow\textsc{RandChoice}(\{1,\dots,N_{C}\})$
+12:     $s_{1}\leftarrow\textsc{RandChoice}(\{1,\dots,N_{C}\}-\{s_{0}\})$
+13:     $q_{0},q_{1}\leftarrow\textsc{RandChoice}(L^{\prime}[s_{0}]),\textsc{RandChoice}(L^{\prime}[s_{1}])$
+14:     $L^{\prime}[s_{0}],L^{\prime}[s_{1}]\leftarrow(L^{\prime}[s_{0}]\cup\{q_{1}\})-\{q_{0}\},(L^{\prime}[s_{1}]\cup\{q_{0}\})-\{q_{1}\}$ $\triangleright$ Swap $q_{0}$ and $q_{1}$
+15:   return $L^{\prime}$
+16: procedure Q2SMappingAN($Circ,N_{C},N_{Q},T_{0},\alpha,\beta$)
+17:   $L\leftarrow$ Initial Layout $\triangleright$ Current Solution
+18:   $R\leftarrow L$ $\triangleright$ Best Solution
+19:   $T\leftarrow T_{0}$ $\triangleright$ Annealing “Temperature”
+20:   while $T>1$ do
+21:     $D\leftarrow\textsc{Ceil}(T\times\alpha)$ $\triangleright$ Number of changes to make
+22:     $L^{\prime}\leftarrow\textsc{Kick}(D,N_{C},N_{Q},L)$
+23:     $\Delta\leftarrow\textsc{Cost}(Circ,L^{\prime})-\textsc{Cost}(Circ,L)$
+24:     $k\leftarrow\textsc{Rand}(0,1)$ $\triangleright$ Random number in $0,1)\subset\mathbb{R}$
+25:     $L\leftarrow L^{\prime}$ if $(\Delta<0)$ or $(k>\exp(\Delta/T))$
+26:     $R\leftarrow L^{\prime}$ if $(\textsc{Cost}(Circ,L^{\prime})<\textsc{Cost}(Circ,R))$
+27:     $T\leftarrow T\times\beta$ $\triangleright$ Decrease temperature
+28:   return $R$
 The STA algorithm is a qubit allocation algorithm designed for trapped-ion quantum systems.
 These devices often feature modular designs, where one device is composed of “traps” arranged in a (typically) line or ring topology, where each trap contains multiple qubits (ions).
-Thus, it seems possible to adapt qubit allocation algorithms designed for trapped-ion systems to support modular transmon systems as well.
-Specifically, one component of STA, denoted as Algorithm 2 in Ovide et al.
-(ovide2024sta), assigns qubits to ion traps, which we can leverage to perform qubit-to-subcircuit mapping and generate an initial subcircuit-to-chiplet mapping.
-Chiplet Allocation
-Algorithm 2 Inter-Chiplet Routing
+Qubit allocation algorithms for modular trapped-ion systems are incentivized to minimize the number of operations between modules due to associated qubit shuttling and relocation overheads (ovide2024sta).
+When compiling for modular transmon architectures, we are similarly incentivized to minimize the number of inter-module interactions.
+Thus, we adapted a component of the STA algorithm (Algorithm 2 in Ovide et al.
+(ovide2024sta)) that assigns qubits to ion traps to perform qubit-to-subcircuit mapping and generate an initial subcircuit-to-module mapping (Algorithm 2).
+The original STA algorithm prioritizes mapping qubits that frequently interact with other qubits (quantified by $R$ in Algorithm 2), or have a large degree of direct or indirect interactions (quantified heuristically as the depth-weighted sum $T$ of interactions between every pair of qubits).
+These are used to derive a sequence of qubit pairs, from which STA constructs a qubit mapping.
+Namely, for each iteration of the STA algorithm, a pair of qubits $(q_{a},q_{b})$ is used to insert at least one qubit into the mapping, based on the shortest topological distance between physical qubits.
+If one of the qubits in the pair is already mapped, say $q_{a}$, then STA also recursively maps any qubits that are more strongly connected to $q_{a}$ than $q_{b}$ is.
+This is done in order to keep strongly connected qubits physically close together.
+Algorithm 2 Qubit-to-Subcircuit Mapping with STA
+1: procedure ComputeRatios($C$)
+2:   $R\leftarrow$ empty map
+3:   for $q\in\textsc{Qubits}(C)$ do
+4:     $R[q]\leftarrow\frac{\textsc{InteractionCnt}(q,C)}{\textsc{NumQubits}(C)}$ $\triangleright$ fraction of qubits $q$ interacts with
+5:   return $R$
+6: procedure ComputeST($C,S$)
+7:   $T\leftarrow$ empty map
+8:   for $(q_{1},q_{2})\in\textsc{AllPairs}(\textsc{Qubits}(C))$ do
+9:     for $(i,s)\in\textsc{enumerate}(S)$ do
+10:      Add $2^{-i}$ to $T[(q_{1},q_{2})]$ if $q_{1}$ and $q_{2}$ interact in $s$
+11:   return $T$
+12: procedure MapQubit($q_{1},T,R,L$)
+13:   $P\leftarrow$ first pair in $T$ containing $q_{1}$; $q_{2}\leftarrow$ other qubit in $P$
+14:   if $\textsc{Index}(q_{2})<\textsc{Index}(q_{1})$ then
+15:     $\textsc{MapQubit}(q_{2},T)$ $\triangleright$ Map $q_{2}$ if it comes first
+16:   if $q_{1}\in R$ or $q_{2}\in R$ then
+17:     $\textsc{PlaceQubits}(q_{1},q_{2},L)$
+18:     $R.\textsc{remove}(q_{1},q_{2})$; $T.\textsc{remove}(P)$ $\triangleright$ Remove placed qubits
+19: procedure Q2SMappingSTA($C$) $\triangleright$ $C$: Quantum Circuit
+20:   $S\leftarrow\textsc{Layers}(C)$; $L\leftarrow$ initial layout
+21:   $R\leftarrow\textsc{ComputeRatios}(C)$; $T\leftarrow\textsc{ComputeST}(S)$
+22:   while $R$ is not empty do
+23:     $\textsc{MapQubit}(R[0],T,R,L)$
+24:   return $L$
+Module Allocation
+Algorithm 3 Inter-Module Routing
+1: procedure FindSwaps($Q,E$)
+2:   return $\textsc{FindSymbioticSwaps}(Q,E)$ if non-empty
+3:   return $\textsc{FindCommensalisticSwaps}(Q,E)$ if non-empty
+4:   return $\textsc{FindParasiticSwaps}(Q,E)$
+5: procedure ValidGates($Q,L$)
+6:   Initialize $V_{virt}$ and $V_{phys}$ to empty sets
+7:   for $g\in Q$ do
+8:     $H\leftarrow\{L[q]|q\in\textsc{Qubits}(g)\}$ $\triangleright$ Get module info
+9:     if $|H|=1$ then $\triangleright$ All qubits on same module
+10:      Add $g$ to $V_{Virt}$ and $(g,H)$ to $V_{phys}$
+11:   return $V_{virt},V_{phys}$
+12: procedure Step($Q,R,L,D$) $\triangleright$ Advances the “wavefront”
+13:   $V_{virt},V_{phys}\leftarrow\textsc{ValidGates}(Q,L)$
+14:   $R\leftarrow R+V_{phys}$
+15:   for $g\in V_{virt}$ do
+16:     Remove $g$ from $Q$
+17:     for $c\in\textsc{Children}(g,D)$ do
+18:      Add $c$ to $Q$ if $p\in R$ for all $p\in\textsc{Parents}(c,D)$
+19: procedure Heuristic($S,R,Circ$)
+20:   $L^{\prime}\leftarrow\textsc{Swap}(\textsc{Copy}(L),S)$; $C\leftarrow 0$
+21:   for $g\in\textsc{TwoQubitGates}(Circ)$ do
+22:     if $g\notin R$ then
+23:      $q_{0},q_{1}\leftarrow\textsc{Qubits}(g)$
+24:      $s_{0},s_{1}\leftarrow L[q_{0}],L[q_{1}]$
+25:      Increment $C$ if $s_{0}\neq s_{1}$
+26:   return $C$
+27: procedure InterModuleRouting($Circ,L,T$) $\triangleright$ $L$ is from Algorithm 2, $T$ is the module topology.
+28:   $D\leftarrow\textsc{CircuitToDAG}(Circ)$
+29:   $E\leftarrow$ All-pairs distances from module topology $T$
+30:   $Q\leftarrow\textsc{FirstLayer}(DAG)$; $R\leftarrow$ empty list
+31:   $\textsc{Step}(Q,R,L,D)$
+32:   while $Q$ is not empty do
+33:     $\mathcal{S}\leftarrow\textnormal{{FindSwaps}}(Q,E)$ $\triangleright$ SWAP Candidates
+34:     $S\leftarrow\operatorname*{arg\,min}_{s\in\mathcal{S}}\textsc{Heuristic}(s,R,Circ)$
+35:     $\textsc{Swap}(L,S)$
+36:     $\textsc{Step}(Q,R,L,D)$
+37:   $D^{\prime}\leftarrow$ DAG built in topological order from $R$ $\triangleright$ $R$ is ordered
 […]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-8:$V_{virt},V_{phys}\leftarrow\textsc{ValidGates}(Q,L)$
-[…]
-[…]
-[…]
-12:$L^{\prime}\leftarrow\textsc{Swap}(\textsc{Copy}(L),S)$
-[…]
-14:$q_{0},q_{1}\leftarrow\textsc{Qubits}(g)$
-15:$s_{0},s_{1}\leftarrow L[q_{0}],L[q_{1}]$
-[…]
-[…]
-18:$D\leftarrow\textsc{CircuitToDAG}(Circ)$
-19:$E\leftarrow$ All-pairs distances from chiplet topology $T$
-20:$Q\leftarrow\textsc{FirstLayer}(DAG)$
-21:$R\leftarrow$ empty list
-[…]
-23:$\mathcal{S}\leftarrow\textnormal{{FindSwaps}}(Q,E)$triangleright SWAP Candidates
-24:$S\leftarrow\operatorname*{arg\,min}_{s\in\mathcal{S}}\textsc{Heuristic}(s,R,Circ)$
-25:$\textsc{Swap}(L,S)$
-26:$\textsc{Step}(Q,R,L,D)$EndWhile
-27:$D^{\prime}\leftarrow$ Empty DAG
-28:Append $g$ to $D^{\prime}$ for each $g\in{R}$triangleright $R$ is ordered
-[…]
-[…]
-Chiplet allocation describes the process of transforming a stratified circuit to be compatible with the inter-chiplet topology of a modular device.
-In Figure 2, we decompose chiplet allocation into subcircuit-to-chiplet mapping and inter-chiplet routing, analogous to how qubit allocation can be decomposed into qubit layout and routing.
-Accordingly, subcircuit-to-chiplet mapping (Figure 3(c)), defines an initial placement of logical subcircuits (lowercase letters) to physical chiplets (uppercase letters).
-Subsequently, inter-chiplet routing (Figure 3(d)) inserts inter-chiplet SWAPs to account for limited cross-chiplet gate connectivity.
-In our design, we use our qubit-to-subcircuit mapping to provide an initial subcircuit-to-chiplet mapping.
-Then, we extend and modify the SABRE algorithm (sabre_swap) to apply to chiplets, placing iterations of inter-chiplet routing alongside randomized qubit permutations to inform a final subcircuit-to-chiplet mapping.
-We did explore a design that extended other components of STA to perform subcircuit-to-chiplet mapping on transmon chiplet systems, but did not produce circuits meaningfully better than this iterative approach and increased compilation time, hence it was abandoned.
-Inter-chiplet routing seeks to consolidate the qubits of (non-SWAP) inter-chiplet gates to a single chiplet by inserting inter-chiplet SWAPs into the stratified circuit.
-For a single two-qubit, inter-chiplet gate, we represent inter-chiplet routing as a shortest-path problem, where the two qubits of the gate correspond to chiplet endpoints in the unweighted inter-chiplet connectivity graph.
-From the shortest path, we can derive an optimal sequence of SWAPs for routing, where the unweighted path length, or chiplet distance, indicates the length of the SWAP chain.
-However, in the context of a stratified circuit, composed of multiple inter-chiplet gates, it becomes possible to interleave SWAP chains (i.e., efficiently reorder their execution) to reduce the depth and/or total number of inter-chiplet SWAPs.
-Like qubit routing, we would expect it to be impractical to derive an optimal solution for this more complex picture of inter-chiplet routing at compile-time.
+Module allocation describes the process of transforming a stratified circuit to be compatible with the inter-module topology of a modular device.
+In Figure 2, we decompose module allocation into subcircuit-to-module mapping and inter-module routing, analogous to how qubit allocation can be decomposed into qubit layout and routing.
+Accordingly, subcircuit-to-module mapping (Figure 3(c)), defines an initial placement of logical subcircuits (lowercase letters) to physical modules (uppercase letters).
+Subsequently, inter-module routing (Figure 3(d)) inserts inter-module SWAPs to account for limited cross-module gate connectivity.
+In our design, we use our qubit-to-subcircuit mapping to provide an initial subcircuit-to-module mapping.
+Then, we extend and modify the SABRE algorithm (sabre_swap) to apply to modular architectures, placing iterations of inter-module routing alongside randomized qubit permutations to inform a final subcircuit-to-module mapping.
+We did explore a design that extended other components of STA to perform subcircuit-to-module mapping on transmon module systems, but it underperformed this iterative approach, and had increased compilation time, hence it was abandoned.
+Inter-module routing seeks to consolidate the qubits of (non-SWAP) inter-module gates to a single module by inserting inter-module SWAPs into the stratified circuit.
+For a single two-qubit, inter-module gate, we represent inter-module routing as a shortest-path problem, where the two qubits of the gate correspond to module endpoints in the unweighted inter-module connectivity graph.
+From the shortest path, we can derive an optimal sequence of SWAPs for routing, where the unweighted path length, or module distance, indicates the length of the SWAP chain.
+However, in the context of a stratified circuit, composed of multiple inter-module gates, it becomes possible to interleave SWAP chains (i.e., efficiently reorder their execution) to reduce the depth and/or total number of inter-module SWAPs.
+Like qubit routing, we would expect it to be impractical to derive an optimal solution for this more complex picture of inter-module routing at compile-time.
 Thus, we employ a similar SWAP selection scheme as SABRE (sabre_swap), where inserted SWAPs are chosen from a pool of SWAP candidates according to a cost heuristic.
-For our inter-chiplet routing algorithm, we categorize SWAPs from our candidate pool into three tiers, based on a concept of the “efficiency” of each SWAP, as shown in Algorithm 2.
-An inter-chiplet SWAP is considered symbiotic if it directly affects two inter-chiplet gates in a “beneficial” manner.
-In other words, each qubit of a symbiotic SWAP corresponds to a different inter-chiplet gate, where the insertion of the SWAP reduces the chiplet distance score of both gates.
-In theory, this outcome represents the most efficient possible insertion for a single SWAP, and thus should be the highest-priority case for our greedy inter-chiplet routing cost heuristic.
-Next, commensalistic SWAPs are those that do not “harm,” or increase that chiplet distance, of any directly impacted gates.
-These SWAPs typically occur when an idle qubit can be used as an ancilla qubit to facilitate the routing of an inter-chiplet gate without affecting other inter-chiplet gates.
-In certain device topologies, a SWAP acting on the qubits of two inter-chiplet gates can also exhibit commensalistic behavior.
-Finally, parasitic SWAPs act on the qubits of two inter-chiplet gates, where one gate benefits while the other is harmed.
+For our inter-module routing algorithm, we categorize SWAPs from our candidate pool into three tiers, based on a concept of the “efficiency” of each SWAP, as shown in Algorithm 3.
+An inter-module SWAP is considered symbiotic if it directly affects two inter-module gates in a “beneficial” manner.
+In other words, each qubit of a symbiotic SWAP corresponds to a different inter-module gate, where the insertion of the SWAP reduces the module distance score of both gates.
+In theory, this outcome represents the most efficient possible insertion for a single SWAP, and thus should be the highest-priority case for our greedy inter-module routing cost heuristic.
+Next, commensalistic SWAPs are those that do not “harm,” or increase that module distance, of any directly impacted gates.
+These SWAPs typically occur when an idle qubit can be used as an ancilla qubit to facilitate the routing of an inter-module gate without affecting other inter-module gates.
+In certain device topologies, a SWAP acting on the qubits of two inter-module gates can also exhibit commensalistic behavior.
+Finally, parasitic SWAPs act on the qubits of two inter-module gates, where one gate benefits while the other is harmed.
 Since this variety of SWAPs is actively counterproductive towards a gate, it should be, and likely will be, avoided in most circumstances.
@@
 3.2.2. Elaboration Stage
-(a) An intra-chiplet topology
-(b) A $3$-qubit subcircuit post-stratification
-(c) Subcircuit after intra-chiplet layout
-(d) Inter-chiplet SWAP lowering
-(e) Intra-chiplet routing
+(a) Intra-module topology
+example
+(b) Stratified 3-qubit subcircuit
+(c) Subcircuit after intra-module layout
+$q_{0},q_{1},q_{2}]\Rightarrow[Q_{2},Q_{0},Q_{1}]$
+(d) Inter-module SWAP lowering
+(e) Intra-module routing
 (f) Basis translation and optimization
-Figure 4. Step-by-step example of intra-chiplet compilation in the Elaboration stage. Gates that cannot be implemented on hardware are shown in red.
-Following the stratification stage, each chiplet can be compiled almost completely independently from one another.
-More precisely, the compilation process within each chiplet is largely identical to that of a monolithic quantum architecture.
-The only exception is the presence of inter-chiplet SWAPs, which must be lowered from the chiplet level to the qubit level.
-Once this assignment is made, these SWAPs must remain fixed and immutable to subsequent compilation stages.
+Figure 4. Step-by-step example of intra-module compilation in the Elaboration stage. Gates that cannot be implemented on hardware are shown in red.
+Following the stratification stage, each module can be compiled almost completely independently from one another.
+More precisely, the compilation process within each module is largely identical to that of a monolithic quantum architecture.
+The only exception is the presence of inter-module SWAPs, which must be lowered from the module level to the qubit level.
+Once this assignment is made, these SWAPs must remain fixed, which we call “qubit pinning” and immutable to subsequent compilation stages.
 Despite this overhead, restricting the task of compilation to a narrower domain of qubits facilitates both parallelization (since each subcircuit can be processed in parallel) and problem size reduction (since there are fewer qubits to consider) of the most computationally intensive compilation stages, namely routing and optimization.
 To accomplish these tasks, we divide the elaboration stage into two substages: qubit allocation and parallel basis translation and optimization.
-Figure 4 illustrates this with an example of elaboration for a single chiplet (Figure 4(a)) and its corresponding subcircuit (Figure 4(b)).
+Figure 4 illustrates this with an example of elaboration for a single module (Figure 4(a)) and its corresponding subcircuit (Figure 4(b)).
 Qubit Allocation
 We employ a modified version of SABRE to perform qubit allocation in the elaboration stage.
-Specifically, we parallelize the SABRE routing algorithm and add constraints to maintain correctness when applied to a stratified quantum circuit with inter-chiplet gates.
-We also explored a layout pass that used portions of the STA algorithm, but the marginal benefits to solution quality were not worth the 56×5-6$\times$5 - 6 × slower elaboration time.
-The placement of inter-chiplet gates is a non-parallelizable task (due to cross-chiplet interactions) that strongly influences the layout and routing passes, which could otherwise be fully parallelized across chiplets.
-To account for this, we introduce a serial stage (Figure 4(d)) between the parallel layout and routing passes that greedily places gates on the nearest valid inter-chiplet edge.
-A greedy policy was chosen because inter-chiplet gates have more restrictions on valid placements compared to intra-chiplet gates, while also experiencing substantially higher error rates (smith_2022_chiplets).
-Once the positions of the inter-chiplet gates are settled, they must remain fixed, imposing an additional constraint on qubit routing.
-To promote this behavior in SABRE, we extend its cost heuristic to weigh the device topology by fidelity.
-This causes higher error links to assume higher costs, discouraging SABRE from moving the high-error inter-chiplet gates.
+Specifically, we parallelize the SABRE routing algorithm and add constraints to maintain correctness when applied to a stratified quantum circuit with inter-module gates.
+We also explored a layout pass that used portions of the STA algorithm, but the marginal benefits to solution quality were not worth the $5-6\times$ slower elaboration time.
+The placement of inter-module gates is a non-parallelizable task (due to cross-module interactions) that strongly influences the layout and routing passes, which could otherwise be fully parallelized across modules.
+To account for this, we introduce a serial stage (Figure 4(d)) between the parallel layout and routing passes that greedily places gates on the nearest valid inter-module edge.
+We choose a greedy policy because inter-module gates have more restrictions on valid placements compared to intra-module gates, while also experiencing substantially higher error rates (smith_2022_chiplets).
+Once the positions of the inter-module gates are settled, they are pinned, imposing an additional constraint on qubit routing.
+To softly encourage this behavior in SABRE, we extend its cost heuristic to weigh the device topology by fidelity.
+This causes higher error links to assume higher costs, discouraging SABRE from moving the high-error inter-module gates.
 As a result, SEQC achieves better solutions in the common case.
-However, there is nothing to prevent SABRE from incorrectly modifying any inter-chiplet SWAPs.
+However, there is nothing to prevent SABRE from incorrectly modifying any inter-module SWAPs.
 Thus, we need to impose hard restrictions to ensure correctness.
-We address this by establishing two disjoint gate sets.
-One gate set operates strictly within a chiplet.
-This is the traditional gate set for a monolithic processor.
-The other gate set is the only one permitted to operate across chiplets, and specifically only implements inter-chiplet SWAPs.
+To address this we establish two disjoint gate sets: one gate set that operates strictly within a module (the traditional gate set for a monolithic processor), and another that is the only gate set permitted to operate across modules (and, specifically for our example implementation, only implements inter-module SWAPs).
 By restricting each gate set to its own disjoint part of the overall device topology, we ensure correctness across all cases.
 Parallel Translation and Optimization
-The remaining compilation stages of basis translation and circuit optimization can be easily parallelized at the chiplet granularity (Figure 4(f)).
+The remaining compilation stages of basis translation and circuit optimization can be easily parallelized at the module granularity (Figure 4(f)).
 Basis translation, in particular, is embarrassingly parallel: gates can be translated independently from each other.
 Meanwhile, as discussed in §2, optimization algorithms already employ partitioning and parallelization to create tractable subproblems from intractably large circuits, facilitating a smaller problem space, lower memory usage, and faster performance.
-In modular quantum architectures, we physically motivate partitions through the natural chiplet boundaries, enabling us to reap the same resource and performance benefits of partitioning without compromising on solution quality.
+In modular quantum architectures, we physically motivate partitions through the natural module boundaries, enabling us to reap the same resource and performance benefits of partitioning without compromising on solution quality.
 4. Methodology
 We evaluate quantum compilers based on the quality of their solutions, their speed of compilation, and their ability to scale to large circuits/devices.
-To experimentally profile these characteristics, we benchmark a compiler on a suite of quantum circuits, detailed in §4.1, targeting a mock chiplet-based quantum backend, as specified in §4.2.
-In total, we evaluate four compilers—a reference compiler, a peephole-augmented compiler, and two versions of SEQC—on simulated backends ranging from 1 to 100 chiplets.
-All compilers are implemented in the Qiskit SDK (qiskit2024) v1.2.41.2.41.2.41.2.4 and Python $3.12.6$, and timing results are taken on a machine with a $128$-core Ampere Altra Max ARM64 processor and $256$ GB of DDR4 RAM.
+To experimentally profile these characteristics, we benchmark a compiler on a suite of quantum circuits, detailed in §4.1, targeting a mock module-based quantum backend, as specified in §4.2.
+In total, we evaluate four compilers—a reference compiler, a peephole-augmented compiler, and two versions of SEQC—on simulated backends ranging from 1 to 100 modules.
+All compilers are implemented in the Qiskit SDK (qiskit2024) v1.2.41.2.4 and Python $3.12.6$, and timing results are taken on a machine with a $128$-core Ampere Altra Max ARM64 processor and $256$ GB of DDR4 RAM.
 For our reference compiler, we use the one built into Qiskit, parameterized with optimization_level=3.
-Because Qiskit is not natively chiplet-aware, we run these experiments on a backend with universal inter-chiplet links (§4.2).
-To construct the peephole compiler we augment the reference Qiskit compiler by inserting our peephole pass between its routing and basis_translation stages, thus making it chiplet-aware.
+Because Qiskit is not natively module-aware, we run these experiments on a backend with universal inter-module links (§4.2).
+To construct the peephole compiler we augment the reference Qiskit compiler by inserting our peephole pass between its routing and basis_translation stages, thus making it module-aware.
 Our two SEQC versions differ in the technique they use for qubit-to-subcircuit mapping.
 Our first SEQC implementation leverages simulated annealing, initialized with a starting temperature $T_{0}=200$, a rate of change of $0.005T$ per round, $\frac{T}{50}$ permutations per round, and a total of $10$ trials per CPU core.
-Meanwhile, our second SEQC implementation leverages STA for subcircuit-to-qubit mapping and subcircuit-to-chiplet mapping.
-We aggregate results in a manner similar to the SPEC family of benchmarks (spec2017), by taking the geometric mean of the relative performance ratio against a baseline technique across a benchmark suite of quantum circuits.
-The resultant score is then derived for each of the evaluated metrics on multiple backends of varying numbers of chiplets.
+Meanwhile, our second SEQC implementation leverages STA for subcircuit-to-qubit mapping and subcircuit-to-module mapping.
+We aggregate results in a manner similar to SPEC benchmarks (spec2017), by taking the geometric mean of the relative performance ratio against a baseline technique across a benchmark suite of quantum circuits.
+The resultant score is then derived for each of the evaluated metrics on multiple backends of varying numbers of modules.
 4.1. Benchmark Suite
+We leverage circuits from Supermarq (supermarq) that are representative of practical, real-world applications for quantum computing, while also being feasible to compile for a large number of qubits.
+In particular, our suite comprises the following quantum circuits.
+[…]
+[…]
 (a) BitCode
@@
 (d) VQE
-(e) Hamiltonian Simulation
+(e) Hamiltonian Sim.
 Figure 5. Examples of benchmark applications for 5-qubit circuits.
-We leverage circuits from Supermarq (supermarq) that are representative of practical, real-world applications for quantum computing, while also being feasible to compile for a large number of qubits.
-In particular, our suite comprises the following quantum circuits.
 […]
-[…]
-Greenberger–Horne–Zeiling (GHZ).
-Entanglement is one of the fundamental properties of quantum states and a major source of quantum advantage.
-The GHZ state represents a maximally entangled $n$-qubit state of $\frac{1}{\sqrt{2}}\left(\ket{|}0>^{\otimes n}+\ket{|}1>^{\otimes n}\right)$.
-The algorithmic circuit structure is essentially a single chain of CNOT gates (Figure 5(c)) From a compilation perspective, this represents a pure exercise in qubit allocation, because the logical circuit is entirely serial.
-An $n$-qubit GHZ circuit that fully utilizes a $k$-chiplet device should necessitate $\lfloor\frac{n}{k}\rfloor-1$ logical CNOT operations to extend between chiplets.
-In turn, this will necessitate at least $2\left(\lfloor\frac{n}{k}\rfloor-1\right)$ inter-chiplet SWAPs to transfer qubit in and out of a chiplet.
 VQE.
@@
 4.2. Simulated Quantum Device Specifications
-We perform our experiments on mock backends that are generated to conform to the chiplet architecture proposed in Smith et al.
-(smith_2022_chiplets).
-The hardware specifications of the qubits and intra-chiplet connections are sourced from Acharya et al.
-(backend_specs).
-Table 1 details the full specifications.
-To generate the expected fidelity of the inter-chiplet SWAPs, we perform a simulated random benchmarking trial and calculate the error rate of a SWAP gate with these backend specifications.
-Following Smith et al.
-(smith_2022_chiplets), we model inter-chiplet SWAPs on this backend with 4×4$\times$4 × the intra-chiplet error rate.
-Similarly, we approximate the duration of the inter-chiplet SWAP gate as 4×4$\times$4 × the duration of an intra-chiplet SWAP.
+(a) Small-scale heavy-hex modules, on grid lattice.
+(b) Large-scale modules with IBM Nighthawk-like topology.
+Figure 6. Modeled topologies.
 T1 (backend_specs) $20\times 10^{-6}$ seconds
@@
 Instruction Duration (ns) Error
-(backend_specs) $25.0$ $0.109\%$
+X (backend_specs) $25.0$ $0.109\%$
 SX (backend_specs) $25.0$ $0.109\%$
 $\text{R}_{\text{z}}(\phi)$ (backend_specs) $0.0$ $0.000\%$
-CZ (intra-chiplet) (backend_specs) $34.0$ $0.605\%$
-CZ (inter-chiplet) $126.0$ $2.420\%$
-SWAP (inter-chiplet) $702.4$ $10.23\%$
+CZ (intra-module) (backend_specs) $34.0$ $0.605\%$
+CZ (inter-module) $126.0$ $2.420\%$
+SWAP (inter-module) $702.4$ $10.23\%$
 Reset (backend_specs) $500.0$ $0.186\%$
@@
 Universal gate set only. Non-universal gate set only.
-Table 1. Specifications of Simulated Chiplet Backend.
-The simulated device has a grid lattice topology of symmetric chiplet modules, where each chiplet features a heavy-hexagon qubit topology (heavy_hex) (Figure 6).
-The intra-chiplet topology is constructed such that the heavy-hex structure is preserved for the global device topology (ignoring chiplet division).
-The tessellation requirements of each chiplet restrict valid chiplet sizes to multiples of $10$; however, for our experiments, we limit the size of each chiplet to the minimum of $10$ qubits due to hardware limitations.
-Finally, note that for each valid chiplet count, there may exist multiple valid chiplet topologies.
-In this work, we generate our backends to use the “most-square” topology for a given chiplet count.
-For example, a $12$-chiplet machine would be generated with a grid of $3\times 4$ chiplets.
-Figure 6. Construction of a grid inter-chiplet lattice from a heavy-hexagon qubit lattice, where the heavy-hex lattice is maintained for each chiplet.
+Table 1. Specifications of Simulated Modular Backend.
+We model two types of simulated devices differing in their inter-module topology and module size.
+To isolate the impact of architecture modularity, we evaluated on symmetrical modules based on a transmon qubit technology.
+One variant, denoted the heavy-hex backend (Figure 6(a)), is generated to conform to the modular architecture proposed in Smith et al.
+(smith_2022_chiplets), and features a 10-qubit, heavy-hexagon topology (heavy_hex).
+The other, denoted the grid backend (Figure 6(b)), uses modules with a 120-qubit grid topology, based on the upcoming IBM Nighthawk processor (IBM_2025_FTQC; Nighthawk_topology).
+Both simulated devices use a grid lattice topology of symmetric modules.
+The hardware specifications of the qubits and intra-module connections (detailed in Table 1) are sourced from Acharya et al.
+(backend_specs).
+To generate the expected fidelity of the inter-module SWAPs, we perform a simulated random benchmarking trial and calculate the error rate of a SWAP gate with these backend specifications.
+Following Smith et al.
+(smith_2022_chiplets), we conservatively model inter-module SWAPs with $4\times$ the error rate and $4\times$ the duration of an intra-module SWAP.
+Finally, note that for each valid module count, there may exist multiple valid module topologies.
+In this work, we generate our backends to use the “most-square” topology for a given module count.
+For example, a $12$-module machine would be a $3\times 4$ grid.
 5. Experimental Results
-We present results for the Qiskit reference compiler, the peephole-augmented Qiskit (PA-Qiskit) compiler, the SEQC compiler using simulated annealing (SEQC-AN), and the SEQC compiler using STA (SEQC-STA).
+We evaluate the Qiskit reference compiler, the peephole-augmented Qiskit (PA-Qiskit) compiler, SEQC using simulated annealing (SEQC-AN), and SEQC using STA (SEQC-STA) on metrics relating to the quality of the compiled circuits, on both the 10-qubit heavy-hex backend, and the 120-qubit grid backend.
 The Qiskit-based and SEQC compilers prioritize different heuristics in compilation.
 We examine how well each compiler is able to minimize its preferred heuristics, and how well that translates to practical solution quality.
-Finally, we analyze compilation time, specifically the recurring and non-recurring compilation times for repeating compilations of a single application.
-Number of inter-chiplet gates
-Figure 7. Relative number of inter-chiplet gates for SEQC normalized to the Qiskit baseline (lower is better).
-The primary heuristic we minimize in SEQC is the number of inter-chiplet gates in a quantum circuit.
-This is because we expect that the inter-chiplet gates of a chiplet architecture contribute the most to error and execution time relative to other gates (at least by a factor of 4×4$\times$4 × (smith_2022_chiplets)).
-As a result, the number of inter-chiplet gates should represent a strong proxy metric for the quality of a compiler solution, and thus minimizing the number of inter-chiplet gates should result in higher-fidelity, faster-executing quantum circuits.
-Accordingly, SEQC consistently produces results that have significantly fewer inter-chiplet gates compared to Qiskit-based compilers (Figure 7).
-Specifically, compilation with SEQC-STA results in $82.43\%$ fewer inter-chiplet gates on average compared to Qiskit (up to $92.94\%$).
-Compared to PA-Qiskit, compilation with SEQC-STA results in $64.92\%$ fewer inter-chiplet gates on average, up to $78.62\%$.
-It is worth mentioning that the inter-chiplet gates in the universal backend would be CZ gates, while the non-universal backend would be SWAP gates.
-For every inter-chiplet (CZ) gate in Qiskit, PA-Qiskit should have 2 inter-chiplet (SWAP) gates.
-However, we observe PA-Qiskit produces circuits with a roughly comparable number of inter-chiplet gates to Qiskit.
-In fact, as size increases, PA-Qiskit tends to result in fewer inter-chiplet gates than Qiskit.
+To demonstrate scalability, we analyze compilation time, specifically the recurring and non-recurring compilation times for repeating compilations of a single application, as well as memory requirements for the compilation.
+Relative metrics are normalized over the Qiskit reference compiler.
+Number of inter-module gates
+The primary heuristic we minimize in SEQC is the number of inter-module gates in a quantum circuit.
+This is because we expect that the inter-module gates of a modular architecture contribute the most to error and execution time relative to other gates (at least by a factor of $4\times$ (smith_2022_chiplets)).
+As a result, the number of inter-module gates should represent a strong proxy metric for the quality of a compiler solution, and thus minimizing the number of inter-module gates should result in higher-fidelity, faster-executing quantum circuits.
+Accordingly, SEQC consistently produces results that have significantly fewer inter-module gates compared to Qiskit-based compilers (Figures 7(a) and 8(a)).
+For example, compilation with SEQC-STA results in $82.43\%$ fewer inter-module gates on average compared to Qiskit (up to $92.94\%$).
+on the heavy-hex backend.
+Meanwhile, compilation with SEQC-STA results in $96.31\%$ fewer inter-module gates on average compared to Qiskit (up to $98.32\%$) on the grid backend.
+Compared to PA-Qiskit, compilation with SEQC-STA results in $64.92\%$ fewer inter-module gates on average, up to $78.62\%$, on the heavy-hex backend.
+Compared to PA-Qiskit, compilation with SEQC-STA results in $94.31\%$ fewer inter-module gates on average, up to $97.83\%$ on the grid backend.
+It is worth mentioning that the inter-module gates in the universal backend would be CZ gates, while the non-universal backend would be SWAP gates.
+For every inter-module (CZ) gate in Qiskit, PA-Qiskit should have 2 inter-module (SWAP) gates.
+However, we observe PA-Qiskit produces circuits with a roughly comparable number of inter-module gates to Qiskit.
+In fact, as size increases, PA-Qiskit tends to result in fewer inter-module gates than Qiskit.
 This outcome may indicate the presence of optimizations that are removing extraneous SWAPs.
+(a) Relative inter-module gate count
+(lower is better).
+(b) Relative estimated fidelity (ESP)
+(higher is better).
+Figure 7. Main circuit performance metrics: inter-module gate count and estimated fidelity, 10-qubit heavy-hex backend.
+(a) Relative inter-module gate count
+(lower is better).
+(b) Relative estimated fidelity (ESP)
+(higher is better).
+Figure 8. Main circuit performance metrics: inter-module gate count and estimated fidelity, 120-qubit grid backend.
+Fidelity
+Practical applications of quantum computing require some guarantee of program output correctness (measurements).
+Thus, the highest priority for quantum circuit compilation is to maximize fidelity.
+We estimate fidelity through estimated success probability (ESP) (esp).
+Specifically, we track the per-qubit ESP of a quantum circuit and report an averaged ESP for each benchmark from the geometric mean of the circuit measurements.
+Figures 7(b) and 8(b) presents the relative estimated fidelity of each compiler compared to the Qiskit compiler with respect to circuit size.
+There is a noticeable benefit in fidelity associated with the SEQC compiler, particularly with SEQC-STA.
+On the heavy-hex backend, we observe an average improvement of $22.13\%$ over PA-Qiskit, and up to $62.87\%$.
+On the grid backend, SEQC-compiled circuits achieve $3.51\%$ higher fidelity on average ($22.50\%$ max) over PA-Qiskit.
+SEQC similarly outperforms Qiskit by a wide margin.
+It is important to emphasize here that SEQC outperforms Qiskit even when the latter is afforded significant allowances (no gate restrictions, and much lower inter-module gate errors).
+The Qiskit trial backend supports universal inter-module links, in contrast to the SEQC backend, and with only a $2.42\%$ error for inter-module CZs, compared to $10.23\%$ error for SEQC’s inter-module SWAPs.
+All these allowances should translate to higher fidelity solutions for Qiskit.
+Nevertheless, even though much more constrained, SEQC-STA demonstrates an average improvement of $9.30\%$, up to a $49.99\%$, on the heavy-hex backend, and $32.30\%$ ($63.36\%$ max) on the grid backend.
+(a) Relative total gate count.
+(b) Relative circuit depth.
+(c) Relative per-shot execution time.
+Figure 9. Total gate count, circuit depth, and per-shot execution time, 10-qubit heavy-hex backend (lower is better).
+(a) Relative total gate count.
+(b) Relative circuit depth.
+(c) Relative per-shot execution time.
+Figure 10. Total gate count, circuit depth, and per-shot execution time, 120-qubit grid backend (lower is better).
 Gate count and circuit depth.
-Figure 8. Relative gate count of SEQC normalized to the Qiskit baseline for devices (lower is better).
-Figure 9. Relative circuit depth of SEQC normalized to the Qiskit baseline for devices (lower is better).
 Gate count and circuit depth are static, backend-agnostic metrics that can serve as indicators for a quantum circuit’s expected fidelity and/or execution time, where circuit depth is the critical path of a quantum circuit in terms of individual gates (mike_and_ike).
 SABRE (zou2024lightsabrelightweightenhancedsabre; sabre_swap), moreover, explicitly minimizes gate count by incorporating it into its cost functions.
-We observe that SEQC typically produces results with higher gate count and circuit depth compared to Qiskit and PA-Qiskit (Figures 8 and 9).
-This is due to SEQC prioritizing the reduction of the number of inter-chiplet gates over total gate count or circuit depth.
+We observe that SEQC typically produces results with higher gate count (Figures 9(a) and 10(a)) and circuit depth (Figures 9(b) and 10(b)) compared to Qiskit and PA-Qiskit.
+This is due to SEQC prioritizing the reduction of the number of inter-module gates over total gate count or circuit depth.
 Nevertheless, SEQC-STA, in particular, is able to demonstrate competitive results in these metrics for larger circuits.
-In terms of relative circuit depth, SEQC-STA has comparable or slightly better results compared to PA-Qiskit at $500$ qubits (50 chiplets) and above.
-In terms of relative gate count, SEQC-STA has comparable or slightly better results compared to PA-Qiskit at $400$ qubits (40 chiplets) and above.
-For the $810$ and $1000$ qubit ($81$ and $100$ chiplet) trials, it achieves $14.11\%$ and $8.87\%$ gate count reduction over Qiskit, respectively, despite the latter using universal inter-chiplet links.
-Fidelity
-Figure 10. Relative estimated fidelity (ESP) of our proposed compilers normalized to the Qiskit baseline (higher is better).
-Practical applications of quantum computing require some guarantee of program output correctness (measurements).
-Thus, the highest priority for quantum circuit compilation is to maximize fidelity.
-Fidelity is impacted by the decoherence constraints of individual qubits, as well as the errors on a gate-by-gate basis (mike_and_ike).
-Due to hardware limitations, we were unable to estimate fidelity though simulation.
-Instead, we leverage estimated success probability (ESP) (esp) to approximate fidelity, which can be computed statically.
-Specifically, we track the per-qubit ESP of a quantum circuit and report an averaged ESP for each benchmark from the geometric mean of the circuit measurements.
-Figure 10 presents the relative estimated fidelity of each of our compilers compared to the Qiskit compiler with respect to circuit size.
-There is a noticeable benefit in fidelity associated with the SEQC compiler, particularly with SEQC-STA.
-We observe an average improvement of $22.13\%$ over PA-Qiskit, up to a $62.87\%$ improvement.
-Moreover, it is important to mention that the Qiskit trial was ran on a backend with universal inter-chiplet links, which should translate to higher fidelity solutions.
-Nevertheless, SEQC-STA demonstrates an average improvement of $9.30\%$, up to a $49.99\%$ improvement.
+SEQC-STA has comparable or slightly better results compared to PA-Qiskit at $500$ qubits (50 modules) and above for both metrics on the heavy-hex backend.
+For the $810$ and $1000$ qubit ($81$ and $100$ module) trials on the heavy-hex backend, SEQC-STA achieves $14.11\%$ and $8.87\%$ gate count reduction over Qiskit, respectively, despite the latter using universal inter-module links.
+For the grid backend, we observe similar behavior to the heavy-hex backend, where SEQC produces increasingly competitive results as the circuit size increases.
+However, we did not perform any trails large enough to witness the point where SEQC-STA would begin to produce better results than the Qiskit baselines
 Estimated execution time.
-Figure 11. Estimated per-shot execution time of circuits produced by SEQC normalized to circuits produced by the Qiskit baseline (lower is better).
 The per-shot execution time of a quantum circuit is determined by the critical path of gate operations weighted by gate duration.
@@
 In other words, it is imperative that quantum circuits execute within the decoherence time to ensure high-fidelity outcomes.
-Figure 11 presents the estimated per-shot execution time of our proposed compilers relative to the reference as a function of the number of qubits per circuit.
-As explained previously, we approximate the duration of an inter-chiplet SWAP as 4×4$\times$4 × the duration of an intra-chiplet SWAP (smith_2022_chiplets).
-Under this model, we observe that the SEQC compilers consistently outperform PA-Qiskit in execution by $24.19\%$ on average ($63.98\%$ max).
-It is notable that, despite the gate set affordances given to the Qiskit compiler, SEQC-STA is able to achieve parity in execution time starting from $600$ qubits ($60$ chiplets).
-For $810$ qubits ($81$ chiplets) and above, SEQC-STA actually compiles faster-executing quantum circuits.
+Figures 9(c) and 10(c) presents the estimated per-shot execution time of our proposed compilers relative to the reference as a function of the number of qubits per circuit.
+As explained previously, we approximate the duration of an inter-module SWAP as $4\times$ the duration of an intra-module SWAP (smith_2022_chiplets).
+Under this model, we observe that the SEQC compilers consistently outperform PA-Qiskit in execution by $24.19\%$ on average ($63.98\%$ max) on the heavy-hex backend.
+It is notable that, despite the gate set affordances given to the Qiskit compiler, SEQC-STA achieves parity in execution time starting from $600$ qubits ($60$ modules).
+For $810$ qubits ($81$ modules) and above, SEQC-STA actually compiles faster-executing quantum circuits.
+Meanwhile, on the grid backend, SEQC-STA-compiled circuits execute up to $151.29\%$ faster over PA-Qiskit.
+Circuit execution time for SEQC-STA-compiled circuits is also faster than the Qiskit baseline from $480$ qubits ($4$ modules) onwards by $28.04\%$ on average, up to $54.82\%$, with SEQC-SA trailing close behind.
 Compilation time
-Figure 12. Compilation time of our evaluated compilers (lower is better).
-Figure 13. Relative recurring compilation time of our proposed compilers normalized to the Qiskit baseline (lower is better).
-Near-term quantum devices feature rapidly fluctuating error characteristics that can change hourly, or even on a run-to-run basis (Tannu_2019_ErrorVariability; Murali_2019_NoiseMapping; dasgupta2021stability).
+Quantum devices feature rapidly fluctuating error characteristics that can change hourly, or even on a run-to-run basis (Tannu_2019_ErrorVariability; Murali_2019_NoiseMapping; dasgupta2021stability).
 Thus, across the entire execution lifetime of an application (quantum circuit), regular re-compilation is necessary to generate the highest-quality circuits with the most up-to-date characteristics possible for every execution.
-Traditional quantum compilers, like Qiskit and PA-Qiskit, approach each compilation tabula rasa, where each new compilation restarts from the original algorithmic circuit.
+Traditional quantum compilers, like Qiskit and PA-Qiskit, approach compilation tabula rasa, where each new compilation restarts from the original algorithmic circuit.
 It is challenging to speed up repeated recompilations without fundamental changes to their design, as noise-aware routing and synthesis algorithms, which leverage updated error characteristics, usually function from a device-agnostic starting point.
@@
 Nevertheless, although stratification is a one-time cost for the SEQC compiler, it must be scalable in order to be practical.
-In Figure 12, we plot the non-recurring compilation time of our evaluated compilers.
+In Figures 11(a) and 12(a), we plot the non-recurring compilation time of our evaluated compilers.
 For our SEQC compilers, we plot stratification time and stratification + elaboration time.
@@
 We demonstrate that all compilation times obey a quadratic trajectory ($R^{2}\geq 0.9883$).
-Figure 13 shows the recurring compilation time.
-Specifically, this refers to the elaboration time of the SEQC compilers and the full compilation time for Qiskit and PA-Qiskit.
+Figures 11(b) and 12(b) show the recurring compilation time.
+Specifically, this refers to the elaboration time of SEQC and the full compilation time for Qiskit and PA-Qiskit.
 Overall, SEQC compilers have a large speedup over Qiskit-based compilers, owing to the use of parallelism and modular compilation strategies.
-Compared to Qiskit, SEQC-STA achieves 3.27×3.27$\times$3.27 × speedup on average (up to 6.74×6.74$\times$6.74 ×), and compared to PA-Qiskit 4.69×4.69$\times$4.69 × average speedup (6.61×6.61$\times$6.61 × max).
+Compared to Qiskit on the heavy-hex backend, SEQC-STA achieves $3.27\times$ speedup on average (up to $6.74\times$), and compared to PA-Qiskit on the heavy-hex backend, $4.69\times$ average speedup ($6.61\times$ max).
+Similarly on the grid backend, SEQC-STA achieves $1.34\times$ speedup on average (up to $3.37\times$) over Qiskit, and $2.52\times$ speedup on average (up to $11.93\times$) over PA-Qiskit.
+(a) Compilation time.
+(b) Relative recurring compilation time.
+(c) Relative peak memory utilization.
+Figure 11. Compilation performance metrics: compilation time, recurring compilation time, and peak memory utilization, 10-qubit heavy-hex backend (lower is better).
+(a) Compilation time.
+(b) Relative recurring compilation time.
+(c) Relative peak memory utilization.
+Figure 12. Compilation performance metrics: compilation time, recurring compilation time, and peak memory utilization, 120-qubit grid backend (lower is better).
+Memory Utilization.
+Figures 11(c) and 12(c) presents the relative peak memory utilization of the evaluated compilers.
+Compared to Qiskit, SEQC-SA exhibits $55.68\%$ higher memory usage on average ($117.71\%$ max) on the heavy-hex backend and $23.71\%$ higher memory usage on average ($38.24\%$ max) on the grid backend.
+This could likely be ameliorated by reducing the number of annealing trials, at the expense of circuit quality.
+SEQC-STA tends to use less memory compared to the Qiskit variants, by $4.22\%$ on average ($17.47\%$ max) on the heavy-hex backend and $10.69\%$ on average ($17.51\%$ max) on the grid backend.
+It is possible that the process of stratification and parallelized elaboration is responsible for the conservation of memory usage, presuming the memory requirements of quantum compilation scales super-linearly with the circuit size.
+This points to the scalability benefits of the SEQC framework in general.
 6. Discussion
@@
 Circuit depth and gate count only consider the number of gates (in the critical path and in total) irrespective of each one’s unique underlying properties.
-It is expected that two-qubit gates will be significantly more noisy than single-qubit gates, and inter-chiplet gates should be significantly more noisy than intra-chiplet gates (smith_2022_chiplets).
-Failure to account for duration and error differences between gates or types of gates in quantum compilation heuristics will inevitably result in sub-optimal solutions when the discrepancies are extreme, e.g., in chiplet architectures.
-Conversely, the hierarchical structure of the SEQC framework with the stratification and elaboration stages imparts some degree of implicit noise-awareness, where minimizing the number of inter-chiplet gates takes precedence over intra-chiplet/total gates.
-As a result, SEQC can better optimize for the practically significant metrics of circuit fidelity and circuit execution time.
+It is expected that two-qubit gates will be significantly more noisy than single-qubit gates, and inter-module gates should be significantly more noisy than intra-module gates (smith_2022_chiplets).
+Failure to account for duration and error differences between gates or types of gates in quantum compilation heuristics will inevitably result in sub-optimal solutions when the discrepancies are extreme, e.g., in modular architectures.
+Conversely, the hierarchical structure of the SEQC framework with the stratification and elaboration stages imparts some degree of implicit noise-awareness, where minimizing the number of inter-module gates takes precedence over intra-module/total gates.
+As a result, SEQC can better optimize for the practically significant metrics of circuit fidelity and execution time.
 Overall, we observe that SEQC yields circuits with more desirable characteristics when executing on modular quantum architectures, while at the same time it achieves faster compilation times.
-Also, as the number of qubits increases, the performance improvement of SEQC widens compared to the baseline, especially for the most important metrics of fidelity, compiled circuit execution time, and number of inter-chiplet gates.
-These trends hold promise in allowing for faster and better compilation for future large-scale modular quantum devices than we can achieve with today’s frameworks.
+Also, as the number of qubits increases, the performance improvement of SEQC widens compared to the baseline, especially for the most important metrics of fidelity, compiled circuit execution time, and number of inter-module gates.
+These trends hold promise in allowing for faster and better compilation for future large-scale, hybrid modular quantum devices than we can achieve with today’s frameworks.
+It is also important to note that SEQC is equally applicable to NISQ as well as future quantum systems.
+The stratification stage can decompose a large circuit into subcircuits that will be mapped to QPUs at the logical level, and then the elaboration stage can compile for each QPU together with an error correction method for that QPU.
+The two stages are distinct, and stratification operates more on the logical level, while the elaboration can be tasked with lowering from the logical to the physical qubits.
+Applicability to NISQ systems is still important, as even with technological breakthroughs, the transition to Fault Tolerant Quantum Computing (FTQC) will be gradual, not instantaneous.
+The hardware that exists now will continue improving and will be widely used for years, as there is practical scientific and industrial value in NISQ, even without the need to wait for fully error-corrected systems (piccinelli_chemistry; Kim2023preFTUtility; sciadv_utility_2025; Abanin2025ConstructiveInterference).
+Moreover, early fault-tolerant demonstrations will run on NISQ hardware, and will heavily rely on NISQ-era methods, as only few qubits can be protected.
+Thus, it remains critical to design architectures and compilers that serve NISQ hardware, both to extract near-term value, and to pave the engineering path toward full FTQC.
 7. Related Work
-MECH (mech) sacrifices some ancillary qubit resources to build an ”adjustable multi-entry communication highway” through software.
-By entangling ancillary qubits and distributing them throughout a chiplet, MECH virtualizes those qubits into fast network-on-chip routing channels.
+The state-of-the-art in distributed quantum computing (DQC) (autocomm; 10228915; 10.1145/3730585; 10821229; 10313645; 10.1145/3579367) requires the generation and distribution of EPR pairs as the main entanglement resource over long-distance optical links, and the deployment of emitter-emitter, emitter-scatterer, or scatterer-scatterer photon-mediated entanglement schemes.
+However, modular architectures have less restrictive distance constraints, and they can often support other hardware technologies, such as microwave links (PhysRevApplied.21.054063).
+As a result, SEQC can model inter-module allocation topologically (as SWAP chains) (LaRacuente2025modelingshortrange), similar to compilation for monolithic (single-die) quantum processors.
+An analogy can be drawn to classical compilation between chiplet-based classical systems and distributed classical systems, where diverse compilation strategies are required due to very different hardware technologies and distances between computation units.
+Compilation for modular transmon systems shares many similar constraints with compilation for quantum charge-coupled devices (QCCDs).
+Namely, these devices features modules, in the form of ion traps, and support movement between traps, via ion shuttling mechanisms (ovide2024sta).
+However, there are some key differences between the hardware technologies that make it difficult to compare SEQC with QCCD compilers on equal footing.
+Notably, qubits in superconducting devices needs to be topologically adjacent to each another to support two-qubit operations.
+However, QCCD devices are less restrictive, and support two-qubit operations so long as qubits are in the same traps / interaction region.
+As a result, QCCD compilers (10.1145/3526241.3530366; Kreppel2023quantumcircuit) can employ simpler routing strategies, or forgo them entirely, compared to modular transmon compilers.
+MECH (mech) entangles ancillary qubits and distributes them throughout a module to build fast network-on-chip routing channels (a ”communication highway”).
+MECH includes a compiler that requires, and optimizes for, the “highway” design.
+Unfortunately, the highway does not generalize to non-universal inter-module links.
+Future work to extend SEQC with MECH’s methodology may enable combining MECH’s techniques with SEQC’s allocation procedures to gain the benefits of each.
 Lin et al.
@@
 Escofet et al.
-(escofet2024revisiting) develop a characterization technique for arbitrary quantum circuits which yields theoretical upper and lower optimization bounds, enabling the optimality assessment of qubit mapping solutions.
-All these techniques are orthogonal to SEQC, and could work synergistically with it to achieve the combined benefits.
-CutQC (cutqc) enables smaller-scale quantum computers to execute quantum circuits that are otherwise too large to fit on the device.
-It partitions large quantum circuits into smaller subcircuits, executes them one at a time, and then merge the individual results at the cost of classical post-processing.
-Piveteau and Sutter (circuit_knitting) similarly allow subcircuits to execute independently, and use quasiprobability simulation to simulate inter-chiplet link behavior where the subcircuits connect, reaching fixed accuracy with less simulation time.
+(escofet2024revisiting) develop a characterization technique for arbitrary quantum circuits, which provides theoretical upper/lower optimization bounds, enabling the optimality assessment of qubit mapping solutions.
+These techniques are orthogonal to SEQC, and could work synergistically with it.
+CutQC (cutqc) partitions large quantum circuits into smaller subcircuits that fit on a device, executes them one at a time, and then merges the individual results through classical post-processing.
+Piveteau and Sutter (circuit_knitting) similarly allow subcircuits to execute independently, and use quasiprobability simulation to simulate inter-module link behavior where the subcircuits connect, reaching fixed accuracy with less simulation time.
 SEQC’s stratification stage was inspired by CutQC, but unlike these works SEQC does not perform any classical “stitching” or simulation.
-It is interesting to note that the first version of SEQC was formulating the problem using constrained optimization as well, but the slowdown of the stratification step was unacceptable and impractical for large quantum circuits, and thus was abandoned in favor of the heuristics we currently employ in our compiler.
+We initially formulated qubit-to-subcircuit mapping as a constraint optimization problem, similar to CutQC, but existing numerical solvers were impractically slow for circuits larger than 20 qubits, corroborating the experience of Baker et al.
+(baker2020).
+Instead, we opted to use the approximate, but practical, solutions we currently employ in SEQC.
 Bandic et al.
@@
 Instead, SEQC takes a “best-effort” approach that works well in practice, and does not require acceleration through quantum means.
-Some works on qubit mapping share a few characteristics with SEQC.
-In a bottom-up approach to qubit mapping, Liu et al.
+Liu et al.
 (liu_2023_tackling) partition a circuit into blocks that span an identical number of qubits, perform permutation-aware synthesis on each block in parallel (solving for all possible qubit topologies), and finally use an augmented SABRE algorithm to map the blocks to the target device and route between blocks.
-SEQC takes a top-down approach instead, where SWAP permutations between chiplets are settled prior to any compilation tasks performed for each chiplet (block).
+SEQC takes a top-down approach instead, where SWAP permutations between modules are settled prior to any compilation tasks performed for each module (block).
 As a result, SEQC only needs to solve for one qubit layout per block, rather than all permutations.
@@
 Pastor et al.
-(pastor2024circuit) use deep reinforcement learning to partition quantum circuits for execution across multiple chiplets.
-Both works aim to reduce inter-chiplet swaps, but assume that all inter-chiplet communications have the same cost, i.e., chiplets are fully connected, which is both impractical for large-scale systems and misses optimization opportunities.
-SEQC, in contrast, can work with arbitrary chiplet topologies.
+(pastor2024circuit) use deep reinforcement learning to partition quantum circuits for execution across multiple modules.
+Both works aim to reduce inter-module swaps, but assume that all inter-module communications have the same cost, i.e., modules are fully connected, which is both impractical for large-scale systems and misses optimization opportunities.
+SEQC, in contrast, can work with arbitrary module topologies.
 Ovide et al.
-(ovide2023mapping) leverage both quantum and classical methods for inter-chiplet communication to address the scalability challenges of monolithic processors.
-While it advocates for a two-level, hierarchical approach for quantum circuit mapping, it leaves the actual problem of qubit placement and routing unsolved.
-SEQC provides an algorithm and implementation, addressing both challenges.
-Parallelizing SEQC’s recurrent elaboration step was inspired by traditional classical compilation techniques, which has a decades-long history and is now commonplace (Mickunas_1978_ParallelCompilation; Gross_1989_ParallelCompilation).
-Besides traditional compilation, though, we can also draw parallels between our work and prior research on data movement in the classical space.
-Data movement is widely understood to be the primary performance bottleneck in the classical architecture space.
-In the last two decades, work has proliferated on dataflow and spatial architectures, much of which focuses on mapping problems.
-Consider neural network architectures, for example, which seek to map their physical topologies as closely as possible to problem topologies.
-Eyeriss (Eyeriss_2016_Chen) approaches the dataflow mapping problem by proposing a row-stationary architecture, and then optimizing the mapping of problem to processing element such that elements which are spatially local in the problem are spatially local in hardware during execution.
-This is not unlike the target optimization for the stratification stage, and in fact is performed for many of the same reasons — like inter-chiplet swaps, data movements between more distant classical programming elements in a dataflow architecture are more expensive, and minimizing their frequency yields performance improvements.
+(ovide2023mapping) leverage both quantum and classical methods for inter-module communication.
+While it advocates for a two-level approach to quantum circuit mapping, it leaves the actual problem of qubit placement and routing unsolved, which SEQC addresses with its algorithm.
 8. Conclusion
-Physical, technological, and economic considerations force future quantum systems to adopt a physically distributed architecture.
-Modular quantum processors already dominate the roadmaps of several major quantum companies.
-However, quantum compilers face a scalability challenge brought by the complexity and increasing scale of chiplet-based quantum devices, and their lack of chiplet awareness impedes their ability to produce high-quality results.
-In this work, we first modify stock contemporary compilers to become aware of hardware modularity, thereby allowing them to correctly compile circuits for modular architectures with limited cross-chiplet gate support.
-We then propose SEQC, a Stratify-Elaborate Quantum Compiler for modular architectures.
-SEQC performs compilation in two stages, with the first stage (stratification, or chiplet splitting) performed only once for a given architecture, and the second stage (elaboration, or chiplet compilation) performed in parallel for each chiplet.
-Unlike contemporary quantum compilers that have to compile a quantum circuit anew before each execution to capitalize on the latest qubit calibration results, only the elaboration stage of SEQC needs to run before each execution.
-As its computational complexity is a function of the quantum chiplet size, not the entire quantum processor size, and chiplet sizes grow slowly, SEQC’s recurrent compilation time is largely unaffected by the relentless growth of qubit counts in future quantum processors.
-SEQC implements several novel methods for qubit placement, qubit routing, and circuit optimization, with both hardware modularity and compilation parallelization in mind, achieving on average $9.3\%$ higher circuit fidelity (up to $49.99\%$), while attaining on average 3.27×3.27$\times$3.27 × faster compilation (up to 6.74×6.74$\times$6.74 ×) compared to a chiplet-unaware Qiskit baseline.
+[…]
 Acknowledgements
@@
 We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.
-[…]